### v1

# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

Abstract
While Graph Neural Networks (GNNs) have proven highly effective at modeling relational data, pairwise connections cannot fully capture multi-way relationships naturally present in complex real-world systems. In response to this, Topological Deep Learning (TDL) leverages more general combinatorial representations –such as simplicial or cellular complexes– to accommodate higher-order interactions. Existing TDL methods often extend GNNs through Higher-Order Message Passing (HOMP), but face critical scalability challenges due to (i) a combinatorial explosion of message-passing routes, and (ii) significant complexity overhead from the propagation mechanism. To overcome these limitations, we propose HOPSE (Higher-Order Positional and Structural Encoder)—a message passing-free framework that uses Hasse graph decompositions to derive efficient and expressive encodings over arbitrary higher-order domains. Notably, HOPSE scales linearly with dataset size while preserving expressive power and permutation equivariance. Experiments on molecular, expressivity and topological benchmarks show that HOPSE matches or surpasses state-of-the-art performance while achieving up to 7× speedups over HOMP-based models, opening a new path for scalable TDL.
1 Introduction
Graph neural networks (GNNs) [46, 11] have demonstrated impressive capabilities in modeling relational data structured as graphs. However, their reliance on edges to encode interactions inherently limits them to pairwise relationships, which may not fully capture the complexity of many real-world systems. For instance, social dynamics often involve group interactions, while physical phenomena like electrostatic interactions in proteins can span multiple atoms simultaneously.
To address such limitations, Topological Deep Learning (TDL) [8, 23] introduces a framework for modeling multi-way interactions using higher-order topological domains such as hypergraphs, simplicial, cell, and combinatorial complexes. Within TDL, standard Topological Neural Networks (TNNs) [18, 12, 22, 21, 38] extend the graph-based message passing paradigm through Higher-Order Message Passing (HOMP), enabling information flow across all of the dimensions of the considered combinatorial representation. [^1] The terms higher-order topological domain and combinatorial representation are used intercheangeably: please see Section A.1 for the formal definitions. This generalization has shown strong performance across diverse fields—including social networks [30], protein biology [28], physics-informed learning [45], and computer networks [6, 5]—where interactions naturally extend beyond pairwise connections.
However, despite their theoretical advantages, TNNs face substantial challenges stemming from the HOMP paradigm. Unlike GNNs, TNNs’ ability to propagate information through the full hierarchy of a higher-order domain –where elements of different ranks (e.g., nodes, edges, faces, etc.) are interconnected– leads to an explosion in the number of possible sequences of message passing routes across these different levels [42, 33, 37]. More specifically, for a combinatorial complex $\mathcal{T}$ with dimension –i.e. maximum rank– $R$, a total of $(R+1)!$ message passing routes arises just considering communication flows between different ranks, and that number skyrockets to $(R+1)!\cdot(R+1)^{R}$ when allowing same-rank connections (please see Appendix B for a detailed analysis).
On top of this vast architectural search space, TNNs operating under the HOMP paradigm are also known to be computationally demanding, message propagation often requiring significant processing resources even for moderately sized inputs [37, 20]. This combination of combinatorial design complexity and high computational cost severely limits existing TNNs’ scalability, posing a major barrier to their deployment in large-scale or time-sensitive real-world applications.
In response to these challenges, this work introduces Higher-Order Positional and Structural Encoder (HOPSE), a novel framework designed to overcome the computational hurdles associated with HOMP. Specifically, HOPSE main features and contributions include:
  • Higher-order topological encodings: By exploiting Hasse graph decompositions –i.e. transformations of combinatorial representations into simpler graph ensembles [23, 38]–, HOPSE facilitates the efficient use of established graph positional and structural encoding techniques (e.g. [13]) to obtain analog representations for higher-order topological domains.
  • General message passing-free method: Inspired by SaNN model [20], the proposed encoding-based methodology effectively captures relational information without explicit message passing mechanisms. Moreover, HOPSE is no longer constrained to simplicial complexes; by design, it can be applied to any higher-order combinatorial representations.
  • Scalability and computational efficiency: In contrast to HOMP-based networks, HOPSE complexity scales linearly with respect to the number of elements in the combinatorial representation (Section 4.3). This is achieved by computing the encodings as a preprocessing step, significantly accelerating both training and inference (see Section 5 for experimental details).
  • Theoretical guarantees: Even without message passing, HOPSE is as expressive as existing state-of-the-art HOMP-based models (e.g. Papillon et al. [38]) in distinguishing non-isomorphic complexes (Theorem 4.1), while preserving cell permutation equivariance (Theorem 4.2).
  • Flexible instantiations: Two HOPSE realizations are presented—HOPSE-M, relying on handcrafted graph-theoretical encodings, and HOPSE-G, using a pretrained GPSE encoder—offering complementary trade-offs between preprocessing cost and representational richness (Section 4.3).
  • Extensive empirical validation: Comprehensive experiments on molecular benchmarks [34], CLS expressivity dataset [36], and MANTRA [2] –recent dataset consisting of pure topological tasks– validate the practical effectiveness of HOPSE, showcasing its ability to efficiently capture complex relational structures and outperform existing baselines –while being up to 7×7$\times$7 × faster than HOMP-based models (Section 5).
Additionally, Appendix B provides, for the first time, a detailed mathematical analysis of the combinatorial explosion of message passing routes in higher-order topological domains.
2 Background
This section introduces key concepts underpinning TDL. For brevity and generality, the following discussion centers on combinatorial complexes, which subsume all of the discrete topological domains leveraged by TDL [23]. Appendix A provides rigorous mathematical definitions of these other domains, such as simplicial and cellular complexes, along with an extended theoretical background.
Combinatorial complex.
A combinatorial complex ${\mathcal{T}}$ is a triple $(\mathcal{V},\mathcal{C},\textrm{rk})$ consisting of a set $\mathcal{V}$, a subset $\mathcal{C}$ of the powerset $\mathcal{P}(\mathcal{V})\backslash{\emptyset}$, and a rank function $\textrm{rk}:\mathcal{C}\rightarrow\mathbb{Z}_{\geq 0}$ with the following properties:
  1. 1.
    For all $v\in\mathcal{V},\{v\}\in\mathcal{C}$ and $\textrm{rk}(\{v\})=0$;
  2. 2.
    The function rk is order-preserving, i.e., if $\sigma,\tau\in\mathcal{C}$ satisfy $\sigma\subseteq\tau$, then $\textrm{rk}(\sigma)\leq$ $\textrm{rk}(\tau)$;
The elements of $\mathcal{V}$ represent the nodes, while the elements of $\mathcal{C}$ are called cells (i.e., groups of nodes). The cell $\sigma\in\mathcal{C}$ of rank $k:=\textrm{rk}(\sigma)$ is referred to as a $k$-cell. $\mathcal{C}$ simplifies notation for $(\mathcal{V},\mathcal{C},\textrm{rk})$. Lastly, the dimension of ${\mathcal{T}}$ is defined as the maximal rank among its cells: $\mathrm{dim}(\mathcal{C}):=\max_{\sigma\in\mathcal{C}}\textrm{rk}(\sigma)$.
Signals over combinatorial complexes.
A combinatorial complex $\mathcal{T}=(\mathcal{V},\mathcal{C},\mathrm{rk})$ can be equipped with signals, generalizing the concept of featured graph [^2] Attributed graph and featured graph are used interchangeably to comply with the literature. (see Definition A.1) to higher-order structures. In this context, a signal is a set of functions =$\mathcal{F}=\{F_{r}\}_{r\geq 0}$, where $F_{r}:\{\sigma\in\mathcal{C}\mid\mathrm{rk}(\sigma)=r\}\to\mathbb{R}^{D_{r}}$, where $D_{r}$ is the dimension of the feature space for rank-$r$ cells. For this work, it is important to emphasize that a complex can have multiple attribute functions over each $r$-cell. More formally, =$\mathcal{F}=\{F_{r,k}\}_{k=1}^{K}$ for all $r\geq 0$, where each $F{r,k}$ shares the same domain: $\text{Dom}(F_{r,k})=\text{Dom}(F_{r})$.
Neighborhood relations: adjacencies and incidences.
Combinatorial complexes can be equipped with a notion of neighborhood among cells that induces rich topological structures. Formally, denoting by $\mathcal{P}(\cdot)$ the power set, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a complex ${\mathcal{T}}$ maps each cell $\sigma\in\mathcal{C}$ to a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$. Traditional neighborhood functions include adjacencies ($\mathcal{A}_{t,s}$), connecting cells of the same rank $t$ –through adjacency relations w.r.t. cells of another rank $s$–, and incidences ($\mathcal{I}_{s\to t}$), linking cells of different ranks –from a source rank $s$ to a target rank $t$. Figure 1 presents some instances, and Section A.2 provides formal definitions.
Figure 1: Neighborhood expansions of a complex. Given a complex ${\mathcal{T}}$ (left), three examples of strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}}$ are illustrated corresponding to 4 neighborhoods: adjacency of nodes w.r.t. edges (${\mathcal{A}}_{0,1}$), incidence from nodes to edges (${\mathcal{I}}_{0\to 1}$), incidence from faces to nodes (${\mathcal{I}}_{2\to 1}$), and incidence from nodes to faces (${\mathcal{I}}_{0\to 2}$).
Strictly augmented Hasse graphs.
Given a complex ${\mathcal{T}}$, each particular neighborhood $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ [39], defined as the directed graph whose nodes and edges are given, respectively, by
$\mathcal{C}_{{\mathcal{N}}}=\{\sigma\in{\mathcal{C}}\mid{\mathcal{N}}(\sigma)\neq\emptyset\}\quad\text{and}\quad\mathcal{E}_{{\mathcal{N}}}=\{(\tau,\sigma)\mid\sigma,\tau\in{\mathcal{C}}_{{\mathcal{N}}},\ \tau\in{\mathcal{N}}(\sigma)\}.$
Figure 1 visually shows some examples of these neighborhood graph expansions. Notably, “strictly” comes from the fact that the set of nodes of $\mathcal{G}_{{\mathcal{N}}}$ is restricted to $\mathcal{C}_{{\mathcal{N}}}$, i.e., to those cells that have at least one neighbor according to ${\mathcal{N}}$. In contrast to this, the original definition of augmented Hasse graph considers the full set of cells $\mathcal{C}$ [23].
3 Related works
Positional and Structural Encodings in GNNs.
In the graph domain, a range of positional and structural encodings (PSEs) have been proposed to enhance node representations. Positional encodings typically capture a node’s location within the graph (e.g., via distances or coordinates), while structural encodings describe the node’s role or function in the topology (e.g., node degrees). Classical approaches of positional encodings include spectral methods based on Laplacian eigenvectors (LapPE) [17], while a common structural encoding is the random-walk embedding (RWSE) [15, 43], which reflects multi-hop connectivity; these handcrafted features help differentiate nodes with otherwise similar local neighborhoods. More recently, transformer-based GNNs have incorporated PSEs into attention mechanisms [14, 40, 31]. Other works focus on decoupling the positional information from the structure and learn the embeddings instead of picking them manually [14]. Building upon this, the Graph Positional and Structural Encoder (GPSE) [13] offers a unified and transferable solution by learning a shared latent space of PSEs through self-supervised pretraining. GPSE’s embeddings can be integrated into downstream GNNs, achieving performance on par with or exceeding task-specific encodings, and advancing the development of general-purpose graph models.
Combinatorial Complex Neural Networks.
To overcome the pairwise limitations of graphs, recent work has turned to explicit representation of higher-order interactions through topological domains. Combinatorial complexes provide a unified framework for modeling higher-order interactions with hierarchical organization, and their associated Combinatorial Complex Neural Networks (CCNNs) enable message passing across cells of varying rank (HOMP paradigm), naturally capturing multi-way relationships [23]. Generalizing and subsuming CCNNs, Generalized CCNNs (GCCNs) [39] offer a more flexible and powerful framework: they systematically decompose a complex into a set of strictly augmented Hasse graphs –each one capturing a particular neighborhood– in a modular way that preserves the higher-order structure. In doing so, GCCNs not only facilitate the design of new TNN architectures but also help bridge the gap between graph deep learning and TDL. However, existing TNNs –CCNNs and GCCNs alike– suffer from the high computational complexity associated to HOMP, limiting their practical use in large real-world applications.
Positional encodings in higher-order domains.
Encoding positional information for higher-order topological elements poses unique challenges. Unlike in graphs, where positional features are typically assigned to nodes, each cell rank of a combinatorial complex –nodes, edges, faces,…– potentially requires its own descriptor. Recent works have begun addressing this, though this topic still remains largely unexplored. For instance, the Cellular Transformer (CT) [1] extends transformer architectures to cell complexes by incorporating attention mechanisms that respect higher-order incidence relations. To support this, the authors introduce a set of manually crafted positional encodings tailored to cellular complexes, though their results did not shed light on how to build robust encodings—different ones perform best on different tasks. Moreover, CT incurs quadratic complexity in the number of cells due to the attention mechanism, limiting its scalability to medium-to-large scenarios. In contrast, Simplicial-aware Neural Networks (SaNN[20] take a more efficient route by bypassing message passing entirely. SaNN precomputes structural and positional summaries (e.g., over lower, upper, and co-neighbors) and feeds them as fixed-length vectors into a standard neural network. This achieves constant runtime and memory overhead while maintaining strong expressivity and performance. Nevertheless, the SaNN architecture is also constrained in terms of the topological domain (simplicial complex-specific), as well as with the choices of neighborhoods –which haves an impact into its theoretical expressivity and applicability.
4 HOPSE: higher-order positional and structural encoder
Section 4.1 outlines the proposed HOPSE pipeline, Section 4.2 establishes its theoretical guarantees, and Section 4.3 presents two practical realizations of the framework.
Let ${\mathcal{T}}$ be an arbitrary combinatorial complex, and ${\mathcal{N}}_{C}\subset{\mathcal{N}}_{\mathcal{T}}$ any subset of neighborhoods on ${\mathcal{T}}$. The following notation is introduced to support the upcoming presentation of the framework:
Strictly augmented multi-attributed Hasse graphs.
For a given neighborhood ${\mathcal{N}}\in{\mathcal{N}}_{\mathcal{C}}$, its strictly augmented multi-attributed Hasse graph –denoted by ${\mathcal{J}}_{{\mathcal{N}}}$– assigns to the resulting strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$ the corresponding signals $\mathcal{F}_{{\mathcal{N}}}$ over the cells present in that neighborhood, i.e. ${\mathcal{J}}_{{\mathcal{N}}}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},\mathcal{F}_{\mathcal{N}})$. The shorthand ${\mathcal{F}}_{{\mathcal{N}},k}=\{F_{r,k}\}_{r\geq 0}$ is the set of all $k$-attribute functions for each rank $r$ that act on the subset of cells in a neighborhood ${\mathcal{C}}_{\mathcal{N}}$ as $F_{r,k}:\{\sigma\in{\mathcal{C}}_{\mathcal{N}}\}\to\mathbb{R}^{D_{r}}$. To streamline the management of multiple associated signals, the restriction of ${\mathcal{J}}_{{\mathcal{N}}}$ to the $k$-th feature is defined as $\evaluated{{\mathcal{J}}_{{\mathcal{N}}}}_{k}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},{\mathcal{F}}_{{\mathcal{N}},k})$. Lastly, $\mathbf{J}_{C}=\left\{{\mathcal{J}}_{{\mathcal{N}}}\right\}_{{\mathcal{N}}\in{\mathcal{N}}_{C}}$ denotes the corresponding Hasse graph decomposition of ${\mathcal{T}}$ w.r.t. a given neighborhood set ${\mathcal{N}}_{C}$.
Rank-targeted neighborhoods.
The rank-targeted neighborhood is defined as the collection of neighborhood functions whose target cells have the same fixed rank $r$: $\mathcal{N}^{\to r}:=\{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}\mid\forall\sigma\in\mathcal{C},\forall\tau\in\mathcal{N}(\sigma):\mathrm{rk}(\tau)=r\}$.
4.1 Methodology
This section describes HOPSE, a method for enriching cell features in combinatorial complexes with higher-order positional and structural encodings derived from Hasse graph decompositions. HOPSE enables expressive higher-order modeling without relying on the computational overhead of HOMP, while retaining its representational power. The full pipeline is illustrated in Figure 2, with components referenced using the corresponding labels (A–F) throughout this section.
Figure 2: HOPSE pipeline. (A) Considering the collection of neighborhoods ${\mathcal{N}}_{C}=\{{\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{0\to 2}\}$, the input combinatorial complex $\mathcal{T}$ is decomposed into the corresponding multi-attributed Hasse graphs expansions $\{\mathcal{J}_{{\mathcal{N}}}\}_{{\mathcal{N}}_{C}}$ –each one with $k$ associated signals. (B) Positional and structural encodings are computed for each Hasse graph, $\mathbf{G}(\mathcal{J}_{\mathcal{N}_{i}})$, as defined in Equation 1. (C) Higher-order PS encodings are derived via the target rank-aware aggregation $\psi_{k}$ (see Equation 2). (D) The PS encodings $\mathbf{X}_{r,k}$ are transformed using $f_{r,k}(\cdot)$, and initial cell features $\mathbf{Z}_{r}$ are embedded via $\varepsilon_{r}(\cdot)$, following Equations 3 and 4. (E) Rank-specific, neighborhood-aware representations $\mathbf{H}_{r}$ are learned using $\Theta_{r}(\cdot)$, as described in Equation 5. (F) The final task-specific readout $\varphi(\cdot)$ integrates all $\mathbf{H}_{r}$ representations (Equation 6).
Extract positional and structural information.
Let ${\mathcal{N}}_{C}$ be a collection of neighborhood functions. For each neighborhood ${\mathcal{N}}\in{\mathcal{N}}_{C}$, the Hasse graph decomposition of ${\mathcal{T}}$ enables the treatment of every ${\mathcal{J}}_{{\mathcal{N}}}$ as an independent component (see Figure 2A), each representing a unique set of connectivity relationships. $\mathbf{G}$ is a set of Positional and Structural Encoders (PSEs) such that
$\mathbf{G}\colon\{1,\dots,K\}\times\mathbf{J}_{\mathcal{C}}\longrightarrow\mathbb{R}^{N\times D_{k}},\quad\left(k,{\mathcal{J}}_{{\mathcal{N}}}\right)\longmapsto g_{k}\left({{\mathcal{J}}_{{\mathcal{N}}}\bigr{|}}_{k}\right).$ (1)
Here $g_{k}(\cdot):\mathbf{J}_{\mathcal{C}}\to\mathbb{R}^{|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}\times D_{k}}$ denotes a particular PSE, with $|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}$ being the number of nodes in the Hasse graph induced by neighborhood ${\mathcal{N}}$ corresponding to the target cells (illustrated in Figure 2B). Note that each $g_{k}(\cdot)$ maps to a different $D_{k}$, allowing PSEs to have different dimensions.
Aggregate positional and structural encodings across neighborhoods.
The proposed neighborhood aggregation mechanism $\psi_{k}(\cdot)$ operates by grouping the encodings of cells with the same target rank $r$ to produce a unique embedding for each cell and each $k\in\{1,\dots,K\}$ (see Figure 2C). The embeddings for cells of rank $r$ are computed as:
$\mathbf{X}_{r,k}=\psi_{k}(\{g_{k}\left({{\mathcal{J}}_{{\mathcal{N}}}\bigr{|}}_{k}\right)\mid{\mathcal{N}}\in{\mathcal{N}}^{\to r}\}),$ (2)
where $\mathbf{X}_{r,k}\in\mathbb{R}^{N_{r}\times D_{k^{\prime}}}$, with $N_{r}$ denoting the number of $r$-cells and $D_{k^{\prime}}$ representing the output feature dimension for attribute $k$. The functions $\psi_{k}(\cdot)$ can also be interpreted as rank-level aggregators, where the aggregation operation can be any standard reduction method. Common choices include concatenation, sum, average, or even learnable transformations such as MLPs. The main advantage of using non-learnable aggregation functions for $\psi_{k}$ is that the embeddings $\mathbf{X}_{r,k}$ can be precomputed during dataset preprocessing, since they do not require learning.
Learning embeddings of cells.
In the HOPSE framework, each $r$-cell is associated with two complementary sources of information: (i) higher-order positional and structural encodings $\mathbf{X}_{r,k}$ aggregated from multiple neighborhoods, and (ii) initial features $\mathbf{Z}_{r}$ provided by the dataset or derived during the construction of the higher-order structure. Both types of features are independently projected into latent spaces through sets of learnable functions (see Figure 2D):
$\hat{\mathbf{X}}_{r,k}$ $=f_{r,k}(\mathbf{X}_{r,k}),$ (3)
$\hat{\mathbf{Z}}_{r}$ $=\varepsilon_{r}(\mathbf{Z}_{r}),$ (4)
where $f_{r,k}(\cdot)$ is applied to each feature $k$ of the $r$-cells, and $\varepsilon_{r}(\cdot)$ to their initial features.
Learning neighborhood-aware representations.
Next, to integrate the structural information with the cells’ initial attributes, the extracted and transformed higher-order PS encodings and the cell features are passed through the aggregator $\Theta_{r}(\cdot)$ (see Figure 2E) to obtain the final neighborhood-aware representations $\mathbf{H}_{r}\in\mathbb{R}^{N_{r}\times D_{r}}$:
$\mathbf{H}_{r}=\Theta_{r}(\hat{\mathbf{Z}}_{r},\hat{\mathbf{X}}_{r,1},\dots,\hat{\mathbf{X}}_{r,K}),$ (5)
Learning the downstream task.
Since the structural information is now contained in the cell features, they can be treated independently. For example, for complex-level tasks where the network needs an output for the whole complex, the global transformation $\varphi(\cdot)$ combines the aggregated embeddings from all ranks (see Figure 2F). The overall output is
$\text{HOPSE}({\mathcal{T}})=\varphi\Bigl{(}\mathbf{H}_{0},\mathbf{H}_{1},\ldots,\mathbf{H}_{R}\Bigr{)}.$ (6)
4.2 Theoretical guarantees
The expressivity of GNNs is commonly analyzed through their correspondence with the Weisfeiler–Leman (WL) isomorphism test [26]. In the context of TDL, a generalization of WL over arbitrary neighborhoods of a combinatorial complex (CC) has been proposed, referred to as the CCWL test [39]. The following theorem establishes the expressivity of HOPSE with respect to it.
Theorem 4.1.
HOPSE is more powerful than the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
The proof idea is that although HOPSE does not rely on message passing, with appropriately chosen neighborhoods, it can compute positional encodings using the same information as CCWL. Given the injectivity assumptions, this implies at least equal expressivity. The proof concludes by showing two CCs distinguished by HOPSE but not by CCWL. Full details are in Appendix C.
Another key property for models working on combinatorial data is cell permutation equivariance—the ability to yield consistent outputs under permutations of the cells. The following theorem establishes that HOPSE preserves this property under reasonable conditions:
Theorem 4.2.
HOPSE is cell permutation equivariant if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are cell permutation equivariant for all $k\in{1,\dots,K}$ and all $r\in{0,\dots,R}$.
This follows from the fact that compositions of permutation-equivariant functions remain equivariant. The proof is in Appendix C. While this requirement may seem strong, standard MLPs—though not equivariant to feature permutations—are equivariant to permutations of input sets, making them suitable functions for $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$.
4.3 Realization of the framework
The practical implementation of the proposed framework requires defining a set of neighborhoods of interest ${\mathcal{N}}_{C}=\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$ (discussed in Section 5), a set of encoding functions $\mathbf{G}$ for deriving structural and positional features, neighborhood aggregation functions $\psi_{k}(\cdot)$, rank-specific processing functions $f_{r,k}(\cdot)$, and a structure mixing function $\Theta_{r}(\cdot)$. In particular, this work adopts $\psi_{k}(\cdot)=\mathrm{CONCAT}(\cdot)$ for aggregation, enabling each positional encoding to be handled independently. For both $f_{r,k}(\cdot)$ and $\Theta_{r}(\cdot)$, trainable multilayer perceptrons (MLPs) are used. The choice of encoding functions $\mathbf{G}$ defines the two proposed instantiations of the framework, HOPSE-M and HOPSE-G, offering a trade-off regarding the informativeness of extracted PS encoding and the computation cost to generate them. Additional implementation details are provided in Appendix D.
HOPSE-M.
This realization incorporates manually selected positional and structural encodings commonly used in graph learning. Specifically, $\mathbf{G}=\{g_{1},g_{2},g_{3},g_{4}\}$ includes: (i) eigenvector-based encodings, (ii) electrostatic potential-based encodings, (iii) random-walk-based encodings, and (iv) heat kernel signatures. These encodings are chosen to provide a fair comparison with HOPSE-G, as the GPSE model used in that variant was trained on the same set of positional and structural descriptors [13]. While informative, several of these encodings—especially those requiring eigendecomposition such as (i), (ii), and (iv)—are computationally intensive, with complexity ${\mathcal{O}}(n^{3})$ with $n$ the number of graph nodes.
HOPSE-G.
This variant leverages the pretrained GPSE network [13] to encode structural and positional information through a lightweight message-passing architecture. In this realization, $\mathbf{G}=\{g_{1}\}$ with $g_{1}=\textsc{GPSE}(\cdot)$. Unlike manual encodings that rely on computationally intensive operations such as eigendecomposition, GPSE produces latent descriptors using standard message passing. This makes HOPSE-G particularly suitable for large combinatorial complexes with many cells, offering a scalable alternative without significant loss in expressiveness.
Computational complexity.
The presented implementation of HOPSE is designed for computational efficiency: no training is required up to and including Equation 2, allowing the matrices $\mathbf{X}_{r,k}$ to be fully precomputed in a preprocessing phase. By offloading all structure-dependent computations to preprocessing, HOPSE has complexity that scales linearly with the number of cells while retaining high expressive power, as established in Theorem 4.1. The efficiency of this design is further validated by experimental runtime results in Table 3. A comprehensive complexity analysis and a comparison with GCCN—a highly general higher-order architecture—are provided in Appendix E.
5 Experiments
This section empirically evaluates the proposed HOPSE framework, benchmarked against state-of-the-art TNNs. Both instantiations of the framework, HOPSE-M and HOPSE-G (see Section 4.3), are assessed on graph molecular datasets, as well as on expressivity and topological tasks. The standardized experimental protocol introduced in TopoBench [42] is adopted to ensure fair and reproducible comparisons through a comprehensive optimization and evaluation suite. The experimental setup is carefully designed to systematically address the following research questions:
  1. Q1.
    Can the proposed HOPSE framework and its realizations (HOPSE-M and HOPSE-G) effectively approximate the performance of TNNs that leverage HOMP on standard TDL datasets?
  2. Q2.
    Can HOPSE achieve competitive performance on synthetic datasets specifically constructed to evaluate the ability to capture and utilize topological properties?
  3. Q3.
    How expressive HOPSE can get in practical scenarios compared with other GNNs and TNNs?
  4. Q4.
    To what extent does the reduced complexity of HOPSE, relative to HOMP-based methods, yield practical improvements in execution cost and scalability on standard TDL benchmarks?
5.1 Evaluation setup
Datasets.
First, the framework is evaluated on five standard benchmarking datasets commonly used in the TDL literature [42], comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^3] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [35]. For regression, the ZINC dataset is used [27]. Following standard practice in TDL, structural lifting is applied to lift these graph datasets into higher-order domains—specifically, cycle-based lifting for the cell domain and clique complex lifting for the simplicial domain. Next, MANTRA family [2] represents purely topological datasets and tasks. It comprises over 43,000 triangulations of 2-manifolds. Among the multiple classification tasks that MANTRA datasets offer, this analysis focuses on the following three: predicting the homeomorphism name, predicting the orientability of a triangulated surface, and assigning a numerical label to the first three Betti numbers $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. Lastly, the evaluation includes the CLS dataset [36], used to assess the expressivity of GNNs [16], which contains 150 graphs across 10 isomorphism classes. For further details, see Appendix F.
Models.
The SaNN architecture is re-implemented due to unavailable source code [19] and included in all benchmarks. For the MANTRA family [2], baseline results are reported for the first time under the TopoBench protocol, ensuring consistency with the broader evaluation framework. Full details of the hyperparameter search and selected configurations are provided in Appendix F.2.
Training and evaluation.
To ensure fair comparisons, five stratified splits with a 50%/25%/25% train/validation/test ratio are considered for each dataset and task –except on ZINC, which uses predefined splits [27]. The optimal hyperparameter configuration is selected based on the average validation performance across five random seeds. Evaluation metrics include classification accuracy (MUTAG, PROTEINS, NCI1, NCI109), mean absolute error (MAE; ZINC), and F1 score (MANTRA). For each dataset, the mean and standard deviation are computed across the five test sets and reported in Table 1. Publicly available results are adopted from Telyatnikov et al. [42] and Papillon et al. [39] where applicable for graph datasets.
5.2 Experimental Results
Table 1: Accuracy on benchmark datasets. Results are shown as mean $\pm$ standard deviation. The best results are highlighted in bold, and those within one standard deviation of the best are in blue.
Graph Tasks Topological Tasks Model MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) Graph GCN 69.79 $\pm$ 6.80 75.70 $\pm$ 2.14 72.86 $\pm$ 0.69 72.20 $\pm$ 1.22 0.62 $\pm$ 0.01 42.14 $\pm$ 2.72 47.94 $\pm$ 0.00 46.86 $\pm$ 4.50 0.00 $\pm$ 0.00 GAT 72.77 $\pm$ 2.77 76.34 $\pm$ 1.66 75.00 $\pm$ 0.99 73.80 $\pm$ 0.73 0.61 $\pm$ 0.01 18.09 $\pm$ 0.65 47.94 $\pm$ 0.00 7.45 $\pm$ 0.05 0.00 $\pm$ 0.00 GIN 79.57 $\pm$ 6.13 75.20 $\pm$ 3.30 74.26 $\pm$ 0.96 74.42 $\pm$ 0.70 0.57 $\pm$ 0.04 76.14 $\pm$ 0.14 56.28 $\pm$ 0.45 88.13 $\pm$ 0.00 0.93 $\pm$ 1.21 Simplicial SCN 73.62 $\pm$ 6.13 75.27 $\pm$ 2.14 74.49 $\pm$ 1.03 75.70 $\pm$ 1.04 0.53 $\pm$ 0.04 79.48 $\pm$ 1.36 69.55 $\pm$ 0.97 76.45 $\pm$ 3.06 5.45 $\pm$ 2.31 SCCNN 76.17 $\pm$ 6.63 74.19 $\pm$ 2.86 76.60 $\pm$ 1.75 77.12 $\pm$ 1.07 0.36 $\pm$ 0.02 95.08 $\pm$ 0.56 86.29 $\pm$ 1.23 90.20 $\pm$ 0.20 65.82 $\pm$ 2.70 SaNN 79.57 $\pm$ 3.56 75.05 $\pm$ 3.52 68.75 $\pm$ 1.64 68.17 $\pm$ 0.88 0.55 $\pm$ 0.01 81.76 $\pm$ 1.37 61.65 $\pm$ 0.55 88.46 $\pm$ 0.09 39.22 $\pm$ 2.80 GCCN 85.96 $\pm$ 4.66 74.91 $\pm$ 2.51 76.85 $\pm$ 1.50 75.76 $\pm$ 1.28 0.35 $\pm$ 0.01 86.76 $\pm$ 1.27 76.60 $\pm$ 1.67 84.20 $\pm$ 4.80 41.82 $\pm$ 20.19 HOPSE-M 86.38 $\pm$ 6.31 75.13 $\pm$ 3.66 77.35 $\pm$ 0.81 76.34 $\pm$ 1.40 0.39 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G 85.53 $\pm$ 2.77 76.27 $\pm$ 2.10 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.43 $\pm$ 0.01 81.75 $\pm$ 1.26 62.17 $\pm$ 0.98 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25 Cell CCCN 77.02 $\pm$ 9.32 73.33 $\pm$ 2.30 76.67 $\pm$ 1.48 75.35 $\pm$ 1.50 0.34 $\pm$ 0.02 - - - - CWN 80.43 $\pm$ 1.78 76.13 $\pm$ 2.70 73.93 $\pm$ 1.87 73.80 $\pm$ 2.06 0.34 $\pm$ 0.01 - - - - GCCN 86.38 $\pm$ 6.49 74.41 $\pm$ 1.77 78.23 $\pm$ 1.47 77.19 $\pm$ 0.21 0.19 $\pm$ 0.00 - - - - HOPSE-M 88.09 $\pm$ 5.55 76.99 $\pm$ 2.63 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.25 $\pm$ 0.01 - - - - HOPSE-G 88.09 $\pm$ 1.90 76.42 $\pm$ 2.19 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01 - - - -
Accuracy comparison.
Table 1 presents a performance comparison across topological domains: graph, simplicial, and cell. The proposed HOPSE-M and HOPSE-G models demonstrate consistently strong performance, particularly in the simplicial domain, answering both Q1. and Q2. affirmatively. On standard graph classification benchmarks used in TDL, both HOPSE variants match or outperform most baselines, with HOPSE-M ranking among the top models and achieving state-of-the-art results in several cases (PROTEINS, NCI1, NCI109). While the performance gap on ZINC highlights the advantage of message-passing in fine-grained molecular regression tasks, HOPSE still remains competitive without relying on the HOMP procedure. The most significant gains are observed on the pure topological datasets. HOPSE-M delivers the strongest results on $\beta_{1}$ and $\beta_{2}$, outperforming all baselines, including models with explicit simplicial message passing. This suggests that HOPSE can effectively encode and leverage topological structure. In summary, the HOPSE framework achieves strong performance across both graph and synthetic topological benchmarks, demonstrating that a carefully designed structural encoding pipeline can approximate or surpass topological message-passing methods. Detailed analysis of the results is provided in Appendix F.
Empirical expressivity.
Table 2: Accuracy on CSL ($\uparrow$).
Model Mean Max Min MP-GNNs[9] 10.00 $\pm$ 0.00 10.00 10.00 SaNN 26.00 $\pm$ 6.83 33.33 20.00 SCCNN 25.33 $\pm$ 7.67 33.33 16.67 GCCN 52.78 $\pm$ 34.28 100.00 3.33 3WLGNN [32] 97.80 $\pm$ 10.92 100.00 30.00 HOPSE-M 100.00 $\pm$ 0.00 100.00 100.00 HOPSE-G 98.00 $\pm$ 4.47 100.00 90.00
Table 2 reports the mean, minimum, and maximum test classification accuracy over five folds for the best-performing baselines, covering all domains. The results show that SaNN and SCCNN slightly outperform baseline MP-GNNs, while GCCN performs better on the cell domain but still falls short of 3WLGNN. HOPSE-M perfectly distinguishes all isomorphism classes, and HOPSE-G achieves the second-best position ( Q3.).
Computation time comparison.
To highlight the efficiency of HOPSE, Table 3 reports the total runtime in seconds. Notably, either HOPSE-G or HOPSE-M achieves the fastest overall runtime across both simplicial and cell domains. On graph datasets, the speedup can reach up to 7x—for example, HOPSE-G runs in 130 seconds compared to 904 seconds for GCCN on NCI1, while also delivering better performance. On average, HOPSE-M and HOPSE-G outperform the most competitive baselines with speedups of 1.5x and 3x over SCCN, and 3x and 5x over GCCN, respectively, while achieving top results on 6 out of 9 datasets. Overall, both models significantly reduce runtime while delivering competitive or superior accuracy (Q4.).
Table 3: Runtime in seconds. Results are shown as mean $\pm$ standard deviation. The fastest model per domain is in bold and those that are within one standard deviation of the best are in blue.
Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ($\uparrow$) ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 4 $\pm$ 1 8 $\pm$ 2 53 $\pm$ 20 37 $\pm$ 9 147 $\pm$ 28 3538 $\pm$ 38 3534 $\pm$ 49 7250 $\pm$ 234 GAT 4 $\pm$ 1 8 $\pm$ 2 57 $\pm$ 17 56 $\pm$ 9 171 $\pm$ 65 3589 $\pm$ 84 3544 $\pm$ 44 3612 $\pm$ 52 GIN 5 $\pm$ 1 9 $\pm$ 2 61 $\pm$ 24 50 $\pm$ 8 168 $\pm$ 107 3568 $\pm$ 44 3536 $\pm$ 46 3674 $\pm$ 176 Simplicial SCN 8 $\pm$ 2 43 $\pm$ 13 276 $\pm$ 63 226 $\pm$ 66 1210 $\pm$ 572 5115 $\pm$ 653 4773 $\pm$ 193 6992 $\pm$ 2049 SCCNN 14 $\pm$ 3 54 $\pm$ 11 307 $\pm$ 83 354 $\pm$ 106 2060 $\pm$ 408 5521 $\pm$ 319 6920 $\pm$ 1524 6720 $\pm$ 210 SaNN 9 $\pm$ 2 50 $\pm$ 12 250 $\pm$ 18 211 $\pm$ 66 1263 $\pm$ 252 5572 $\pm$ 1053 4619 $\pm$ 103 5220 $\pm$ 421 GCCN 61 $\pm$ 7 66 $\pm$ 21 904 $\pm$ 180 538 $\pm$ 39 3603 $\pm$ 475 5898 $\pm$ 1961 4564 $\pm$ 631 7689 $\pm$ 2522 HOPSE-M 7 $\pm$ 2 38 $\pm$ 8 164 $\pm$ 79 219 $\pm$ 26 1061 $\pm$ 405 3596 $\pm$ 519 3409 $\pm$ 162 5377 $\pm$ 185 HOPSE-G 5 $\pm$ 0 33 $\pm$ 8 130 $\pm$ 22 107 $\pm$ 28 478 $\pm$ 107 3324 $\pm$ 1113 2795 $\pm$ 112 3876 $\pm$ 1141 Cell CCCN 12 $\pm$ 2 42 $\pm$ 7 372 $\pm$ 109 272 $\pm$ 21 1621 $\pm$ 142 - - - CWN 11 $\pm$ 1 54 $\pm$ 18 302 $\pm$ 63 295 $\pm$ 46 1390 $\pm$ 97 - - - GCCN 61 $\pm$ 18 59 $\pm$ 18 523 $\pm$ 119 386 $\pm$ 76 3301 $\pm$ 440 - - - HOPSE-M 7 $\pm$ 1 37 $\pm$ 7 147 $\pm$ 24 122 $\pm$ 18 1313 $\pm$ 132 - - - HOPSE-G 7 $\pm$ 1 32 $\pm$ 5 153 $\pm$ 8 104 $\pm$ 38 1230 $\pm$ 527 - - -
Choice of ${\mathcal{N}}_{C}$.
The neighborhood set ${\mathcal{N}}_{C}$ is a critical design choice in the HOPSE framework, influencing both the expressivity and computational behavior of the model. Performed ablation studies (Appendix F) reveal that richer (i.e. larger and diverse) neighborhood configurations consistently yield top or near-top accuracy across complex datasets—e.g. MANTRA’s topological tasks, ZINC—as shown in Tables 6 and 7. Importantly, these gains come without significant runtime penalties: as demonstrated in Tables 10 and 11, per-epoch runtime remains almost stable, and total runtime often decreases due to faster convergence. Notably, the cell domain strongly aligns with expressivity and efficiency, especially on molecular tasks. These findings suggest a practical guideline: practitioners aiming for strong performance without exhaustive search over all neighborhood configurations can safely opt for larger ${\mathcal{N}}_{C}$ sets. This choice is especially advantageous for complex tasks, balancing predictive power, convergence speed, and runtime efficiency (Q1., Q2., Q4.).
6 Concluding remarks
Conclusions.
This work introduces HOPSE, a scalable framework designed to extract and leverage higher-order structural and positional information for arbitrary combinatorial representations. Theoretically, HOPSE matches the expressivity of state-of-the-art TNNs while avoiding the HOMP paradigm, thereby substantially reducing the computational cost of training and inference. Moreover, model scaling is linear with dataset size. These claims are supported by experiments on molecular graph benchmarks, as well as expressivity and topological tasks, where both instantiations—HOPSE-M and HOPSE-G—match or surpass state-of-the-art performance while achieving up to 7×7$\times$7 × speed-ups over HOMP-based models. By offering a compelling trade-off between scalability and expressivity, this work not only enables broader applicability of TDL but also facilitates the integration of higher-order interactions into other areas of deep learning—mirroring the widespread adoption of graph-based methods [47, 44, 24, 25].
Limitations & future work.
A deeper investigation of HOPSE encodings—along with a comprehensive comparison against alternative methods across diverse domains—lies beyond the scope of this work but remains an important and intriguing direction for future research. Other promising avenues include exploring graph-based strategies for generating higher-order positional and structural encodings, or combining encodings derived from multiple topological domains for enhanced expressivity. Additionally, this work does not address whether integrating the proposed encodings into GNNs, TNNs, or transformer-based architectures leads to substantial improvements across a range of downstream tasks. Finally, and from a more practical standpoint, HOPSE-G realization of the framework leverages a pretrained model to extract latent structural and positional features, but it would be interesting to explore if a joint end-to-end training may further enhance performance –even at the cost of additional computational overhead.
Appendix A Extended background
A.1 Topological domains
This section draws primarily from Appendix A of Telyatnikov et al. [42]. It begins with the fundamental concept of featured graphs, which forms the basis for understanding more intricate structures. Then it proceeds to explore higher-order domains — including simplicial complexes, cell complexes, and combinatorial complexes — each providing distinct capabilities for representing various types of relationships and hierarchies within data.
Definition A.1.
Let $\mathcal{G}=(V,E)$ be a graph, with node set $V$ and edge set $E$. A featured graph is a tuple $\mathcal{G}_{F}=(V,E,F_{V},F_{E})$, where $F_{V}:V\to\mathbb{R}^{d_{v}}$ is a function that maps each node to a feature vector in $\mathbb{R}^{d_{v}}$ and $F_{E}:E\to\mathbb{R}^{d_{e}}$ is a function that maps each edge to a feature vector in $\mathbb{R}^{d_{e}}$.
A topological domain is a graph generalization that captures pairwise and higher-order relationships between entities [7, 4]. When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions [23, 38].
Definition A.2 (Set-type relation).
A relation in a topological domain is called a set-type relation if another relation in the domain does not imply its existence.
Definition A.3 (Rank function).
A rank function on a higher-order domain $\mathcal{X}$ is an order-preserving function $rk\colon\mathcal{X}\to\mathbb{Z}_{\geq 0}$ such that $x\subseteq y$ implies $rk(x)\leq rk(y)$ for all $x,y\in\mathcal{X}$.
Set-type relations emphasize the independence of connections within a domain, allowing for flexible representation of complex interactions. In contrast, rank functions introduce a hierarchical (also called part-whole) organization that facilitates the representation and analysis of nested relationships.
Simplicial complexes.
Simplicial complexes extend graphs by incorporating hierarchical part-whole relationships through the multi-scale construction of cells. In this structure, nodes correspond to rank 00-cells, which can be combined to form edges (rank $1$-cells). Edges can then be grouped to form faces (rank 2 cells), and faces can be combined to create volumes (rank $3$-cells), continuing in this manner. Consequently, the faces of a simplicial complex are triangles, volumes are tetrahedra, and higher-dimensional cells follow the same pattern. A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of $(k-1)$-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
Definition A.4 (Simplicial Complex).
A simplicial complex (SC) in a non-empty set $S$ is a pair $SC=(S,\mathcal{X})$, where $\mathcal{X}\subset\mathcal{P}(S)\setminus\{\emptyset\}$ satisfies: if $x\in SC$ and $y\subseteq x$, then $y\in SC$. The elements of $\mathcal{X}$ are called simplices.
Example 1 (3D Surface Meshes).
3D models of objects, such as those used in computer graphics or for representing anatomical structures, are often constructed using triangular meshes. These meshes naturally form simplicial complexes, where the vertices of the triangles are 0-simplices, the edges are 1-simplices, and the triangular faces themselves are 2-simplices.
Cell complexes
Cell complexes provide a hierarchical interior-to-boundary structure, offering clear topological and geometric interpretations, but they are not based on set-type relations. Unlike simplicial complexes, cell complexes are not limited to simplexes; faces can involve more than three nodes, allowing for a more flexible representation. This increased flexibility grants cell complexes greater expressivity compared to simplicial complexes [10, 8].
Definition A.5 (Cell complex).
A regular cell complex is a topological space $S$ partitioned into subspaces (cells) $\{x_{\alpha}\}_{\alpha\in P_{S}}$, where $P_{S}$ is an index set, satisfying:
  1. 1.
    $S=\cup_{\alpha\in P_{S}}\text{int}(x_{\alpha})$, where $\text{int}(x)$ denotes the interior of cell $x$.
  2. 2.
    For each $\alpha\in P_{S}$, there exists a homeomorphism $\psi_{\alpha}$ (attaching map) from $x_{\alpha}$ to $\mathbb{R}^{n_{\alpha}}$ for some $n_{\alpha}\in\mathbb{N}$. The integer $n_{\alpha}$ is the dimension of cell $x_{\alpha}$.
  3. 3.
    For each cell $x_{\alpha}$, the boundary $\partial x_{\alpha}$ is a union of finitely many cells of strictly lower dimension.
Example 2 (Molecular structures.).
Molecules admit natural representations as cell complexes by considering atoms as nodes (i.e., cells of rank zero), bonds as edges (i.e., cells of rank one), and rings as faces (i.e., cells of rank two).
Combinatorial complexes.
Combinatorial complexes combine hierarchical structure with set-type relations, enabling a flexible yet comprehensive representation of higher-order networks.
Definition A.6 (Combinatorial complex).
A combinatorial complex (CC) is a triple $(\mathcal{V},\mathcal{C},{\textnormal{k}})$ consisting of a set $\mathcal{V}$, a subset $\mathcal{C}\subset\mathcal{P}(\mathcal{V})\setminus\{\emptyset\}$, and a function $rk\colon\mathcal{C}\to\mathbb{Z}_{\geq 0}$ satisfying:
  1. 1.
    For all $v\in\mathcal{V}$, $\{v\}\in\mathcal{C}$ and $rk(v)=0$.
  2. 2.
    The function $rk$ is order-preserving: if $x,y\in\mathcal{C}$ with $x\subseteq y$, then $rk(x)\leq rk(y)$.
Example 3 (Geospatial structures.).
Geospatial data, comprised of grid points (0-cells), road polylines (1-cells), and census tract polygons (2-cells), can be effectively represented using combinatorial complexes. A visual example is provided in Figure 2 (Right) of Battiloro et al. [3].
Featured topological domains.
A featured graph is a graph whose nodes or edges are equipped with feature functions [41]. Similarly, each topological element (e.g., simplex or cell) can carry feature vectors to represent some information.
Definition A.7 (Featured topological domain).
A featured topological domain is a pair $(\mathcal{X},F)$, where $\mathcal{X}$ is a topological domain and $F=\{F_{i}\}_{i\geq 0}$ is a collection of feature functions. Each function $F_{i}$ maps the $i$-dimensional elements of $\mathcal{X}$, denoted $\mathcal{X}_{i}$, to a feature space $\mathbb{R}^{k_{i}}$:
$F_{i}\colon\mathcal{X}_{i}\to\mathbb{R}^{k_{i}}.$
A.2 Neighborhood functions and augmented Hasse graphs
Discrete topological domains are equipped with a notion of neighborhood among cells that confers on them a topological structure; this section revisits the main neighborhood-related concepts in the general combinatorial complex case.
Definition A.8 (Neighborhood function).
Given a combinatorial complex ${\mathcal{T}}=(\mathcal{V},\mathcal{C},{\textnormal{k}})$, a neighborhood $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a combinatorial complex is a function that assigns to each cell $\sigma\in\mathcal{C}$ a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$.
Each neighborhood can be used to create a Strictly Augmented Hasse graph.
Definition A.9 (Strictly augmented Hasse graph).
Each particular neighborhood $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ on $\mathcal{T}$ [39, 23], defined as the directed graph whose nodes and edges are given, respectively, by $\mathcal{C}_{\mathcal{N}}=\{\sigma\in\mathcal{C}\,|\,\mathcal{N}(\sigma)\neq\emptyset\}$ and $\mathcal{E}_{\mathcal{N}}=\{(\tau,\sigma)\,|\,\tau\in\mathcal{N}(\sigma)\}$.
Examples of neighborhood functions are incidences (connecting cells with different ranks) and adjacencies (connecting cells with the same rank), although other neighborhoods can be defined for specific tasks [3]. More in detail, given a combinatorial complex ${\mathcal{T}}$:
Definition A.10.
(Incidences) The incidence neighborhood of a $t$-cell $\sigma\in\mathcal{T}$ with respect to a rank $s$ is defined as
$\begin{split}\mathcal{I}_{s\to t}(\sigma)=\left\{\begin{array}[]{ll}\{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=s,\sigma\subset\tau\}&\text{ if $t<s$;}\\ \{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=s,\tau\subset\sigma\}&\text{ if $t>s$.}\end{array}\right.\end{split}$ (7)
A $s$-cell $\tau\in\mathcal{T}$ is said to be co-incident of $\sigma$ if it contains $\sigma$ and $s>t$; analogously, a $s$-cell $\tau\in\mathcal{T}$ is incident of $\sigma$ if it is is contained by $\sigma$ and $s<t$.
Incidences in turn induce adjacency neighborhoods of cells of the same rank as follows:
Definition A.11 (Adjacencies).
The adjacency neighborhood of a $t$-cell $\sigma\in\mathcal{T}$ with respect to a rank $s$ is defined as
$\begin{split}\mathcal{A}_{s,t}(\sigma)=\{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=t,\mathcal{I}_{s\to t}(\sigma)\cap\mathcal{I}_{s\to t}(\tau)\neq\emptyset\}\end{split}$ (8)
In this case, two $t$-cells $\sigma$ and $\tau$ are said to be adjacent if they are both contained in a $s$-cell $\delta\in\mathcal{T}$.
Appendix B Counting neighborhoods and message passing routes
(b) ${\mathcal{I}}_{1\to 0}$
(c) ${\mathcal{I}}_{2\to 0}$
(d) ${\mathcal{A}}_{0,1}$
(e) ${\mathcal{A}}_{0,2}$
(f) ${\mathcal{I}}_{0\to 1}$
(g) ${\mathcal{I}}_{2\to 1}$
(h) ${\mathcal{A}}_{1,0}$
(i) ${\mathcal{A}}_{1,2}$
(j) ${\mathcal{I}}_{0\to 2}$
(k) ${\mathcal{I}}_{1\to 2}$
(l) ${\mathcal{A}}_{2,0}$
(m) ${\mathcal{A}}_{2,1}$
Figure 3: Illustration of a combinatorial complex (top) and different neighborhood types, grouped by the target entity: nodes (second row), edges (third row), and faces (bottom row).
This section provides a detailed discussion of two key concepts in Topological Deep Learning: expanded neighborhoods and the combinatorial explosion of message passing routes. This discussion, while emphasizing the expressive power of combinatorial representations, highlights the challenges they pose when scaling to real-world problems.
B.1 Neighborhood Types in Combinatorial Complexes
Combinatorial representations allow for a variety of neighborhood types, depending on how cells of different ranks interact. For a 2-dimensional combinatorial complex (with 00-, $1$-, and $2$-cells), Figure 3 illustrates all possible neighborhood types. These relations can be grouped into two main groups:
  • Incidence Relations:
    • (b) ${\mathcal{I}}_{1\to 0}$: edges $\rightarrow$ nodes
    • (c) ${\mathcal{I}}_{2\to 0}$: faces $\rightarrow$ nodes
    • (f) ${\mathcal{I}}_{0\to 1}$: nodes $\rightarrow$ edges
    • (g) ${\mathcal{I}}_{2\to 1}$: faces $\rightarrow$ edges
    • (j) ${\mathcal{I}}_{0\to 2}$: nodes $\rightarrow$ faces
    • (k) ${\mathcal{I}}_{1\to 2}$: edges $\rightarrow$ faces
  • Adjacency Relations:
    • (d) ${\mathcal{A}}_{0,1}$: nodes $\rightarrow$ nodes (via edges)
    • (e) ${\mathcal{A}}_{0,2}$: nodes $\rightarrow$ nodes (via faces)
    • (h) ${\mathcal{A}}_{1,0}$: edges $\rightarrow$ edges (via nodes)
    • (i) ${\mathcal{A}}_{1,2}$: edges $\rightarrow$ edges (via faces)
    • (l) ${\mathcal{A}}_{2,0}$: faces $\rightarrow$ faces (via nodes)
    • (m) ${\mathcal{A}}_{2,1}$: faces $\rightarrow$ faces (via edges)
B.2 Counting neighborhoods
The combinatorial complex depicted in Figure 3 has rank $R=\dim({\mathcal{T}})=2$, which results in 12 distinct neighborhood types. These neighborhoods can be divided into the two categories from above: (i) incidence relations and (ii) adjacency relations.
Under these assumptions, the number of incidence relations is $|I|=(R+1)R$, as each of the $R+1$ ranks can be connected to any of the remaining $R$ ranks. Similarly, each rank gives rise to $R$ possible adjacency relations (each defined through a different intermediary rank), also resulting in $|A|=(R+1)R$. Combining these yields a total of
$|{\mathcal{N}}_{{\mathcal{T}}}|=2(R+1)R$
distinct neighborhood types. For the example in Figure 3, where $R=2$, this formula recovers the 12 neighborhoods observed.
B.3 Message passing routes
The number of neighborhoods increases quadratically with the maximum rank, which may initially seem manageable, especially since most practical datasets involve low-rank structures (e.g., $R=2$ or $R=3$). However, the true combinatorial complexity arises when considering how these neighborhoods can be composed. To formalize this notion, the following definition is introduced.
Definition B.1 (Message passing route).
A message passing route is an ordered set of neighborhoods that a model uses to propagate information across a data structure.
Information may flow intra-rank, through adjacency relations, or inter-rank, through incidence relations. For example, a network might use the message passing route $[{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ to transmit information from faces to edges, then from edges to nodes, from nodes back to faces, and finally between nodes connected by an edge.
Given a fixed combinatorial complex, the number of possible message passing routes is infinite, since any sequence can reuse the same neighborhood relations and grow to arbitrary length. In what follows, the analysis is restricted to a subset of message passing routes that are particularly relevant for the design of higher-order neural architectures.
Definition B.2 (Minimal message passing route).
A minimal message passing route is a message passing route of minimal length that updates all ranks in a higher-order representation at least once. Furthermore, it ensures that, after a sufficient number of message passing steps, the representation of each rank depends on all other ranks.
Minimal routes are of particular interest, as they guarantee full interdependence between all cells—a desirable property for expressive models. Moreover, minimal length implies simplicity in design, yielding the most compact routes that preserve this full dependency. For instance, the previously introduced route $[{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ can be transformed into a minimal message passing route by removing the adjacency ${\mathcal{A}}_{0,1}$. In this reduced route, information from all cells becomes accessible to all others within two message passing steps. In fact, consider the case of 00-cells: after the first step, they depend on $1$-cells; after the second step, since $1$-cells now incorporate information from $2$-cells, the 00-cells become indirectly connected to them as well.
By definition, minimal message passing routes rely exclusively on incidence relations, as adjacencies do not propagate information inter-rank. Since each incidence targets a specific rank, and every rank must be updated once, minimal routes necessarily have length $R+1$. To enumerate all possible minimal routes, consider the set of ranks $T=\{0,1,\dots,R\}$, of size $R+1$. There are $(R+1)!$ distinct permutations of these ranks, each defining a unique order in which the ranks are updated. Given such a permutation $\{r_{0},r_{1},\dots,r_{R}\}$, a corresponding minimal message passing route can be constructed by selecting the sequence of incidence relations ${\mathcal{I}}_{r_{j-1}\to r_{j}}$ for $j=1,\dots,R$, and closing the cycle with ${\mathcal{I}}_{r_{R}\to r_{0}}$. This construction yields exactly one valid minimal route per permutation, resulting in a total of $N=(R+1)!$ such routes. In the case of graphs, where $R=1$, this results in exactly two minimal message passing routes—an intuitively expected outcome.
B.4 Beyond minimal routes: including adjacencies
Minimal message passing routes represent only a small subset of the possible routes that can be constructed using the full set of neighborhoods in a combinatorial complex. A natural extension involves augmenting these minimal routes with adjacency operations—a common practice in many TDL architectures. For simplicity, the analysis here is restricted to the case where adjacencies are applied strictly after all incidence relations.
As previously discussed, for a given rank $r$, there are $R$ possible adjacency relations of the form ${\mathcal{A}}_{r,x}$. Since adjacencies operate intra-rank, they do not affect inter-rank dependencies and merely update the representation of the corresponding rank. As a consequence, the order in which these adjacencies are applied does not impact the outcome of the message passing route. For example, applying ${\mathcal{A}}_{0,0}$ followed by ${\mathcal{A}}_{1,1}$ produces the same result as applying them in reverse order.
Given that each of the $R+1$ ranks can be updated with any of their $R$ possible adjacencies, there are $(R+1)^{R}$ combinations of adjacency updates. When combined with the $(R+1)!$ minimal incidence-based routes, the total number of extended message passing routes becomes $\hat{N}=(R+1)!\cdot(R+1)^{R}$.
Appendix C Proofs
Before proving Theorem 4.1, the following paragraphs show that HOPSE is at least as powerful as the WL test. Although this is a weaker result than the one from Theorem 4.1, the proof is more intuitive, making it a valuable tool to understand the HOPSE model.
First, a beneficial result from Gurugubelli and Chepuri [20] regarding the WL test is reported.
Lemma C.1.
For a graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, the update of WL coloring of a node $\sigma\in{\mathcal{V}}$ at iteration K can be represented by the following equation
$c_{\sigma}^{(K)}=\phi\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{U}}^{(1)}({\sigma}),...,{\mathcal{U}}^{(K)}({\sigma})\}\}\mskip-5.0mu\}\right),$ (9)
where $\phi(\cdot)$ is some injective function and ${\mathcal{U}}^{(k)}(\sigma)$ denotes the set of upper adjacent $t$-hop neighbors of node $\sigma$.
Using this lemma, the following theorem can now be tackled.
Theorem C.2.
HOPSE is at least as powerful as the WL test in distinguishing non-isomorphic graphs if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
Proof.
Let ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ be two graphs. By considering their nodes as zero rank cells and their edges as rank-1 cells, they can be considered as CC without loss of generalization. Given two nodes (zero rank cells) $\sigma_{1}\in{\mathcal{T}}_{1}$ and $\sigma_{2}\in{\mathcal{T}}_{2}$ let $c_{\sigma_{1}}^{(K)}$, $c_{\sigma_{2}}^{(K)}$ be the WL coloring of the two nodes after $K$ steps of the WL algorithm, and $h_{\sigma_{1}}^{(K)}$ and $h_{\sigma_{2}}^{(K)}$ be the embeddings obtained with HOPSE using a set of neighborhoods $\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$. Showing that for all graphs ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ and for all nodes $\sigma_{1}$ and $\sigma_{2}$, there exists a version of HOPSE such that $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies $c_{\sigma_{1}}^{(K)}=c_{\sigma_{2}}^{(K)}$ proves that HOPSE is at least as powerful as the WL isomorphism test. This is because it would mean that the WL test fails to assign different embeddings whenever HOPSE fails.
As the focus is solely on the embeddings of the nodes in the considered graphs, the target rank is fixed to $r_{k}=0$, which is omitted in the remainder of the proof for clarity. Furthermore, since only neighborhoods ${\mathcal{N}}_{k}$ with target rank 0 are relevant, the proof is restricted to the case ${\mathcal{N}}_{k}={\mathcal{U}}^{(k)}$ for $k\in{1,\dots,K}$, where ${\mathcal{U}}^{(k)}$ denotes the upper adjacent $k$-hop neighborhood.
Without loss of generality, the analysis can be restricted to strictly augmented multi-attributed Hasse graphs defined inSection 4 of the form ${\mathcal{J}}_{{\mathcal{N}}}=({\mathcal{C}}_{{\mathcal{N}}},{\mathcal{E}}_{{\mathcal{N}}},{\mathcal{F}}_{{\mathcal{N}}})$, where ${\mathcal{F}}_{{\mathcal{N}}}={\mathcal{F}}$ for all ${\mathcal{N}}\in N^{\rightarrow 0}$, and ${\mathcal{F}}$ denotes the initial node features of the graph. As only a single feature type is considered, the index $k=0$ is omitted in the remainder of the proof for clarity.
For any node $\sigma$, by combining equation 4 to equation 5 the HOPSE embeddings can be written as
$h_{\sigma}^{(K)}=\Theta\Big{(}\varepsilon(\mathbf{Z}),\{f\big{(}\psi(\{g(\sigma,{\mathcal{J}}_{{\mathcal{N}}_{j}})|{\mathcal{N}}_{j}\in N^{\rightarrow 0}\})\big{)}\}\Big{)}.$ (10)
Since $g({\mathcal{J}}_{{\mathcal{N}}})$ is the function that creates the encodings for all target cells in the graph by creating a row for each cell, $g(\sigma,{\mathcal{J}}_{{\mathcal{N}}})$ is defined as the row-wise version of $g({\mathcal{J}}_{{\mathcal{N}}})$.
The functions $g(\sigma,{\mathcal{J}}_{{\mathcal{N}}})$ can be any injective functions, but by considering only functions that depend on the 1-hop neighbors of $\sigma$ in the corresponding Hasse graph, it is immediate to see that $g(\sigma,{\mathcal{J}}_{{\mathcal{U}}_{j}})=\xi(\{\mskip-5.0mu\{{\mathcal{F}}[\tau]|\forall\tau\in{\mathcal{U}}^{(j)}(\sigma)\}\mskip-5.0mu\})$, given that ${\mathcal{N}}_{j}={\mathcal{U}}^{(j)}$. Since by hypothesis $g(\cdot)$ is injective, also $\xi(\cdot)$ is. Since $\varepsilon(\cdot)$ is also injective by hypothesis equation 10 can be rewritten as
$\begin{split}h_{\sigma}^{(K)}&=\hat{\Theta}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in{\mathcal{U}}^{(1)}(\sigma)\}\mskip-5.0mu\},...,\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in{\mathcal{U}}^{(K)}(\sigma)\}\mskip-5.0mu\}\right)=\\ &=\hat{\Theta}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{U}}^{(1)}(\sigma),...,{\mathcal{U}}^{(K)}(\sigma)\}\}\mskip-5.0mu\}\right),\end{split}$ (11)
where the fact that $c_{\sigma}^{(0)}=\mathbf{Z}$ was used, which means that for the WL coloring algorithm the nodes’ features are the initial features. Also, since the functions $\Theta(\cdot)$, $f(\cdot)$, $\psi(\cdot)$, and $\xi(\cdot)$ are injective, $\hat{\Theta}(\cdot)$ is also injective since it is a composition of injective functions.
Since $\hat{\Theta}$ is injective $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies that the arguments of $\hat{\Theta}$ are equal. But those arguments are the exact arguments for the WL colorings of nodes $\sigma_{1}$ and $\sigma_{2}$ reported in equation 9, which means that $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies $c_{\sigma_{1}}^{(K)}=c_{\sigma_{2}}^{(K)}$. As discussed previously, this implies that HOPSE is at least as powerful as the WL test. ∎
The definition of the CCWL test, which generalizes the WL test to combinatorial complexes, is taken from [39].
Definition C.3 (The CC Weisfeiler-Leman (CCWL) test on labeled combinatorial complexes).
Let $({\mathcal{T}},F)$ be a labelled combinatorial complex. Let ${\mathcal{N}}$ be a neighborhood on ${\mathcal{C}}$. The scheme proceeds as follows:
  • Initialization: Cells $\sigma$ are initialized with the labels given by F, i.e.: for all $\sigma\in{\mathcal{C}}$: $c_{\sigma,F}^{(0)}=F(\sigma)$.
  • Refinement: Given colors of cells at iteration $k$, the refinement step computes the color of cell $\sigma$ at the next iteration $c_{\sigma,F}^{(k+1)}$ using a perfect HASH function as follows:
    $\begin{split}c_{\mathcal{N}}^{(k)}(\sigma)&=\{\mskip-5.0mu\{c_{\tau,F}^{(k)}|\forall\tau\in{\mathcal{N}}(\sigma)\}\mskip-5.0mu\},\\ c_{\sigma,F}^{(k+1)}&=\text{HASH}\left(c_{\sigma,F}^{(k)},c_{\mathcal{N}}^{(k)}(\sigma)\right).\end{split}$ (12)
  • Termination: The algorithm stops when an iteration leaves the coloring unchanged.
To show that HOPSE is more powerful than the CCWL test, the following lemma is proven first.
Lemma C.4.
For a combinatorial complex $({\mathcal{T}},F)$ the update of CCWL coloring of a cell $\sigma\in{\mathcal{C}}$ at iteration $T$ can be represented as the following equation
$c_{\sigma}^{(K)}=\phi\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}({\sigma}),...,{\mathcal{N}}^{(K)}({\sigma})\}\}\mskip-5.0mu\}\right),$ (13)
where $\phi(\cdot)$ is some injective function, and ${\mathcal{N}}^{(k)}$ denotes the $t$-hops version of the neighborhood ${\mathcal{N}}$.
Proof.
The proof is by induction. For $k=0$, the equation holds trivially, since the initial coloring for cell $\sigma$ is $c_{\sigma}^{(0)}$. Assuming that the $k-1$ update is true
$c_{\sigma}^{(t-1)}=\phi_{0}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\right).$
By substituting this in equation 12 the $k$ update can be written as
$\begin{split}c_{\sigma}^{(k)}=\text{HASH}\Big{(}&\phi_{0}\big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\\ \{\mskip-5.0mu\{&\phi_{0}\big{(}c_{\upsilon}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\forall\upsilon\in{\mathcal{N}}^{(1)}(\sigma)\}\mskip-5.0mu\}\Big{)}.\end{split}$
Now, since HASH and $\phi_{0}$ are injective functions, it follows that
$\begin{split}c_{\sigma}^{(k)}&=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\},\\ &\hskip 14.22636pt\{\mskip-5.0mu\{\big{(}c_{\upsilon}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau 1\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\forall\upsilon\in{\mathcal{N}}^{(1)}(\sigma)\}\mskip-5.0mu\}\Big{)}\\ &=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(k)}(\sigma)\}\}\mskip-5.0mu\}\Big{)}\\ &=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(k)}(\sigma)\}\}\mskip-5.0mu\}\Big{)}\end{split}$
This is the result needed, so the theorem is proven by induction. ∎
This result is the combinatorial complex equivalent of Lemma C.1, and it can now be used to prove the following theorem.
Theorem C.5.
HOPSE is at least as powerful as the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $i\in{0,...,R}$.
Proof.
Given that Lemma C.4 is equivalent to Lemma C.1 with ${\mathcal{N}}_{i}={\mathcal{N}}^{(i)}$ for $i\in 0,...,K$, this proof follows the same steps as the proof for Theorem C.2. ∎
Figure 4: a. Two combinatorial complexes with nodes, edges, and faces. ${\mathcal{T}}_{1}$ is a icosahedron while ${\mathcal{T}}_{2}$ is composed of five tetrahedrons. b. Their respective Hasse graphs given the choice of the neighborhood ${\mathcal{A}}_{2,1}$, which connects faces that share an edge. Image adapted from [39].
Now the two main theorems of the paper can be proven.
Theorem (Theorem 4.1).
HOPSE is more powerful than the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
Proof.
Since HOPSE matches the expressive power of CCWL, as established in Theorem C.5, to prove this theorem is sufficient to present a case in which two combinatorial complexes are indistinguishable by the CCWL test but are successfully distinguished by the HOPSE algorithm. The example in Figure 4 illustrates two such CCs. Using the neighborhood ${\mathcal{A}}_{2,1}$ (the faces are adjacent if they share an edge), the corresponding Hasse graphs ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$ are constructed, both of which are 3-regular. It is known that the WL test cannot distinguish between regular graphs of the same degree [29]. According to Proposition B.11 in [39], the CCWL test on these CCs is equivalent to the WL test on ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$, and thus also fails to distinguish them. However, the HOPSE algorithm can separate the two cases. In fact, by choosing the function $g_{1}(\cdot)$ to return the size of the connected component to which a node belongs, the difference between the two HOPSE encodings is apparent. ∎
Theorem (Theorem 4.2).
HOPSE is cell permutation equivariant if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are cell permutation equivariant for all $k\in{1,\dots,K}$ and all $r\in{0,\dots,R}$.
Proof.
To demonstrate that HOPSE is cell permutation equivariant, it suffices to show that the composition of permutation equivariant functions remains permutation equivariant. Since, by assumption, all functions that compose HOPSE are permutation equivariant, the result follows directly. Let $\pi(\cdot)$ be the permutation operator. Then $p(\cdot)$ is permutation equivariant if $\pi(p(\cdot))=p(\pi(\cdot))$. Given two permutaiton equivariant functions $p(\cdot)$ and $q(\cdot)$ their composition $r(\cdot)=(p\circ q)(\cdot)$ is also permutation equivariant:
$r(\pi(\cdot))=(p\circ q)(\pi(\cdot))=p(q(\pi(\cdot)))=p(\pi(q(\cdot)))=\pi(p(q(\cdot)))=\pi(r(\cdot)).$
Appendix D Framework realization
This section contains an in-depth description of the algorithmic procedure for extracting structural and positional information according two the proposed realizations HOPSE-M and HOPSE-G. Complementary to Section 4.3 the following is an in-depth explanation of the procedure for computing the positional and structural encodings of both realizations.
The first step in the pipeline is the definition of the family of functions $(g_{0},\dots,g_{K})$ and the same target aggregations $\psi_{k}(\cdot)$. Both the variations have a different way of calculating the output embeddings so both of the cases are shown.
D.1 Precomputation of HOPSE-M
In this case four hand-picked structural and positional encodings were used. Let ${\mathcal{G}}=(V,E)$ be an undirected graph with $|V|=n$ nodes. The individual channels $g_{k}({\mathcal{G}})$ are defined as follows:
  1. 1.
    $g_{1}(\cdot)=LapPE_{i}=[\phi_{1},\dots,\phi_{i}]\in\mathbb{R}^{n\times i},$ the first $i$ nontrivial eigenvectors of the graph Laplacian (up to frequency $i$).
  2. 2.
    $g_{2}(\cdot)=RWSE_{K}\in\mathbb{R}^{n\times K},$ the Random Walk Structural Encodings, obtained by stacking the diagonal entries of the $t$-step random-walk transition matrices for $t=1,\dots,K$ (here $K=60$).
  3. 3.
    $g_{3}(\cdot)=HKdiagSE_{K}\in\mathbb{R}^{n\times K},$ the Heat Kernel Diagonal Structural Encodings, given by $[\exp(-\lambda_{j}\,t)]_{j=1,\dots,n;\,t=1,\dots,K}$ for the Laplacian eigenvalues $\{\lambda_{j}\}$.
  4. 4.
    $g_{4}(\cdot)=ElstaticPE\in\mathbb{R}^{n\times d_{4}},$ the Electrostatic Potential Positional Encodings, computed by solving the discrete Poisson equation on ${\mathcal{G}}$ with suitable boundary conditions.
Importantly, since each $g_{k}$ depends only on the adjacency structure of ${\mathcal{G}}$, none of these channels requires access to any node feature vectors. For more indepth explanation of each PS encoding, see [13].
D.2 Precomputation of HOPSE-G
As mentioned previously, this method relies on the use of GPSE. This is a pre-trained graph PSE presented by  Cantürk et al. [13]. The pre-trained componenet consists of a GNN model pre-trained on a set of different datasets: ZINC, MolPCBA, GEOM, PCQM4Mv2 and ChEMBL. During the pre-training, a combination of losses targeting positional and structural properties is used to learn a latent representation of them.
In this case only one function $g_{1}=\text{GPSE}(\cdot)$ is used to calculate the encodings. It is worth noting that random noise initialization is used in GPSE to help the model learn some of the more challenging PEs. To emulate this procedure, for all neighborhoods ${\mathcal{N}}\in{\mathcal{N}}_{C}$, ${\mathcal{N}}_{C}\subseteq{\mathcal{N}}_{\mathcal{T}}$, and their corresponding Hasse graphs ${\mathcal{J}}_{{\mathcal{N}}}{\bigr{|}}_{k}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},{\mathcal{F}}_{\mathcal{N}})$, the features $F_{r,k}\in{\mathcal{F}}_{\mathcal{N}}$ are initialized as $F_{r,k}\sim N(0,\mathbf{I})\in\mathbb{R}^{20}$.
D.3 Learnable embedding functions
For the functions $f_{r,k}$, $\varepsilon_{r}$, and $\Theta_{r}$ HOPSE uses $L$-layer multilayer perceptrons (MLPs) with independent parameters. Each layer calculates the next embedding as:
$\mathbf{h}_{r,k}^{(\ell)}=\sigma\!\Biggl{(}\mathrm{LayerNorm}\Bigl{(}(\mathbf{h}_{r,k}^{(\ell-1)}\mathbf{W}_{r,k}^{(\ell)}+\mathbf{b}_{r,k}^{(\ell)})+\mathbf{h}_{r,k}^{(\ell-1)});\,\varepsilon=10^{-6}\Bigr{)}\Biggr{)},$ (14)
where each $\mathbf{W}_{r,k}^{(\ell)}\in\mathbb{R}^{D_{r}\times D_{r}}$ and $\mathbf{b}_{r,k}^{(\ell)}\in\mathbb{R}^{D_{r}}$ are learned, an additional $\mathbf{h}_{r,k}^{(\ell-1)}$ is added as a skip-connection, $\mathrm{LayerNorm}(\cdot;\varepsilon)$ normalizes to zero mean and unit variance with $\varepsilon=10^{-6}$ for numerical stability, and $\sigma$ is the pointwise nonlinearity LeakyReLU. The final output $\mathbf{h}_{r,k}^{(L)}$ yields the updated channel feature.
All MLPs share depth $L$ but may vary in hidden-layer widths $\{d_{r,\ell}\}$. This unified design—with layer normalization before activation—ensures stable gradient flow and flexible expressivity across channels and ranks.
Appendix E Complexity Analysis
To analyze the time complexity (in terms of FLOPs) of the HOPSE framework, distinguishing between two components is needed: (i) operations that can be precomputed once before training, and (ii) operations that must be performed during each training iteration.
Preprocessing overhead.
For positional and structural encodings, although the functions $g_{k}(\cdot)$ can, in principle, be any function—including trainable neural networks—the implementation is restricted to either graph-theoretic positional encodings or those generated by a pretrained GPSE model. This allows the tensors $\mathbf{X}_{r,k}$ (as defined in equation 2) to be pre-computed before training begins.
This is helpful because certain positional encodings can be computationally expensive. For instance, using the eigenvectors of the graph Laplacian requires eigendecomposition of a matrix $n\times n$, which has a complexity of ${\mathcal{O}}(n^{3})$. This makes such encodings impractical for large graphs if computed during training.
Train time computation.
The trainable components of the HOPSE framework are the functions $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$, for all $r=1,\dots,R$ and $k=1,\dots,K$. For these functions, HOPSE uses MLPs with $L$ layers
An MLP with $L$ layers applied to $N$ elements with input and output dimensions $D_{\text{in}}$ and $D_{\text{out}}$ has a computational complexity of ${\mathcal{O}}(ND_{\text{in}}D_{\text{out}})$. For simplicity, let us consider fixed input and output dimensions for each rank. The complexity of generating the initial features $\hat{\mathbf{Z}}_{r}$ in equation 4 is
${\mathcal{O}}(N_{r}D_{r}^{2}L),$
where $N_{r}$ is the number of cells of rank $r$, and $D_{r}$ is the feature dimension.
Next, the cost of computing $\mathbf{X}_{r,k}$ for all $K$ positional encodings (as in equation 3) is:
${\mathcal{O}}(N_{r}D_{r}^{2}KL),$
assuming that each function $f_{r,k}$ preserves the feature dimensionality.
Assuming concatenation is used as the aggregation operation in equation 5, and that the output features have dimension $D_{r}$, the complexity of computing the final embeddings $\mathbf{H}_{r}$ is:
${\mathcal{O}}(N_{r}D_{r}^{2}(K+1)L).$
Combining all components, the total complexity for rank-$r$ cells is:
$C_{r}={\mathcal{O}}\left(N_{r}D_{r}^{2}L+N_{r}D_{r}^{2}KL+N_{r}D_{r}^{2}(K+1)L\right)={\mathcal{O}}(N_{r}D_{r}^{2}(K+1)L).$
The final complexity, assuming that all cells have the same input and hidden dimensions, is
$C_{\text{HOPSE}}=\sum_{r=1}^{R}C_{r}.$
This complexity scales linearly with the number of rank-$r$ cells and is independent of their connectivity. This is in sharp contrast with most higher-order neural architectures. For example, the Generalized Combinatorial Complex Neural Network (GCCN) analyzed in [39] has the following complexity:
$C_{\text{GCCN}}=\sum_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}{\mathcal{O}}(\|\mathcal{N}\|_{0}D^{2}+\text{nrows}(\mathcal{N})d_{r^{\prime}}D+\text{nrows}(\mathcal{N}))+\sum_{r^{\prime}\in[0,R^{\prime}]}n_{r^{\prime}}n_{\mathcal{N}_{r^{\prime}}}D,$
where:
  • ${\mathcal{N}}_{C}$ is the set of neighborhoods considered,
  • $\mathcal{N}$ is a neighborhood from rank-$r$ to rank-$r^{\prime}$ cells,
  • $\|\mathcal{N}\|_{0}$ is the number of nonzero entries in $\mathcal{N}$,
  • $D$ is the feature dimension,
  • $d_{r^{\prime}}$ is the degree of rank-$r^{\prime}$ cells (assumed constant),
  • $n_{r^{\prime}}$ is the number of rank-$r^{\prime}$ cells,
  • $n_{\mathcal{N}_{r^{\prime}}}$ is the number of neighborhoods with target rank $r^{\prime}$.
Unlike HOPSE, the GCCN complexity depends on the number of cells and the structure of their relationships.
Summary.
This analysis shows that HOPSE achieves high expressive power (as demonstrated in Theorem 4.1) while maintaining a computational complexity that scales linearly with the number of cells. This is possible by offloading the structure-dependent computations to a one-time preprocessing step. The resulting efficiency, confirmed by the experimental results in Table 3, makes HOPSE well-suited for large-scale graph data.
Appendix F Extended results
All experiments were performed on a Linux system with 256 CPU cores, 1TB of RAM, and 8 NVIDIA A30 GPUs, each with 24GB of GPU memory. The implementation code and reproduction materials will be released upon manuscript acceptance.
F.1 Training and evaluation.
Five splits are generated for each dataset, with 50%/25%/25% of the data going to the training, validation, and test sets, respectively; the exception is ZINC, for which the predefined splits are used [27]. If available, this work considers the results reported in TopoBench for the baseline models. In the case of the HOPSE architecture and SaNN baseline, the benchmarking of the corresponding architectures follows the predefined TopoBench suite.
Following the literature, different performance metrics are reported per dataset. Specifically, predictive accuracy is used for MUTAG, PROTEINS, NCI1, and NCI109. Mean absolute error (MAE) is used for ZINC. F1 score is used for MANTRA datasets. For each dataset, the mean and standard deviation of the chosen metric are computed across the five test sets.
F.2 Hyperparameter search
The HOPSE architecture is optimized over eight variations in the neighborhood set. Each variation specifies a combination of neighborhoods. The considered neighborhood combinations are as follows:
  • ${\mathcal{A}}_{0,1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$, ${\mathcal{A}}_{2,1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{I}}_{0\to 1}$, ${\mathcal{I}}_{1\to 2}$
  • ${\mathcal{A}}_{0,1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{I}}_{0\to 1}$, ${\mathcal{I}}_{1\to 2}$, ${\mathcal{I}}_{1\to 0}$, ${\mathcal{I}}_{2\to 1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$, ${\mathcal{A}}_{1,0}$, ${\mathcal{A}}_{2,1}$, ${\mathcal{I}}_{0\to 1}$, ${\mathcal{I}}_{1\to 2}$, ${\mathcal{I}}_{1\to 0}$, ${\mathcal{I}}_{2\to 1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$, ${\mathcal{A}}_{0,2}$, ${\mathcal{A}}_{1,0}$, ${\mathcal{A}}_{2,1}$, ${\mathcal{A}}_{2,0}$
The first three configurations rely solely on adjacency-based neighborhoods. The following four integrate adjacency and incidence relations, while the final configuration includes jumps across ranks.
For both HOPSE-M and HOPSE-G, the set of neighborhood configurations described above is incorporated into the architecture search space, allowing the models to explore diverse neighborhoods.
HOPSE-G leverages the GPSE model to extract both positional and structural information. The evaluation is on four variants of the GPSE model, each pre-trained on a different dataset: ZINC, GEOM, MOLPCBA, and PCQM4MV2. These pre-trained models are treated as hyperparameters during the hyperparameter search. Overall, results show that GPSE models pre-trained on different datasets generalize well and consistently capture meaningful positional and structural features across all evaluated benchmarks, with no single pre-trained model consistently outperforming the others.
For the SaNN model, the number of hops needed to define both positional and structural encodings was treated as an additional hyperparameter.
Finally, the search space over architecture parameters is restricted to hyperparameters that are common across all models: SaNN, HOPSE, SCCNN, and SCN. A grid-search strategy is employed to identify the optimal parameters for each model-dataset combination. Specifically:
  • Encoder hidden dimension: {32, 64, 128, 256}
  • Encoder dropout: {0.25, 0.5}
  • Number of backbone layers: {1, 2, 3, 4}
  • Learning rate: {0.01, 0.001}
  • Weight decay: {0, 0.0001}
  • Batch size: {128, 256}
For SCN and SCCNN models, the considered readout types are direct readout (DR) and signal down-propagation (SDP) (refer to Telyatnikov et al. [42] for in-depth discussion on these options).
F.3 Datasets
The datasets considered are five standard benchmarking datasets commonly used in the TDL literature, comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^4] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [35]. For regression, the ZINC dataset is used [27]. Following standard practice in the TDL literature [42], liftings are applied to transform these graph datasets into higher-order domains. Specifically, the cycle-based lifting was used for the cell domain, and the clique complex lifting for the simplicial domain.
Additionally, the MANTRA dataset comprises different triangulations of manifolds. A triangulation of a manifold ${\mathcal{M}}$ is a pair $(K,\iota)$ where $K$ is a simplicial complex and $\iota$ is a homeomorphism between ${\mathcal{M}}$ and the geometric realization of $K$. Concretely, triangulations encode higher-order topological information that the underlying graph, or 00-skeleton, does not encode. Since some triangulations are non-homeomorphic but share an underlying graph, the ability to distinguish triangulations requires that the model can access the topological information to make the correct predictions. This dataset emphasizes benchmarking a model’s predictive ability on purely topological properties. Thus, it aims to assess how much the model encodes topological information beyond the base graph. In contrast to the MANTRA paper, this work does not enhance node features with any additional manually derived features but instead initializes features for all cells as ones.
Regarding $\beta_{0},\beta_{1},\beta_{2}$, following the Ballester et al. [2] work, in the case of predicting Betti numbers, this work focuses on training a single model as a multivariate regression task to report corresponding results for each Betti number. The choice of using F1 over AUROC is based on class imbalances of the Betti numbers $\beta_{0},\beta_{1},\beta_{2}$ and of the predicted properties name (NAME) and orientability ORIENT.
F.4 Runtime per epoch
In these experiments, the runtime per epoch is shown for each dataset. Table 4 shows the average runtime per epoch on the benchmarked datasets. Our method has the lowest runtime per epoch across all datasets among HOMP methods. These results validate the theoretical findings presented. In addition, HOPSE-G is faster than HOPSE-M in all benchmarks except NCI109. This aligns with complexity and scales linearly with the number $K$ of positional encodings considered.
Compared to graph message passing methods, such as GCN, GAT, and GIN, the HOPSE framework performs competitively on smaller datasets such as MUTAG, PROTEINS, NCI1, and NCI109. On larger datasets, it speeds up up to 4x in ZINC and closer to 2x in MANTRA.
Table 4: Runtime per epoch in seconds. Results are shown as mean and standard deviation. The best result per domain and per model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. The HOPSE framework is more efficient than all other HOMP-based methods across all datasets. In the larger datasets such as ZINC and MANTRA, it is more efficient than traditional message passing over graphs. The variation between the simplicial and cell domain for both models is relatively low
Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 0.03 $\pm$ 0.00 0.05 $\pm$ 0.00 0.26 $\pm$ 0.01 0.27 $\pm$ 0.02 1.24 $\pm$ 0.01 6.24 $\pm$ 0.11 6.24 $\pm$ 0.10 6.43 $\pm$ 0.08 GAT 0.04 $\pm$ 0.00 0.07 $\pm$ 0.00 0.34 $\pm$ 0.02 0.33 $\pm$ 0.02 1.24 $\pm$ 0.01 6.38 $\pm$ 0.14 6.16 $\pm$ 0.12 6.44 $\pm$ 0.15 GIN 0.03 $\pm$ 0.00 0.06 $\pm$ 0.00 0.27 $\pm$ 0.02 0.27 $\pm$ 0.02 1.19 $\pm$ 0.00 6.37 $\pm$ 0.12 6.18 $\pm$ 0.06 6.46 $\pm$ 0.23 Simplicial SCN 0.09 $\pm$ 0.00 0.42 $\pm$ 0.01 1.60 $\pm$ 0.03 1.59 $\pm$ 0.02 6.93 $\pm$ 0.09 6.83 $\pm$ 0.08 7.35 $\pm$ 0.38 6.92 $\pm$ 0.08 SCCNN 0.09 $\pm$ 0.00 0.43 $\pm$ 0.01 1.65 $\pm$ 0.02 1.66 $\pm$ 0.03 10.40 $\pm$ 0.06 8.47 $\pm$ 0.63 8.25 $\pm$ 0.12 9.24 $\pm$ 0.15 SaNN 0.03 $\pm$ 0.00 0.18 $\pm$ 0.00 0.64 $\pm$ 0.02 0.65 $\pm$ 0.02 0.63 $\pm$ 0.03 7.21 $\pm$ 0.08 7.77 $\pm$ 0.15 7.34 $\pm$ 0.10 GCCN 0.04 $\pm$ 0.00 3.06 $\pm$ 0.02 0.43 $\pm$ 0.02 0.39 $\pm$ 0.02 - 6.79 $\pm$ 0.24 6.13 $\pm$ 0.09 7.70 $\pm$ 0.05 HOPSE-M 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00 0.32 $\pm$ 0.00 0.47 $\pm$ 0.03 0.35$\pm$ 0.01 5.13 $\pm$ 0.14 5.60 $\pm$ 0.55 5.54 $\pm$ 0.05 HOPSE-G 0.02 $\pm$ 0.00 0.10 $\pm$ 0.02 0.31 $\pm$ 0.00 0.37 $\pm$ 0.02 0.28 $\pm$ 0.01 4.48 $\pm$ 0.14 4.23 $\pm$ 0.12 4.48 $\pm$ 0.11 Cell CCCN 0.11 $\pm$ 0.00 0.32 $\pm$ 0.01 1.33 $\pm$ 0.02 1.33 $\pm$ 0.02 6.14 $\pm$ 0.04 - - - CWN 0.10 $\pm$ 0.00 0.37 $\pm$ 0.02 1.38 $\pm$ 0.03 1.37 $\pm$ 0.01 5.86 $\pm$ 0.05 - - - HOPSE-M 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.34 $\pm$ 0.00 0.33 $\pm$ 0.01 0.47 $\pm$ 0.02 - - - HOPSE-G 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.40 $\pm$ 0.02 0.37 $\pm$ 0.00 0.32 $\pm$ 0.01 - - -
F.5 Ablation on neighborhood set ${\mathcal{N}}_{C}$
To gain more insight into how the choice of neighborhoods ${\mathcal{N}}_{C}$ impacts the performance of the HOPSE framework, this section shows an ablation over the considered sets of neighborhoods. The neighborhood sets listed in the tables are defined in Table 5. Each configuration differs in the number and type of neighborhoods considered, directly influencing computational complexity and model expressivity. Specifically, the Adj family explores adjacency-based neighborhoods of increasing order: Adj-1 uses a single neighborhood, while Adj-2 and Adj-3 consider two and three neighborhoods, respectively. The Inc configurations mix the base adjacency (${\mathcal{A}}_{0,1}$) and incidence, with Inc-1 and Inc-2 using three neighborhoods, and Inc-3 extending to five. The Mix configurations, Mix-1 and Mix-2, are the biggest, combining a total of eight and six neighborhoods, respectively. Mix-1 uses all previous adjacencies and incidences combined while Mix-2 includes more complex neighborhoods. One of such is ${\mathcal{A}}_{0,2}$, which means the adjacencies of 00-cells with $2$-cells through $1$-cells. Effectively, this means that the structure of neighborhoods between non-adjacent ranks (i.e., a difference of $1$ in the rank number) is considered. As will be seen later in the section, considering these non-standard neighborhoods does account for a difference in performance for the more complex benchmarks.
Table 5: Taxonomy of neighborhood combinations
Group Neighborhood Combination
Adj-1 ${\mathcal{A}}_{0,1}$
Adj-2 ${\mathcal{A}}_{0,1},A_{1,2}$
Adj-3 ${\mathcal{A}}_{0,1},A_{1,2},A_{2,1}$
Inc-1 ${\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2}$
Inc-2 ${\mathcal{A}}_{0,1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
Inc-3 ${\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
Mix-1 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1}$
Mix-2 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{0,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{A}}_{2,0}$
F.5.1 Analysis of performance across neighborhoods
Analysis of performance for different domains.
The results in Table 6 and Table 7 demonstrate insightful patterns in how neighborhood selection impacts performance depending on domain type and dataset complexity. Comparing performance between HOPSE-M and HOPSE-G variants across different neighborhood configurations further suggests that the pre-training strategy employed in GPSE provides rich neighborhood information, particularly for the molecular datasets, as the GPSE encoder has been pre-trained on molecular data. This is most evident in the cell domain results for ZINC, where HOPSE-G with Mix-2 achieves a 16% improvement compared to the best HOPSE-M configuration.
Table 6: Performance per neighborhood on simplicial. Results are shown as mean $\pm$ standard deviation. The best result per neighborhood is in bold and results within one standard deviation of the best per neighborhood are highlighted in blue. The performance differences between neighborhood configurations are minimal, with most results falling within one standard deviation of each other. As datasets grow bigger and more complex, such as ZINC and the MANTRA benchmarks, the neighborhood choice starts to matter. The results show that for simplicial models, the most informative neighborhoods for ZINC are the extended groups of adjacencies Adj-2 and Adj-3. However, for the MANTRA tasks, the most important neighborhood is Mix-2
Graph Tasks Topological Tasks Nbhd MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) HOPSE-M Adj-1 86.38 $\pm$ 6.31 74.70 $\pm$ 4.04 77.35 $\pm$ 0.81 75.41 $\pm$ 1.17 0.40 $\pm$ 0.01 79.58 $\pm$ 2.88 60.93 $\pm$ 3.03 88.35 $\pm$ 0.09 28.25 $\pm$ 2.71 Adj-2 85.53 $\pm$ 7.12 73.19 $\pm$ 3.32 76.26 $\pm$ 1.59 75.59 $\pm$ 0.41 0.39 $\pm$ 0.02 80.35 $\pm$ 3.96 74.63 $\pm$ 1.77 89.97 $\pm$ 0.18 64.06 $\pm$ 2.27 Adj-3 85.96 $\pm$ 5.75 75.13 $\pm$ 3.66 76.44 $\pm$ 1.44 75.22 $\pm$ 2.06 0.39 $\pm$ 0.02 84.72 $\pm$ 1.23 78.19 $\pm$ 2.70 90.13 $\pm$ 0.49 70.10 $\pm$ 1.86 Inc-1 85.11 $\pm$ 6.20 73.98 $\pm$ 2.97 76.40 $\pm$ 0.99 76.09 $\pm$ 1.50 0.40 $\pm$ 0.01 84.01 $\pm$ 6.84 65.10 $\pm$ 9.66 90.06 $\pm$ 0.09 65.02 $\pm$ 1.58 Inc-2 84.26 $\pm$ 6.13 73.69 $\pm$ 2.76 76.40 $\pm$ 1.45 75.39 $\pm$ 1.13 0.42 $\pm$ 0.05 80.67 $\pm$ 2.88 69.51 $\pm$ 6.86 89.51 $\pm$ 0.94 64.28 $\pm$ 3.08 Inc-3 85.11 $\pm$ 5.83 73.55 $\pm$ 3.38 76.56 $\pm$ 1.40 76.34 $\pm$ 1.40 0.42 $\pm$ 0.02 82.56 $\pm$ 4.15 77.35 $\pm$ 2.12 89.89 $\pm$ 0.13 63.12 $\pm$ 1.68 Mix-1 84.26 $\pm$ 5.12 73.69 $\pm$ 3.21 76.44 $\pm$ 1.62 75.57 $\pm$ 1.09 0.40 $\pm$ 0.01 90.37 $\pm$ 1.29 72.64 $\pm$ 7.53 90.21 $\pm$ 0.18 67.14 $\pm$ 2.18 Mix-2 83.83 $\pm$ 6.49 73.41 $\pm$ 2.95 75.14 $\pm$ 1.59 74.87 $\pm$ 0.94 0.41 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G Adj-1 82.98 $\pm$ 7.22 74.05 $\pm$ 2.17 77.18 $\pm$ 1.45 76.21 $\pm$ 0.65 0.44 $\pm$ 0.01 80.43 $\pm$ 1.03 60.92 $\pm$ 1.18 88.18 $\pm$ 0.03 24.59 $\pm$ 1.87 Adj-2 85.53 $\pm$ 2.77 74.05 $\pm$ 3.08 77.51 $\pm$ 1.15 74.91 $\pm$ 1.17 0.43 $\pm$ 0.00 80.19 $\pm$ 1.13 61.88 $\pm$ 0.83 88.13 $\pm$ 0.02 20.31 $\pm$ 2.57 Adj-3 85.11 $\pm$ 3.36 73.26 $\pm$ 4.20 77.32 $\pm$ 1.19 74.79 $\pm$ 0.91 0.43 $\pm$ 0.01 79.11 $\pm$ 1.53 61.86 $\pm$ 1.10 88.11 $\pm$ 0.02 19.08 $\pm$ 3.57 Inc-1 84.68 $\pm$ 3.50 73.19 $\pm$ 3.52 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.44 $\pm$ 0.00 80.70 $\pm$ 0.80 62.17 $\pm$ 0.98 87.31 $\pm$ 1.17 20.47 $\pm$ 3.22 Inc-2 82.13 $\pm$ 5.75 76.27 $\pm$ 2.10 75.66 $\pm$ 1.52 74.17 $\pm$ 1.99 0.45 $\pm$ 0.02 78.92 $\pm$ 1.25 59.57 $\pm$ 1.56 88.08 $\pm$ 0.08 20.28 $\pm$ 3.48 Inc-3 77.87 $\pm$ 4.90 75.48 $\pm$ 3.15 76.15 $\pm$ 1.77 75.14 $\pm$ 0.75 0.44 $\pm$ 0.01 78.78 $\pm$ 0.74 60.35 $\pm$ 1.05 88.12 $\pm$ 0.04 18.82 $\pm$ 3.54 Mix-1 79.15 $\pm$ 7.88 75.70 $\pm$ 3.24 76.50 $\pm$ 1.09 75.35 $\pm$ 0.97 0.45 $\pm$ 0.01 79.76 $\pm$ 0.51 60.33 $\pm$ 1.60 88.07 $\pm$ 0.09 23.48 $\pm$ 3.44 Mix-2 83.40 $\pm$ 8.56 74.05 $\pm$ 3.39 77.02 $\pm$ 0.92 74.75 $\pm$ 1.47 0.46 $\pm$ 0.01 81.75 $\pm$ 1.26 61.32 $\pm$ 2.67 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25
Efficacy of neighborhoods by domain.
One important observation is how neighborhood efficacy varies between the simplicial and cell domains. While in the simplicial domain, adding mixed neighborhoods occasionally leads to performance degradation for some datasets (e.g., PROTEINS where Inc-1 outperforms Mix configurations for HOPSE-M), in the cell domain, more comprehensive neighborhood sets either improve or do not significantly impact performance. This domain-dependent behavior suggests that the optimal neighborhood configuration should be ideally selected based on task complexity and the topological representation chosen for the data; however, choosing comprehensive neighborhood sets on average allows for achieving good performance while not harming the computational cost (see Section F.5.2 for details on computational analysis). These findings demonstrate that the choice of neighborhood configuration represents a critical design decision in topological learning frameworks. While simpler configurations may suffice for some graph benchmarks, an effective strategy is to consider a greater neighbourhood set ${\mathcal{N}}_{C}$ to capture the full expressive power of topological structures. It is especially evident on complex tasks, topological benchmarking datasets (MANTRA), and relatively large datasets like ZINC.
Table 7: Performance per neighborhood on cell. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. In smaller and simpler datasets such as MUTAG, PROTEINS, NCI1, and NCI109, the best neighborhoods are not always the most encompassing, but they perform closely to others. However, on a more complex dataset such as ZINC there is a significant difference between simpler neighborhoods like the Adj[1-3] and Mix-2, particularly in the case of HOPSE-G.
Graph Tasks Nbhd MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) HOPSE-M Adj-1 88.09 $\pm$ 5.55 75.77 $\pm$ 2.67 76.85 $\pm$ 0.83 76.34 $\pm$ 1.02 0.29 $\pm$ 0.01 Adj-2 82.13 $\pm$ 4.15 74.55 $\pm$ 2.65 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.27 $\pm$ 0.01 Adj-3 85.53 $\pm$ 3.50 75.27 $\pm$ 2.19 77.57 $\pm$ 1.48 75.37 $\pm$ 0.79 0.28 $\pm$ 0.01 Inc-1 86.81 $\pm$ 4.61 76.99 $\pm$ 2.63 77.32 $\pm$ 1.62 75.70 $\pm$ 1.36 0.27 $\pm$ 0.02 Inc-2 83.40 $\pm$ 3.81 75.56 $\pm$ 2.26 75.89 $\pm$ 2.14 75.43 $\pm$ 1.98 0.29 $\pm$ 0.02 Inc-3 82.98 $\pm$ 3.36 74.70 $\pm$ 3.05 77.14 $\pm$ 0.46 74.46 $\pm$ 1.52 0.25 $\pm$ 0.01 Mix-1 81.28 $\pm$ 5.90 75.13 $\pm$ 3.09 77.26 $\pm$ 1.70 76.13 $\pm$ 1.38 0.25 $\pm$ 0.01 Mix-2 84.26 $\pm$ 6.13 72.62 $\pm$ 2.55 76.69 $\pm$ 1.25 76.57 $\pm$ 1.55 0.26 $\pm$ 0.01 HOPSE-G Adj-1 82.13 $\pm$ 7.76 76.27 $\pm$ 2.70 78.31 $\pm$ 0.84 77.48 $\pm$ 0.99 0.26 $\pm$ 0.00 Adj-2 87.66 $\pm$ 6.96 76.42 $\pm$ 2.19 78.83 $\pm$ 1.40 77.48 $\pm$ 0.64 0.24 $\pm$ 0.00 Adj-3 88.09 $\pm$ 1.90 75.91 $\pm$ 1.89 78.35 $\pm$ 1.11 76.94 $\pm$ 1.46 0.23 $\pm$ 0.00 Inc-1 85.11 $\pm$ 6.56 74.62 $\pm$ 3.99 77.57 $\pm$ 1.67 76.81 $\pm$ 1.23 0.25 $\pm$ 0.01 Inc-2 82.55 $\pm$ 8.82 74.91 $\pm$ 2.11 78.15 $\pm$ 0.75 76.40 $\pm$ 1.86 0.27 $\pm$ 0.01 Inc-3 84.68 $\pm$ 7.88 74.12 $\pm$ 3.39 76.81 $\pm$ 0.59 75.26 $\pm$ 1.16 0.26 $\pm$ 0.00 Mix-1 88.09 $\pm$ 3.23 76.20 $\pm$ 2.62 79.03 $\pm$ 0.73 76.88 $\pm$ 0.95 0.23 $\pm$ 0.00 Mix-2 85.11 $\pm$ 6.02 75.63 $\pm$ 2.35 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01
F.5.2 Analysis of time efficiency across neighborhoods
This section presents the effects of the choice of ${\mathcal{N}}_{C}$ on the total runtime and per-epoch performance of the HOPSE framework. The results in Table 8 and Table 9 report the per-epoch runtime for each neighborhood configuration across both the simplicial and cell domains, under the HOPSE-M and HOPSE-G models, respectively. While Table 10 and Table 11 present the total runtime to convergence for eight different neighbourhoods for simplicial and cell domains, and HOPSE-M and HOPSE-G correspondingly.
Table 8: Runtime per epoch per neighborhood on simplicial. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. Simpler configurations such as Adj-1 consistently yield the lowest runtime across domains, particularly for HOPSE-M, where Adj-1 achieves the fastest epoch times in 5 out of 8 datasets,
Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.32 $\pm$ 0.00 0.32 $\pm$ 0.00 0.30 $\pm$ 0.01 4.12 $\pm$ 0.08 4.41 $\pm$ 0.11 4.48 $\pm$ 0.16 Adj-2 0.03 $\pm$ 0.01 0.12 $\pm$ 0.00 0.36 $\pm$ 0.01 0.36 $\pm$ 0.00 0.35 $\pm$ 0.01 4.37 $\pm$ 0.13 4.15 $\pm$ 0.09 4.54 $\pm$ 0.21 Adj-3 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.37 $\pm$ 0.00 0.39 $\pm$ 0.00 0.37 $\pm$ 0.01 4.42 $\pm$ 0.11 4.41 $\pm$ 0.22 4.91 $\pm$ 0.07 Inc-1 0.03 $\pm$ 0.00 0.14 $\pm$ 0.01 0.44 $\pm$ 0.01 0.40 $\pm$ 0.01 0.40 $\pm$ 0.01 6.28 $\pm$ 1.34 5.14 $\pm$ 0.24 5.41 $\pm$ 0.17 Inc-2 0.03 $\pm$ 0.01 0.12 $\pm$ 0.01 0.36 $\pm$ 0.00 0.39 $\pm$ 0.00 0.34 $\pm$ 0.01 4.41 $\pm$ 0.05 4.29 $\pm$ 0.06 4.87 $\pm$ 0.12 Inc-3 0.02 $\pm$ 0.00 0.16 $\pm$ 0.00 0.46 $\pm$ 0.01 0.47 $\pm$ 0.03 0.42 $\pm$ 0.01 5.89 $\pm$ 0.31 5.71 $\pm$ 0.23 5.70 $\pm$ 0.23 Mix-1 0.03 $\pm$ 0.00 0.17 $\pm$ 0.01 0.51 $\pm$ 0.00 0.51 $\pm$ 0.02 0.48 $\pm$ 0.01 6.00 $\pm$ 0.29 6.25 $\pm$ 0.31 6.51 $\pm$ 0.23 Mix-2 0.02 $\pm$ 0.00 0.15 $\pm$ 0.00 0.45 $\pm$ 0.01 0.44 $\pm$ 0.00 0.40 $\pm$ 0.01 5.13 $\pm$ 0.14 5.60 $\pm$ 0.55 5.54 $\pm$ 0.05 HOPSE-G Adj-1 0.02 $\pm$ 0.00 0.10 $\pm$ 0.01 0.28 $\pm$ 0.01 0.28 $\pm$ 0.00 0.26 $\pm$ 0.02 3.19 $\pm$ 0.09 3.18 $\pm$ 0.12 3.55 $\pm$ 0.12 Adj-2 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.30 $\pm$ 0.00 0.31 $\pm$ 0.01 0.28 $\pm$ 0.01 3.53 $\pm$ 0.12 3.41 $\pm$ 0.10 3.67 $\pm$ 0.12 Adj-3 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.33 $\pm$ 0.02 0.34 $\pm$ 0.01 0.29 $\pm$ 0.02 3.79 $\pm$ 0.14 3.84 $\pm$ 0.14 3.98 $\pm$ 0.07 Inc-1 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.38 $\pm$ 0.01 0.37 $\pm$ 0.02 0.34 $\pm$ 0.01 4.44 $\pm$ 0.14 4.23 $\pm$ 0.12 4.64 $\pm$ 0.08 Inc-2 0.02 $\pm$ 0.00 0.10 $\pm$ 0.00 0.31 $\pm$ 0.00 0.31 $\pm$ 0.00 0.28 $\pm$ 0.01 3.74 $\pm$ 0.17 3.66 $\pm$ 0.10 3.78 $\pm$ 0.06 Inc-3 0.02 $\pm$ 0.00 0.13 $\pm$ 0.00 0.41 $\pm$ 0.00 0.40 $\pm$ 0.01 0.38 $\pm$ 0.03 5.17 $\pm$ 0.40 4.75 $\pm$ 0.23 5.13 $\pm$ 0.06 Mix-1 0.02 $\pm$ 0.00 0.16 $\pm$ 0.01 0.46 $\pm$ 0.02 0.46 $\pm$ 0.01 0.44 $\pm$ 0.01 5.50 $\pm$ 0.24 5.74 $\pm$ 0.22 5.84 $\pm$ 0.38 Mix-2 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.38 $\pm$ 0.00 0.38 $\pm$ 0.01 0.36 $\pm$ 0.02 4.48 $\pm$ 0.14 4.55 $\pm$ 0.19 4.84 $\pm$ 0.11
Table 9: Runtime per epoch per neighborhood on cell. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.31 $\pm$ 0.01 0.32 $\pm$ 0.01 0.30 $\pm$ 0.01 Adj-2 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.34 $\pm$ 0.00 0.33 $\pm$ 0.00 0.30 $\pm$ 0.01 Adj-3 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.35 $\pm$ 0.00 0.34 $\pm$ 0.00 0.32 $\pm$ 0.01 Inc-1 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.40 $\pm$ 0.01 0.40 $\pm$ 0.02 0.37 $\pm$ 0.02 Inc-2 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00 0.35 $\pm$ 0.01 0.34 $\pm$ 0.01 0.32 $\pm$ 0.01 Inc-3 0.03 $\pm$ 0.01 0.15 $\pm$ 0.00 0.42 $\pm$ 0.01 0.43 $\pm$ 0.01 0.41 $\pm$ 0.01 Mix-1 0.03 $\pm$ 0.00 0.16 $\pm$ 0.01 0.50 $\pm$ 0.01 0.54 $\pm$ 0.01 0.47 $\pm$ 0.02 Mix-2 0.02 $\pm$ 0.00 0.13 $\pm$ 0.00 0.47 $\pm$ 0.02 0.46 $\pm$ 0.00 0.39 $\pm$ 0.01 HOPSE-G Adj-1 0.02 $\pm$ 0.00 0.09 $\pm$ 0.01 0.25 $\pm$ 0.01 0.26 $\pm$ 0.01 0.24 $\pm$ 0.00 Adj-2 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.29 $\pm$ 0.01 0.27 $\pm$ 0.01 0.26 $\pm$ 0.01 Adj-3 0.02 $\pm$ 0.00 0.10 $\pm$ 0.00 0.32 $\pm$ 0.01 0.32 $\pm$ 0.02 0.27 $\pm$ 0.01 Inc-1 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.37 $\pm$ 0.01 0.34 $\pm$ 0.00 0.33 $\pm$ 0.00 Inc-2 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.32 $\pm$ 0.01 0.30 $\pm$ 0.05 0.28 $\pm$ 0.02 Inc-3 0.02 $\pm$ 0.00 0.14 $\pm$ 0.01 0.40 $\pm$ 0.00 0.37 $\pm$ 0.00 0.36 $\pm$ 0.02 Mix-1 0.03 $\pm$ 0.00 0.14 $\pm$ 0.00 0.44 $\pm$ 0.01 0.45 $\pm$ 0.00 0.42 $\pm$ 0.00 Mix-2 0.02 $\pm$ 0.00 0.14 $\pm$ 0.00 0.40 $\pm$ 0.02 0.37 $\pm$ 0.01 0.32 $\pm$ 0.01
Runtime per epoch per neighborhood.
As expected Table 8 and Table 9 show that the runtime per epoch increases as the number of neighborhoods in ${\mathcal{N}}_{\mathcal{C}}$ increases. This aligns with the complexity analysis presented in Appendix E. Still, the relative increase in computational time is at most 80%, which is acceptable when considering how many more higher-order relations are being considered.
The efficiency-expressivity trade-off revealed in these results has important implications for the practical deployment. While intuition might suggest that more complex neighborhood configurations would always incur substantial computational penalties, our findings demonstrate that this is not necessarily true.
Despite incorporating significantly more structural information, complex neighborhood configurations like Mix-1 and Mix-2 introduce only a modest increase in per-epoch runtime, typically less than 2×2$\times$2 × compared to the simplest configurations. This makes them viable options even in computationally constrained environments.
Total runtime per neighborhood.
The total runtime required to reach convergence, presented in Table 10 and Table 11, reveals more nuanced patterns. Unlike per-epoch runtime, total runtime encapsulates both computational cost per epoch and the number of epochs to convergence, which are indirectly influenced by the expressiveness of the neighborhood structure. Interestingly, richer neighborhood configurations, despite their potentially higher per-epoch cost, often lead to a similar relative increase in time needed as the per-epoch case. This means that the number of epochs needed for convergence stays relatively constant.
Table 10: Total runtime per neighborhood on simplicial. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. HOPSE configurations, especially with mixed neighborhood sets, provide an excellent balance between efficiency and expressiveness.
Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 6.76 $\pm$ 1.38 44.60 $\pm$ 17.17 164.10 $\pm$ 79.48 125.46 $\pm$ 19.54 1238.09 $\pm$ 378.91 2652.62 $\pm$ 67.51 3028.14 $\pm$ 41.05 5327.90 $\pm$ 1334.91 Adj-2 8.66 $\pm$ 0.51 31.26 $\pm$ 15.10 165.10 $\pm$ 31.84 131.47 $\pm$ 24.84 1512.34 $\pm$ 310.43 3036.30 $\pm$ 1105.71 2535.98 $\pm$ 45.80 4053.39 $\pm$ 718.36 Adj-3 6.87 $\pm$ 1.54 38.41 $\pm$ 8.07 175.02 $\pm$ 49.19 146.19 $\pm$ 27.88 $1$$\pm$ 404.64 2734.33 $\pm$ 156.56 2689.60 $\pm$ 38.86 4266.24 $\pm$ 651.78 Inc-1 8.52 $\pm$ 1.47 41.92 $\pm$ 5.67 194.90 $\pm$ 64.73 206.44 $\pm$ 16.44 1183.26 $\pm$ 135.82 4351.87 $\pm$ 1203.66 3342.21 $\pm$ 88.20 5172.86 $\pm$ 1084.20 Inc-2 10.37 $\pm$ 2.21 40.03 $\pm$ 20.40 161.40 $\pm$ 39.23 180.17 $\pm$ 22.35 895.12 $\pm$ 295.76 2704.44 $\pm$ 23.84 2998.32 $\pm$ 790.12 4859.76 $\pm$ 819.26 Inc-3 8.81 $\pm$ 3.00 39.99 $\pm$ 6.31 209.70 $\pm$ 38.55 218.69 $\pm$ 25.83 1332.58 $\pm$ 316.15 3493.64 $\pm$ 62.70 4076.33 $\pm$ 671.42 5465.95 $\pm$ 1816.61 Mix-1 9.81 $\pm$ 2.93 54.94 $\pm$ 13.50 214.50 $\pm$ 7.54 198.45 $\pm$ 54.21 1288.78 $\pm$ 590.16 4201.85 $\pm$ 851.36 4315.76 $\pm$ 1097.48 6542.09 $\pm$ 1272.01 Mix-2 7.63 $\pm$ 1.71 43.67 $\pm$ 18.29 229.05 $\pm$ 65.67 189.19 $\pm$ 45.39 1434.43 $\pm$ 547.60 3596.38 $\pm$ 519.37 3408.67 $\pm$ 162.45 5377.27 $\pm$ 185.49 HOPSE-G Adj-1 4.56 $\pm$ 0.62 25.49 $\pm$ 6.16 92.97 $\pm$ 27.48 104.94 $\pm$ 28.21 542.03 $\pm$ 94.68 1891.07 $\pm$ 17.81 1885.35 $\pm$ 28.06 2194.20 $\pm$ 42.70 Adj-2 4.92 $\pm$ 0.37 32.03 $\pm$ 9.91 118.32 $\pm$ 17.84 111.18 $\pm$ 18.27 480.02 $\pm$ 100.70 2172.54 $\pm$ 216.62 2096.15 $\pm$ 33.94 2323.75 $\pm$ 222.16 Adj-3 5.58 $\pm$ 1.28 39.85 $\pm$ 17.90 120.88 $\pm$ 31.47 134.65 $\pm$ 49.81 478.06 $\pm$ 107.13 2302.00 $\pm$ 47.08 2385.63 $\pm$ 176.04 2997.61 $\pm$ 824.53 Inc-1 5.80 $\pm$ 0.27 32.71 $\pm$ 8.43 130.06 $\pm$ 22.37 106.98 $\pm$ 27.98 649.05 $\pm$ 114.28 2580.45 $\pm$ 47.63 2471.89 $\pm$ 47.46 3716.82 $\pm$ 1192.83 Inc-2 6.44 $\pm$ 0.54 32.98 $\pm$ 8.25 128.01 $\pm$ 22.39 99.39 $\pm$ 33.07 661.57 $\pm$ 133.52 2489.63 $\pm$ 444.47 2240.50 $\pm$ 43.84 2568.75 $\pm$ 76.80 Inc-3 6.85 $\pm$ 0.98 51.18 $\pm$ 11.23 174.20 $\pm$ 57.02 148.67 $\pm$ 37.09 778.78 $\pm$ 117.23 3038.25 $\pm$ 65.40 2930.43 $\pm$ 430.59 3602.56 $\pm$ 555.32 Mix-1 6.85 $\pm$ 0.72 51.94 $\pm$ 9.59 209.07 $\pm$ 28.40 150.07 $\pm$ 36.64 911.69 $\pm$ 165.22 3408.18 $\pm$ 515.16 3379.79 $\pm$ 111.99 4321.64 $\pm$ 1094.92 Mix-2 5.16 $\pm$ 0.36 35.97 $\pm$ 4.38 168.30 $\pm$ 38.75 142.36 $\pm$ 34.54 716.46 $\pm$ 53.19 3324.16 $\pm$ 1113.48 2794.73 $\pm$ 345.73 3876.34 $\pm$ 1141.15
Table 11: Total runtime per neighborhood on cell. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. On HOPSE-G, some more complex neighborhoods will result in faster time, even though they contain more information and thus parameters to be learned. This is one of the effects of convergence speed. This can be seen in neighborhoods Mix-1 and Mix-2. On HOPSE-M, there is no such drastic difference
Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 7.03 $\pm$ 0.86 37.52 $\pm$ 2.06 144.11 $\pm$ 32.07 142.87 $\pm$ 45.01 1128.14 $\pm$ 166.57 Adj-2 7.66 $\pm$ 0.82 35.68 $\pm$ 7.88 147.04 $\pm$ 24.07 121.89 $\pm$ 18.30 901.44 $\pm$ 203.17 Adj-3 9.78 $\pm$ 2.94 37.85 $\pm$ 10.19 153.94 $\pm$ 29.39 131.39 $\pm$ 36.31 1414.14 $\pm$ 325.91 Inc-1 7.86 $\pm$ 0.54 37.06 $\pm$ 7.05 174.19 $\pm$ 33.02 190.50 $\pm$ 35.45 1173.43 $\pm$ 198.41 Inc-2 6.41 $\pm$ 1.43 36.82 $\pm$ 11.82 220.47 $\pm$ 55.22 128.32 $\pm$ 14.86 1020.04 $\pm$ 131.15 Inc-3 8.30 $\pm$ 0.78 48.24 $\pm$ 18.26 169.79 $\pm$ 41.60 169.78 $\pm$ 37.03 1183.44 $\pm$ 217.72 Mix-1 8.83 $\pm$ 1.18 51.25 $\pm$ 13.72 205.54 $\pm$ 25.51 210.50 $\pm$ 11.62 1313.43 $\pm$ 132.16 Mix-2 7.42 $\pm$ 1.37 44.34 $\pm$ 13.40 203.46 $\pm$ 48.75 193.17 $\pm$ 44.83 1676.92 $\pm$ 398.90 HOPSE-G Adj-1 5.09 $\pm$ 0.93 26.71 $\pm$ 4.79 88.31 $\pm$ 11.27 95.95 $\pm$ 28.16 683.86 $\pm$ 125.70 Adj-2 4.98 $\pm$ 0.47 31.85 $\pm$ 5.43 96.33 $\pm$ 19.91 93.98 $\pm$ 15.00 887.83 $\pm$ 160.74 Adj-3 6.64 $\pm$ 0.89 31.13 $\pm$ 5.56 100.65 $\pm$ 14.14 113.39 $\pm$ 18.75 595.39 $\pm$ 107.46 Inc-1 5.75 $\pm$ 1.45 34.11 $\pm$ 7.89 110.19 $\pm$ 19.33 105.67 $\pm$ 27.00 767.42 $\pm$ 85.39 Inc-2 7.65 $\pm$ 2.32 24.65 $\pm$ 6.74 105.02 $\pm$ 28.61 126.14 $\pm$ 37.05 721.57 $\pm$ 60.10 Inc-3 7.87 $\pm$ 1.47 44.42 $\pm$ 9.81 181.18 $\pm$ 96.45 100.22 $\pm$ 11.50 896.86 $\pm$ 162.01 Mix-1 7.61 $\pm$ 0.68 46.14 $\pm$ 10.02 239.59 $\pm$ 38.24 168.35 $\pm$ 50.01 671.40 $\pm$ 948.29 Mix-2 5.27 $\pm$ 0.57 36.29 $\pm$ 13.05 152.73 $\pm$ 8.26 104.12 $\pm$ 37.76 1229.56 $\pm$ 527.27
Optimal configurations.
For simpler graph classification tasks, simpler neighborhood configurations may offer the best efficiency. In contrast, more expressive neighborhood sets often provide better efficiency-accuracy trade-offs for complex topological tasks and larger datasets. These findings demonstrate that the choice of neighborhood configuration represents a critical design decision not just for model accuracy, but also for computational efficiency. The provided analysis presents a principled approach to neighborhood selection based on task requirements and domain characteristics, establishing an essential foundation for applying TDL to new fields.
F.5.3 Preprocessing time per neighborhood
Preprocessing time analysis
The computational costs of extracting positional and structural encodings are summarized in Tables 12 and 13. While preprocessing times scale with the complexity of the selected neighborhoods and not only with the number of cells, two points need to be emphasized:
Table 12: Preprocessing time per neighborhood on simplicial. Results are shown as mean time in seconds. The best result per domain and model is in bold. All Topological Tasks have the same time as they share the same preprocessing time. Times increase with neighborhood since with Adj-1 being the fastest as it’s the neighborhood of the underlying graph.
Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 2.42 35.15 53.37 53.82 142.01 333.65 333.65 333.65 Adj-2 3.24 49.49 72.94 72.80 195.71 823.68 823.68 823.68 Adj-3 3.26 58.49 74.16 73.40 199.37 973.49 973.49 973.49 Inc-1 3.62 79.37 92.40 92.20 229.92 1121.78 1121.78 1121.78 Inc-2 4.39 80.91 111.65 111.49 280.18 1544.42 1544.42 1544.42 Inc-3 5.58 123.97 147.31 148.10 367.95 2233.27 2233.27 2233.27 Mix-1 7.27 165.71 191.94 193.45 489.55 3724.43 3724.43 3724.43 Mix-2 4.99 95.34 118.46 117.35 318.12 2495.14 2495.14 2495.14 HOPSE-G Adj-1 4.54 45.94 103.91 103.47 297.61 913.38 913.38 913.38 Adj-2 7.81 65.54 172.20 172.97 496.73 1676.79 1676.79 1676.79 Adj-3 7.76 92.15 174.23 174.49 502.85 2742.26 2742.26 2742.26 Inc-1 8.05 93.32 181.82 183.44 532.93 2654.36 2654.36 2654.36 Inc-2 11.21 90.87 245.87 248.52 714.05 2813.64 2813.64 2813.64 Inc-3 14.53 142.22 324.62 326.91 952.95 4432.76 4432.76 4432.76 Mix-1 21.02 210.41 474.58 472.70 1381.15 7329.51 7329.51 7329.51 Mix-2 14.28 160.55 314.86 319.78 932.91 5241.30 5241.30 5241.30
Table 13: Preprocessing time per neighborhood on cell. Results are shown as mean time in seconds. The best result per domain and model is in bold. Similarly to the case of simplicial, the fastest neighborhoods to precompute are the smallest ones.
Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 2.81 27.16 59.68 60.07 163.66 Adj-2 3.75 40.70 81.96 82.60 221.18 Adj-3 4.44 50.97 96.94 97.38 267.01 Inc-1 4.97 74.99 120.83 121.33 310.83 Inc-2 4.98 75.22 121.62 122.38 312.12 Inc-3 7.07 116.00 182.49 179.43 461.41 Mix-1 9.66 162.30 242.08 244.40 627.80 Mix-2 7.04 89.45 159.44 160.62 441.84 HOPSE-G Adj-1 5.08 38.04 113.21 108.81 315.07 Adj-2 8.34 58.50 189.71 184.54 525.01 Adj-3 11.10 75.56 253.63 248.51 727.66 Inc-1 11.89 85.26 264.02 259.75 734.69 Inc-2 11.81 120.32 292.64 257.43 740.39 Inc-3 18.51 132.25 420.45 403.67 1163.78 Mix-1 28.02 188.99 641.07 621.33 1776.32 Mix-2 20.47 140.60 472.04 460.79 1326.40
  1. 1.
    One time preprocessing: Encodings are precomputed once per dataset, enabling efficient experimentation. For instance, preprocessing all neighborhoods for ZINC in the cell domain takes 25–30 minutes (e.g., 1,776 seconds for Mix-1 in HOPSE-G), but this cost is amortized over all subsequent training runs. The practitioners can freely test configurations like Adj-3 or Inc-2 without recomputing features, streamlining hyperparameter optimization. Finally, the precomputed encodings can be shared, reducing redundant computation.
  2. 2.
    Parallel computation: The reported precomputation time results can be easily optimized by parallel precomputation.
F.6 Runtime vs accuracy
Figure 5 provides a comparative overview of model performance in terms of F1 score against total training runtime for four benchmark tasks: $\beta_{1}$ (a), $\beta_{2}$ (b), NAME (c), and ORIENT (d). Figure 6 shows the same arrangement but from the perspective of average runtime per epoch against performance. Additionally, Figure 6 shows the convergence speed of any particular model. These results clearly show that HOPSE is the best model for an optimal trade-off between runtime and accuracy.
[…]
Appendix G Licenses
This work does not distribute existing datasets. However, the dataset it uses will be mentioned here along with their licenses. The Graph datasets MUTAG, PROTEINS, NCI1, NCI109 do not specify a license and are distributed as part of ’TUDatasets’ accessed via https://pubs.acs.org/doi/full/10.1021/acs.jcim.5b00559. The ZINC dataset from https://pubs.acs.org/doi/full/10.1021/acs.jcim.5b00559 is freely available. The MANTRA benchmarking dataset is licensed under the 3-Clause BSD license. The CSL dataset does not specify a particular license aside from the CC BY 4.0 from [17].
Appendix H Author contributions.
M.C. led the implementation, developed the theoretical framework, performed data analysis, and drafted the manuscript. M.M. primarily contributed to the theoretical framework and manuscript writing. N.M. provided infrastructure and project support, and contributed to writing. L.T. and G.B. conceived the research idea, designed the study, analyzed the results, and contributed to manuscript preparation. Additionally, L.T. participated in the implementation and theoretical development, conducted experiments with M.C., and supervised the project.

### v2

@@
 While Graph Neural Networks (GNNs) have proven highly effective at modeling relational data, pairwise connections cannot fully capture multi-way relationships naturally present in complex real-world systems.
-In response to this, Topological Deep Learning (TDL) leverages more general combinatorial representations –such as simplicial or cellular complexes– to accommodate higher-order interactions.
-Existing TDL methods often extend GNNs through Higher-Order Message Passing (HOMP), but face critical scalability challenges due to (i) a combinatorial explosion of message-passing routes, and (ii) significant complexity overhead from the propagation mechanism.
-To overcome these limitations, we propose HOPSE (Higher-Order Positional and Structural Encoder)—a message passing-free framework that uses Hasse graph decompositions to derive efficient and expressive encodings over arbitrary higher-order domains.
-Notably, HOPSE scales linearly with dataset size while preserving expressive power and permutation equivariance.
-Experiments on molecular, expressivity and topological benchmarks show that HOPSE matches or surpasses state-of-the-art performance while achieving up to 7× speedups over HOMP-based models, opening a new path for scalable TDL.
+In response to this, Topological Deep Learning (TDL) leverages more general combinatorial representations —such as simplicial or cellular complexes— to accommodate higher-order interactions.
+Existing TDL methods often extend GNNs through Higher-Order Message Passing (HOMP), but face critical scalability challenges due to (i) a combinatorial explosion of message-passing routes, and (ii) significant complexity overhead from the propagation mechanism.This work presents HOPSE (Higher-Order Positional and Structural Encoder), an alternative method to solve tasks involving higher-order interactions without message passing.
+Instead, HOPSE breaks apart arbitrary higher-order domains into it’s neighborhood relationships using a Hasse graph decompositions.
+This method shows that decoupling the representation learning of the topology of each neighborhood and the attributes results in a lower computational complexity, putting into question the need for HOMP.The experiments on molecular graph tasks and topological benchmarks show that HOPSE matches performance in traditional TDL datasets and outperforms HOMP methods in topological tasks, all while achieving up to $7\times$ speedups over HOMP-based models, opening a new path for scalable TDL.
 1 Introduction
-Graph neural networks (GNNs) [46, 11] have demonstrated impressive capabilities in modeling relational data structured as graphs.
+Graph neural networks (GNNs) (Zhou et al., 2020; Bronstein et al., 2021) have demonstrated impressive capabilities in modeling relational data structured as graphs.
 However, their reliance on edges to encode interactions inherently limits them to pairwise relationships, which may not fully capture the complexity of many real-world systems.
 For instance, social dynamics often involve group interactions, while physical phenomena like electrostatic interactions in proteins can span multiple atoms simultaneously.
-To address such limitations, Topological Deep Learning (TDL) [8, 23] introduces a framework for modeling multi-way interactions using higher-order topological domains such as hypergraphs, simplicial, cell, and combinatorial complexes.
-Within TDL, standard Topological Neural Networks (TNNs) [18, 12, 22, 21, 38] extend the graph-based message passing paradigm through Higher-Order Message Passing (HOMP), enabling information flow across all of the dimensions of the considered combinatorial representation.
+To address such limitations, Topological Deep Learning (TDL) (Bodnar, 2023; Hajij et al., 2023) introduces a framework for modeling multi-way interactions using higher-order topological domains such as hypergraphs, simplicial, cell, and combinatorial complexes.
+Within TDL, standard Topological Neural Networks (TNNs) (Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; 2022; Papillon et al., 2023) extend the graph-based message passing paradigm through Higher-Order Message Passing (HOMP), enabling information flow across all of the dimensions of the considered combinatorial representation.
 [^1] The terms higher-order topological domain and combinatorial representation are used intercheangeably: please see Section A.1 for the formal definitions.
-This generalization has shown strong performance across diverse fields—including social networks [30], protein biology [28], physics-informed learning [45], and computer networks [6, 5]—where interactions naturally extend beyond pairwise connections.
+This generalization has shown strong performance across diverse fields including social networks (Knoke and Yang, 2019), protein biology (Jha et al., 2022), physics-informed learning (Wei and Fink, 2024), and computer networks (Bernárdez et al., 2025; 2023) where interactions naturally extend beyond pairwise connections.
 However, despite their theoretical advantages, TNNs face substantial challenges stemming from the HOMP paradigm.
-Unlike GNNs, TNNs’ ability to propagate information through the full hierarchy of a higher-order domain –where elements of different ranks (e.g., nodes, edges, faces, etc.) are interconnected– leads to an explosion in the number of possible sequences of message passing routes across these different levels [42, 33, 37].
-More specifically, for a combinatorial complex $\mathcal{T}$ with dimension –i.e.
-maximum rank– $R$, a total of $(R+1)!$ message passing routes arises just considering communication flows between different ranks, and that number skyrockets to $(R+1)!\cdot(R+1)^{R}$ when allowing same-rank connections (please see Appendix B for a detailed analysis).
-On top of this vast architectural search space, TNNs operating under the HOMP paradigm are also known to be computationally demanding, message propagation often requiring significant processing resources even for moderately sized inputs [37, 20].
-This combination of combinatorial design complexity and high computational cost severely limits existing TNNs’ scalability, posing a major barrier to their deployment in large-scale or time-sensitive real-world applications.
-In response to these challenges, this work introduces Higher-Order Positional and Structural Encoder (HOPSE), a novel framework designed to overcome the computational hurdles associated with HOMP.
-Specifically, HOPSE main features and contributions include:
-  • Higher-order topological encodings: By exploiting Hasse graph decompositions –i.e.
-transformations of combinatorial representations into simpler graph ensembles [23, 38]–, HOPSE facilitates the efficient use of established graph positional and structural encoding techniques (e.g.
-[13]) to obtain analog representations for higher-order topological domains.
-  • General message passing-free method: Inspired by SaNN model [20], the proposed encoding-based methodology effectively captures relational information without explicit message passing mechanisms.
-Moreover, HOPSE is no longer constrained to simplicial complexes; by design, it can be applied to any higher-order combinatorial representations.
-  • Scalability and computational efficiency: In contrast to HOMP-based networks, HOPSE complexity scales linearly with respect to the number of elements in the combinatorial representation (Section 4.3).
-This is achieved by computing the encodings as a preprocessing step, significantly accelerating both training and inference (see Section 5 for experimental details).
-  • Theoretical guarantees: Even without message passing, HOPSE is as expressive as existing state-of-the-art HOMP-based models (e.g.
-Papillon et al.
-[38]) in distinguishing non-isomorphic complexes (Theorem 4.1), while preserving cell permutation equivariance (Theorem 4.2).
-  • Flexible instantiations: Two HOPSE realizations are presented—HOPSE-M, relying on handcrafted graph-theoretical encodings, and HOPSE-G, using a pretrained GPSE encoder—offering complementary trade-offs between preprocessing cost and representational richness (Section 4.3).
-  • Extensive empirical validation: Comprehensive experiments on molecular benchmarks [34], CLS expressivity dataset [36], and MANTRA [2] –recent dataset consisting of pure topological tasks– validate the practical effectiveness of HOPSE, showcasing its ability to efficiently capture complex relational structures and outperform existing baselines –while being up to 7×7$\times$7 × faster than HOMP-based models (Section 5).
-Additionally, Appendix B provides, for the first time, a detailed mathematical analysis of the combinatorial explosion of message passing routes in higher-order topological domains.
+Unlike GNNs, TNNs’ ability to propagate information through the full hierarchy of a higher-order domain, where elements of different ranks (e.g., nodes, edges, faces, etc.) are interconnected, leads to an explosion in the number of possible sequences of message passing routes across these different levels (Telyatnikov et al., 2024; Montagna et al., 2024; Papamarkou et al., 2024).
+More specifically, for a combinatorial complex $\mathcal{T}$ with dimension —i.e.
+maximum rank— $R$, a total of $(R+1)!$ message passing routes arises just considering communication flows between different ranks, and that number skyrockets to $(R+1)!\cdot(R+1)^{R}$ when allowing same-rank connections (please see Appendix˜C for a detailed analysis).
+On top of this vast architectural search space, TNNs operating under the HOMP paradigm are also known to be computationally demanding, message propagation often requiring significant processing resources even for moderately sized inputs (Papamarkou et al., 2024; Gurugubelli and Chepuri, 2024).
+This combination of combinatorial design complexity and high computational cost severely limits existing TNNs’ scalability and is not required for many tasks, even ones that require leveraging higher-order relationship.
+Appendix˜B contains theoretical motivation for this claim.
+In response to these challenges, this work introduces the Higher-Order Positional and Structural Encoder (HOPSE), a novel and scalable framework that bypasses the HOMP paradigm altogether.
+While the message-passing preprocessing philosophy has been successfully applied within the simplicial domain (SaNN (Gurugubelli and Chepuri, 2024)), HOPSE does not rely on it altogether and achieves a critical generalization to arbitrary higher-order domains.
+By employing the Hasse graph decomposition as a methodological bridge, this framework enables the rich literature of powerful graph-based positional and structural encodings (PSEs) to be systematically and flexibly applied to any higher-order structure.
+More specifically, HOPSE’s main features and contributions include:
+  • Higher-order topological characterization: By decomposing a combinatorial complex into simpler graph ensembles, called Hasse graphs  (Hajij et al., 2023; Papillon et al., 2023), HOPSE efficiently uses established graph PSE techniques (e.g.
+(Cantürk et al., 2023)) to obtain characterizations for higher-order topological domains.
+  • Scalability and computational efficiency: In contrast to HOMP networks, HOPSE complexity scales linearly with respect to the number of elements in the combinatorial representation (Section˜4.2).
+This is achieved by computing the encodings as a preprocessing step, significantly accelerating both training and inference (see Section˜5 for experimental details).
+  • Extensive empirical validation: Comprehensive experiments on graph  (Morris et al., 2020) and topological  (Ballester et al., 2025) datasets validate the practical effectiveness of HOPSE.
+It is able to perform on-par to HOMP methods on graph tasks while showing stronger characterization capabilities required for topological tasks with a speed up of up to $7\times$ with respect to HOMP models (Section˜5).
+Additionally, Appendix˜C provides, for the first time, a detailed mathematical analysis of the combinatorial explosion of message passing routes in higher-order topological domains.
 2 Background
 This section introduces key concepts underpinning TDL.
-For brevity and generality, the following discussion centers on combinatorial complexes, which subsume all of the discrete topological domains leveraged by TDL [23].
+For brevity and generality, the following discussion centers on combinatorial complexes (CCs), which subsume all of the discrete topological domains leveraged by TDL Hajij et al.
+(2023).
 Appendix A provides rigorous mathematical definitions of these other domains, such as simplicial and cellular complexes, along with an extended theoretical background.
@@
 The cell $\sigma\in\mathcal{C}$ of rank $k:=\textrm{rk}(\sigma)$ is referred to as a $k$-cell.
-$\mathcal{C}$ simplifies notation for $(\mathcal{V},\mathcal{C},\textrm{rk})$.
 Lastly, the dimension of $\mathcal{T}$} is defined as the maximal rank among its cells: $\mathrm{dim}(\mathcal{C}):=\max_{\sigma\in\mathcal{C}}\textrm{rk}(\sigma)$.
-Signals over combinatorial complexes.
-A combinatorial complex $\mathcal{T}=(\mathcal{V},\mathcal{C},\mathrm{rk})$ can be equipped with signals, generalizing the concept of featured graph [^2] Attributed graph and featured graph are used interchangeably to comply with the literature.
-(see Definition A.1) to higher-order structures.
-In this context, a signal is a set of functions =$\mathcal{F}=\{F_{r}\}_{r\geq 0}$, where $F_{r}:\{\sigma\in\mathcal{C}\mid\mathrm{rk}(\sigma)=r\}\to\mathbb{R}^{D_{r}}$, where $D_{r}$ is the dimension of the feature space for rank-$r$ cells.
-For this work, it is important to emphasize that a complex can have multiple attribute functions over each $r$-cell.
-More formally, =$\mathcal{F}=\{F_{r,k}\}_{k=1}^{K}$ for all $r\geq 0$, where each $F{r,k}$ shares the same domain: $\text{Dom}(F_{r,k})=\text{Dom}(F_{r})$.
 Neighborhood relations: adjacencies and incidences.
-Combinatorial complexes can be equipped with a notion of neighborhood among cells that induces rich topological structures.
-Formally, denoting by $\mathcal{P}(\cdot)$ the power set, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a complex ${\mathcal{T}}$ maps each cell $\sigma\in\mathcal{C}$ to a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$.
-Traditional neighborhood functions include adjacencies ($\mathcal{A}_{t,s}$), connecting cells of the same rank $t$ –through adjacency relations w.r.t.
-cells of another rank $s$–, and incidences ($\mathcal{I}_{s\to t}$), linking cells of different ranks –from a source rank $s$ to a target rank $t$.
-Figure 1 presents some instances, and Section A.2 provides formal definitions.
+Combinatorial complexes can be equipped with a notion of neighborhood among cells that induces topological structures.
+Formally, denoting by $\mathcal{P}(\cdot)$ the power set, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a complex $\mathcal{T}$} maps each cell $\sigma\in\mathcal{C}$ to a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}$.
+Traditional neighborhood functionsme include adjacencies ($\mathcal{A}_{t,s}$), connecting cells of the same rank $t$ —through adjacency relations w.r.t.
+cells of another rank ss—, and incidences ($\mathcal{I}_{s\to t}$), linking cells of different ranks from a source rank $s$ to a target rank $t$.
+Figure˜1 presents some instances, and Section˜A.2 provides formal definitions.
 Figure 1: Neighborhood expansions of a complex.
-Given a complex ${\mathcal{T}}$ (left), three examples of strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}}$ are illustrated corresponding to 4 neighborhoods: adjacency of nodes w.r.t.
+Given a complex $\mathcal{T}$} (left), three examples of strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}}$ are illustrated corresponding to 4 neighborhood functions: adjacency of nodes w.r.t.
 edges (${\mathcal{A}}_{0,1}$), incidence from nodes to edges (${\mathcal{I}}_{0\to 1}$), incidence from faces to nodes (${\mathcal{I}}_{2\to 1}$), and incidence from nodes to faces (${\mathcal{I}}_{0\to 2}$).
 Strictly augmented Hasse graphs.
-Given a complex ${\mathcal{T}}$, each particular neighborhood $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ [39], defined as the directed graph whose nodes and edges are given, respectively, by
+Given a complex $\mathcal{T}$}, each particular neighborhood function $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ (Papillon et al., 2024), defined as the directed graph whose nodes and edges are given, respectively, by
 $\mathcal{C}_{{\mathcal{N}}}=\{\sigma\in{\mathcal{C}}\mid{\mathcal{N}}(\sigma)\neq\emptyset\}\quad\text{and}\quad\mathcal{E}_{{\mathcal{N}}}=\{(\tau,\sigma)\mid\sigma,\tau\in{\mathcal{C}}_{{\mathcal{N}}},\ \tau\in{\mathcal{N}}(\sigma)\}.$
-Figure 1 visually shows some examples of these neighborhood graph expansions.
-Notably, “strictly” comes from the fact that the set of nodes of $\mathcal{G}_{{\mathcal{N}}}$ is restricted to $\mathcal{C}_{{\mathcal{N}}}$, i.e., to those cells that have at least one neighbor according to ${\mathcal{N}}$.
-In contrast to this, the original definition of augmented Hasse graph considers the full set of cells $\mathcal{C}$ [23].
+Figure˜1 visually shows some examples of the graph expansions induced by these neighborhood functions.
+Notably, strictly comes from the fact that the set of nodes of $\mathcal{G}_{{\mathcal{N}}}$ is restricted to $\mathcal{C}_{{\mathcal{N}}}$, i.e., to those cells that have at least one neighbor according to $\mathcal{N}$}.
+In contrast to this, the original definition of augmented Hasse graph considers the full set of cells $\mathcal{C}$ (Hajij et al., 2023).
+Rank-targeted neighborhood functions.
+The rank-targeted neighborhood is defined as the collection of neighborhood functions whose target cells have the same fixed rank $r$: $\mathcal{N}^{\to r}:=\{\mathcal{N}\in\mathcal{N}_{S}\mid\forall\sigma\in\mathcal{C},\forall\tau\in\mathcal{N}(\sigma):\mathrm{rk}(\tau)=r\}$.
 3 Related works
-Positional and Structural Encodings in GNNs.
-In the graph domain, a range of positional and structural encodings (PSEs) have been proposed to enhance node representations.
+Graph characterization in GNNs.
+In the graph domain, a range of positional and structural encodings (PSEs) have been proposed to enhance node representations and characterize graphs.
 Positional encodings typically capture a node’s location within the graph (e.g., via distances or coordinates), while structural encodings describe the node’s role or function in the topology (e.g., node degrees).
-Classical approaches of positional encodings include spectral methods based on Laplacian eigenvectors (LapPE) [17], while a common structural encoding is the random-walk embedding (RWSE) [15, 43], which reflects multi-hop connectivity; these handcrafted features help differentiate nodes with otherwise similar local neighborhoods.
-More recently, transformer-based GNNs have incorporated PSEs into attention mechanisms [14, 40, 31].
-Other works focus on decoupling the positional information from the structure and learn the embeddings instead of picking them manually [14].
-Building upon this, the Graph Positional and Structural Encoder (GPSE) [13] offers a unified and transferable solution by learning a shared latent space of PSEs through self-supervised pretraining.
+Classical approaches of positional encodings include spectral methods based on Laplacian eigenvectors (LapPE) Dwivedi et al.
+(2023), while a common structural encoding is the random-walk embedding (RWSE) Dwivedi et al.
+(2022); Wang et al.
+(2022), which reflects multi-hop connectivity.
+These handcrafted features help characterize graphs with otherwise similar local neighborhoods.
+More recently, transformer-based GNNs have incorporated PSEs into attention mechanisms Dwivedi et al.
+(2021); Rampášek et al.
+(2022); Ma et al.
+(2023).
+Other works focus on decoupling the positional information from the structure and learn the embeddings instead of picking them manually (Dwivedi et al., 2021).
+Building upon this, the Graph Positional and Structural Encoder (GPSE) Cantürk et al.
+(2023) offers a unified and transferable solution by learning a shared latent space of PSEs through self-supervised pretraining.
 GPSE’s embeddings can be integrated into downstream GNNs, achieving performance on par with or exceeding task-specific encodings, and advancing the development of general-purpose graph models.
-Combinatorial Complex Neural Networks.
-To overcome the pairwise limitations of graphs, recent work has turned to explicit representation of higher-order interactions through topological domains.
-Combinatorial complexes provide a unified framework for modeling higher-order interactions with hierarchical organization, and their associated Combinatorial Complex Neural Networks (CCNNs) enable message passing across cells of varying rank (HOMP paradigm), naturally capturing multi-way relationships [23].
-Generalizing and subsuming CCNNs, Generalized CCNNs (GCCNs) [39] offer a more flexible and powerful framework: they systematically decompose a complex into a set of strictly augmented Hasse graphs –each one capturing a particular neighborhood– in a modular way that preserves the higher-order structure.
-In doing so, GCCNs not only facilitate the design of new TNN architectures but also help bridge the gap between graph deep learning and TDL.
-However, existing TNNs –CCNNs and GCCNs alike– suffer from the high computational complexity associated to HOMP, limiting their practical use in large real-world applications.
-Positional encodings in higher-order domains.
-Encoding positional information for higher-order topological elements poses unique challenges.
-Unlike in graphs, where positional features are typically assigned to nodes, each cell rank of a combinatorial complex –nodes, edges, faces,…– potentially requires its own descriptor.
-Recent works have begun addressing this, though this topic still remains largely unexplored.
-For instance, the Cellular Transformer (CT) [1] extends transformer architectures to cell complexes by incorporating attention mechanisms that respect higher-order incidence relations.
-To support this, the authors introduce a set of manually crafted positional encodings tailored to cellular complexes, though their results did not shed light on how to build robust encodings—different ones perform best on different tasks.
-Moreover, CT incurs quadratic complexity in the number of cells due to the attention mechanism, limiting its scalability to medium-to-large scenarios.
-In contrast, Simplicial-aware Neural Networks (SaNN[20] take a more efficient route by bypassing message passing entirely.
-SaNN precomputes structural and positional summaries (e.g., over lower, upper, and co-neighbors) and feeds them as fixed-length vectors into a standard neural network.
-This achieves constant runtime and memory overhead while maintaining strong expressivity and performance.
-Nevertheless, the SaNN architecture is also constrained in terms of the topological domain (simplicial complex-specific), as well as with the choices of neighborhoods –which haves an impact into its theoretical expressivity and applicability.
+Efficient message passing methods
+There are many different ways to achieve expressive and computationally efficient GNNs but the same cannot be said for TNNs.
+Various works in the GNN domain have approached message passing over subgraphs or partitions of the original graph (Zeng et al., 2021; Li et al., 2021).
+Additionally, other approaches focus on decoupling learning the features from the structure (Frasca et al., 2020; Doshi and Chepuri, 2022; Chen et al., 2020).
+In higher-order domains partitions are not trivial, so other methods have been applied to achieve efficiency.Simplicial Scattering Networks  (Madhu et al., 2024), leverage random walks over the intitial simplicial complex which give node structure-informed embeddings resulting in a parameter-free network.
+However, their experimental framework is mostly focused in node-level tasks.
+To the best of our knowledge Simplicial-aware Neural Networks (SaNN(Gurugubelli and Chepuri, 2024) is the only method that proposes efficient computations for higher-order networks in a learning scenario.
+They precompute a deterministic transformation over both the structure and the features of neighborhoods in a simplicial complex.
+Effectively, this results in a precomputation of message passing.
+However, it has several drawbacks.
+First, there is no clear indication of which transformation of the features and structure of a neighborhood should be used.
+Since this is a preprocessing step, this function cannot be learned.
+Second, this approach is limited to only simplicial complexes.
+HOPSE joins the ideas from (Frasca et al., 2020; Chen et al., 2020) and (Gurugubelli and Chepuri, 2024) to partially decouple learning the structure from the attributes.
+Additionally, HOPSE bypasses the theoretical limitations of (Gurugubelli and Chepuri, 2024) (and HOMP-based models), which is at most as expressive as SWL.
 4 HOPSE: higher-order positional and structural encoder
-Section 4.1 outlines the proposed HOPSE pipeline, Section 4.2 establishes its theoretical guarantees, and Section 4.3 presents two practical realizations of the framework.
-Let ${\mathcal{T}}$ be an arbitrary combinatorial complex, and ${\mathcal{N}}_{C}\subset{\mathcal{N}}_{\mathcal{T}}$ any subset of neighborhoods on ${\mathcal{T}}$.
-The following notation is introduced to support the upcoming presentation of the framework:
-Strictly augmented multi-attributed Hasse graphs.
-For a given neighborhood ${\mathcal{N}}\in{\mathcal{N}}_{\mathcal{C}}$, its strictly augmented multi-attributed Hasse graph –denoted by ${\mathcal{J}}_{{\mathcal{N}}}$– assigns to the resulting strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$ the corresponding signals $\mathcal{F}_{{\mathcal{N}}}$ over the cells present in that neighborhood, i.e.
-${\mathcal{J}}_{{\mathcal{N}}}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},\mathcal{F}_{\mathcal{N}})$.
-The shorthand ${\mathcal{F}}_{{\mathcal{N}},k}=\{F_{r,k}\}_{r\geq 0}$ is the set of all $k$-attribute functions for each rank $r$ that act on the subset of cells in a neighborhood ${\mathcal{C}}_{\mathcal{N}}$ as $F_{r,k}:\{\sigma\in{\mathcal{C}}_{\mathcal{N}}\}\to\mathbb{R}^{D_{r}}$.
-To streamline the management of multiple associated signals, the restriction of ${\mathcal{J}}_{{\mathcal{N}}}$ to the $k$-th feature is defined as $\evaluated{{\mathcal{J}}_{{\mathcal{N}}}}_{k}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},{\mathcal{F}}_{{\mathcal{N}},k})$.
-Lastly, $\mathbf{J}_{C}=\left\{{\mathcal{J}}_{{\mathcal{N}}}\right\}_{{\mathcal{N}}\in{\mathcal{N}}_{C}}$ denotes the corresponding Hasse graph decomposition of ${\mathcal{T}}$ w.r.t.
-a given neighborhood set ${\mathcal{N}}_{C}$.
-Rank-targeted neighborhoods.
-The rank-targeted neighborhood is defined as the collection of neighborhood functions whose target cells have the same fixed rank $r$: $\mathcal{N}^{\to r}:=\{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}\mid\forall\sigma\in\mathcal{C},\forall\tau\in\mathcal{N}(\sigma):\mathrm{rk}(\tau)=r\}$.
+Section˜4.1 outlines the proposed HOPSE pipelineand Section˜4.2 presents two practical realizations of the framework.
 4.1 Methodology
-This section describes HOPSE, a method for enriching cell features in combinatorial complexes with higher-order positional and structural encodings derived from Hasse graph decompositions.
-HOPSE enables expressive higher-order modeling without relying on the computational overhead of HOMP, while retaining its representational power.
-The full pipeline is illustrated in Figure 2, with components referenced using the corresponding labels (A–F) throughout this section.
+This section describes HOPSE, a method for learning representations of (CCs) through the caracterization of it’s pieces.
+Characterizing a CC amounts to determining properties or invariants.
+These are shared among equivalnt CCs and allow distinguishing non-equivalent CCs from each toher, however similar they may be.
+The Hasse decomposition facilitates this characterization as a CC is determined by it’s Hasse decomposition.
+This decomposition results in a set of graphs that determines the initial object.
+For the theoretical motivation refer to Appendix˜B.
+This work leverages the previous theoretical insights and the extensive research in graph characterization, for example through PSEs, to employ Hasse graphs as proxy objects to characterize a CC and model higher-order interactions.
+In the spirit of (Frasca et al., 2020), HOPSE partially decouples the structural from the functional representation.
+The structure is given by the Hasse decomposition while the functional representation is given by the attributes over the complex.
+Learning both structure and features in parallel circumvents the high cost of mixing features through neighborhood relationships, the pillar behind HOMP.
+The full pipeline is illustrated in Figure˜2, with components referenced using the corresponding labels (A–F) throughout this section.
 Figure 2: HOPSE pipeline.
-(A) Considering the collection of neighborhoods ${\mathcal{N}}_{C}=\{{\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{0\to 2}\}$, the input combinatorial complex $\mathcal{T}$ is decomposed into the corresponding multi-attributed Hasse graphs expansions $\{\mathcal{J}_{{\mathcal{N}}}\}_{{\mathcal{N}}_{C}}$ –each one with $k$ associated signals.
-(B) Positional and structural encodings are computed for each Hasse graph, $\mathbf{G}(\mathcal{J}_{\mathcal{N}_{i}})$, as defined in Equation 1.
-(C) Higher-order PS encodings are derived via the target rank-aware aggregation $\psi_{k}$ (see Equation 2).
-(D) The PS encodings $\mathbf{X}_{r,k}$ are transformed using $f_{r,k}(\cdot)$, and initial cell features $\mathbf{Z}_{r}$ are embedded via $\varepsilon_{r}(\cdot)$, following Equations 3 and 4.
-(E) Rank-specific, neighborhood-aware representations $\mathbf{H}_{r}$ are learned using $\Theta_{r}(\cdot)$, as described in Equation 5.
-(F) The final task-specific readout $\varphi(\cdot)$ integrates all $\mathbf{H}_{r}$ representations (Equation 6).
+(A) Considering the collection of neighborhood functions ${\mathcal{N}}_{C}=\{{\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{0\to 2}\}$, the input combinatorial complex $\mathcal{T}$ is decomposed into the corresponding Hasse graphs expansions $\{\mathcal{G}_{{\mathcal{N}}}\}_{{\mathcal{N}}_{C}}$.
+(B) PSEs are computed for each Hasse graph, $\mathbf{G}(\mathcal{G}_{\mathcal{N}_{i}})$, as defined in Equation˜1.
+(C) Higher-order PSEs are derived via the target rank-aware aggregation $\psi_{k}$ (see Equation˜2).
+(D) The PSEs $\mathbf{X}_{r,k}$ are transformed using $f_{r,k}(\cdot)$, and initial cell features $\mathbf{Z}_{r}$ are embedded via $\varepsilon_{r}(\cdot)$, following Equations˜3 and 4.
+(E) Rank-specific, structure-aware representations $\mathbf{H}_{r}$ are learned using $\Theta_{r}(\cdot)$, as described in Equation˜5.
+(F) The final task-specific readout $\varphi(\cdot)$ integrates all $\mathbf{H}_{r}$ representations (Equation˜6).
 Extract positional and structural information.
-Let ${\mathcal{N}}_{C}$ be a collection of neighborhood functions.
-For each neighborhood ${\mathcal{N}}\in{\mathcal{N}}_{C}$, the Hasse graph decomposition of ${\mathcal{T}}$ enables the treatment of every ${\mathcal{J}}_{{\mathcal{N}}}$ as an independent component (see Figure 2A), each representing a unique set of connectivity relationships.
+Let ${\mathcal{N}}_{S}$ be a collection of neighborhood functions.
+For each neighborhood function ${\mathcal{N}}\in{\mathcal{N}}_{S}$, the Hasse graph decomposition of $\mathcal{T}$} enables the treatment of every ${\mathcal{G}}_{{\mathcal{N}}}$ as an independent component (see Figure˜2A), each representing a unique set of connectivity relationships.
 $\mathbf{G}$ is a set of Positional and Structural Encoders (PSEs) such that
-$\mathbf{G}\colon\{1,\dots,K\}\times\mathbf{J}_{\mathcal{C}}\longrightarrow\mathbb{R}^{N\times D_{k}},\quad\left(k,{\mathcal{J}}_{{\mathcal{N}}}\right)\longmapsto g_{k}\left({{\mathcal{J}}_{{\mathcal{N}}}\bigr{|}}_{k}\right).$ (1)
-Here $g_{k}(\cdot):\mathbf{J}_{\mathcal{C}}\to\mathbb{R}^{|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}\times D_{k}}$ denotes a particular PSE, with $|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}$ being the number of nodes in the Hasse graph induced by neighborhood ${\mathcal{N}}$ corresponding to the target cells (illustrated in Figure 2B).
+$\mathbf{G}\colon\{1,\dots,K\}\times\mathcal{G}_{S}\longrightarrow\mathbb{R}^{N\times D_{k}},\quad\left(k,\mathcal{G}_{\mathcal{N}}\right)\longmapsto g_{k}\left(\mathcal{G}_{\mathcal{N}}\right).$ (1)
+Here $g_{k}(\cdot):\mathcal{G}_{S}\to\mathbb{R}^{|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}\times D_{k}}$ denotes a particular PSE, with $|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}$ being the number of nodes in the Hasse graph constructed from the neighborhood function $\mathcal{N}$} corresponding to the target cells (illustrated in Figure˜2B).
 Note that each $g_{k}(\cdot)$ maps to a different $D_{k}$, allowing PSEs to have different dimensions.
 Aggregate positional and structural encodings across neighborhoods.
-The proposed neighborhood aggregation mechanism $\psi_{k}(\cdot)$ operates by grouping the encodings of cells with the same target rank $r$ to produce a unique embedding for each cell and each $k\in\{1,\dots,K\}$ (see Figure 2C).
+The function $\psi_{k}(\cdot)$ groups encodings of cells with the same target rank $r$ to produce a embedding that characterizes each cell and for each $k\in\{1,\dots,K\}$ (see Figure˜2C).
 The embeddings for cells of rank $r$ are computed as:
-$\mathbf{X}_{r,k}=\psi_{k}(\{g_{k}\left({{\mathcal{J}}_{{\mathcal{N}}}\bigr{|}}_{k}\right)\mid{\mathcal{N}}\in{\mathcal{N}}^{\to r}\}),$ (2)
+$\mathbf{X}_{r,k}=\psi_{k}(\{\!\!\{g_{k}(\mathcal{G}_{\mathcal{N}})\mid{\mathcal{N}}\in{\mathcal{N}}^{\to r}\}\!\!\}),$ (2)
 where $\mathbf{X}_{r,k}\in\mathbb{R}^{N_{r}\times D_{k^{\prime}}}$, with $N_{r}$ denoting the number of $r$-cells and $D_{k^{\prime}}$ representing the output feature dimension for attribute $k$.
-The functions $\psi_{k}(\cdot)$ can also be interpreted as rank-level aggregators, where the aggregation operation can be any standard reduction method.
-Common choices include concatenation, sum, average, or even learnable transformations such as MLPs.
-The main advantage of using non-learnable aggregation functions for $\psi_{k}$ is that the embeddings $\mathbf{X}_{r,k}$ can be precomputed during dataset preprocessing, since they do not require learning.
-Learning embeddings of cells.
-In the HOPSE framework, each $r$-cell is associated with two complementary sources of information: (i) higher-order positional and structural encodings $\mathbf{X}_{r,k}$ aggregated from multiple neighborhoods, and (ii) initial features $\mathbf{Z}_{r}$ provided by the dataset or derived during the construction of the higher-order structure.
-Both types of features are independently projected into latent spaces through sets of learnable functions (see Figure 2D):
+The functions $\psi_{k}(\cdot)$ can also be interpreted as rank-level aggregators, where the aggregation operation can be any standard reduction method over a multiset  [^2] Common choices include concatenation, sum, average, or even learnable transformations such as MLPs..
+The main advantage of using non-learnable aggregation functions for $\psi_{k}$ is to precompute the embeddings $\mathbf{X}_{r,k}$ during dataset preprocessing.
+Learning structure and attributes.
+In the HOPSE framework, each $r$-cell is associated with two complementary sources of information: (i) neighborhood-aware PSEs $\mathbf{X}_{r,k}$ , and (ii) initial features $\mathbf{Z}_{r}$ provided by the dataset or derived during the construction of the higher-order structure.
+Both types of features are projected in parallel into latent spaces through sets of learnable functions (see Figure˜2D):
 $\hat{\mathbf{X}}_{r,k}$ $=f_{r,k}(\mathbf{X}_{r,k}),$ (3)
 $\hat{\mathbf{Z}}_{r}$ $=\varepsilon_{r}(\mathbf{Z}_{r}),$ (4)
-where $f_{r,k}(\cdot)$ is applied to each feature $k$ of the $r$-cells, and $\varepsilon_{r}(\cdot)$ to their initial features.
+where $f_{r,k}(\cdot)$ is applied to each structural feature matrix $\mathbf{X}_{r,k}$ of each $r$-cell, and $\varepsilon_{r}(\cdot)$ to their initial features.
 Learning neighborhood-aware representations.
-Next, to integrate the structural information with the cells’ initial attributes, the extracted and transformed higher-order PS encodings and the cell features are passed through the aggregator $\Theta_{r}(\cdot)$ (see Figure 2E) to obtain the final neighborhood-aware representations $\mathbf{H}_{r}\in\mathbb{R}^{N_{r}\times D_{r}}$:
+Next, to integrate the structural information with the cells’ initial attributes, $\Theta_{r}(\cdot)$ aggregates their latent representations (see Figure˜2E) to obtain the final neighborhood-aware representations $\mathbf{H}_{r}\in\mathbb{R}^{N_{r}\times D_{r}}$:
 $\mathbf{H}_{r}=\Theta_{r}(\hat{\mathbf{Z}}_{r},\hat{\mathbf{X}}_{r,1},\dots,\hat{\mathbf{X}}_{r,K}),$ (5)
@@
 Since the structural information is now contained in the cell features, they can be treated independently.
-For example, for complex-level tasks where the network needs an output for the whole complex, the global transformation $\varphi(\cdot)$ combines the aggregated embeddings from all ranks (see Figure 2F).
+For example, for complex-level tasks where the network needs an output for the whole complex, the global transformation $\varphi(\cdot)$ combines the aggregated embeddings from all ranks (see Figure˜2F).
 The overall output is
-$\text{HOPSE}({\mathcal{T}})=\varphi\Bigl{(}\mathbf{H}_{0},\mathbf{H}_{1},\ldots,\mathbf{H}_{R}\Bigr{)}.$ (6)
-4.2 Theoretical guarantees
-The expressivity of GNNs is commonly analyzed through their correspondence with the Weisfeiler–Leman (WL) isomorphism test [26].
-In the context of TDL, a generalization of WL over arbitrary neighborhoods of a combinatorial complex (CC) has been proposed, referred to as the CCWL test [39].
-The following theorem establishes the expressivity of HOPSE with respect to it.
-Theorem 4.1.
-HOPSE is more powerful than the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
-The proof idea is that although HOPSE does not rely on message passing, with appropriately chosen neighborhoods, it can compute positional encodings using the same information as CCWL.
-Given the injectivity assumptions, this implies at least equal expressivity.
-The proof concludes by showing two CCs distinguished by HOPSE but not by CCWL.
-Full details are in Appendix C.
-Another key property for models working on combinatorial data is cell permutation equivariance—the ability to yield consistent outputs under permutations of the cells.
-The following theorem establishes that HOPSE preserves this property under reasonable conditions:
-Theorem 4.2.
-HOPSE is cell permutation equivariant if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are cell permutation equivariant for all $k\in{1,\dots,K}$ and all $r\in{0,\dots,R}$.
-This follows from the fact that compositions of permutation-equivariant functions remain equivariant.
-The proof is in Appendix C.
-While this requirement may seem strong, standard MLPs—though not equivariant to feature permutations—are equivariant to permutations of input sets, making them suitable functions for $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$.
+$\text{HOPSE}({\mathcal{T}})=\varphi\Bigl(\mathbf{H}_{0},\ldots,\mathbf{H}_{R}\Bigr).$ (6)
 4.2 Realization of the framework
-The practical implementation of the proposed framework requires defining a set of neighborhoods of interest ${\mathcal{N}}_{C}=\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$ (discussed in Section 5), a set of encoding functions $\mathbf{G}$ for deriving structural and positional features, neighborhood aggregation functions $\psi_{k}(\cdot)$, rank-specific processing functions $f_{r,k}(\cdot)$, and a structure mixing function $\Theta_{r}(\cdot)$.
+The practical implementation of the proposed framework requires defining a set of $T$ neighborhood functions of interest ${\mathcal{N}}_{S}=\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$ (discussed in Section˜5), a set of encoding functions $\mathbf{G}$ for deriving structural and positional features, neighborhood aggregation functions $\psi_{k}(\cdot)$, rank-specific processing functions $f_{r,k}(\cdot)$, and a structure mixing function $\Theta_{r}(\cdot)$.
 In particular, this work adopts $\psi_{k}(\cdot)=\mathrm{CONCAT}(\cdot)$ for aggregation, enabling each positional encoding to be handled independently.
-For both $f_{r,k}(\cdot)$ and $\Theta_{r}(\cdot)$, trainable multilayer perceptrons (MLPs) are used.
-The choice of encoding functions $\mathbf{G}$ defines the two proposed instantiations of the framework, HOPSE-M and HOPSE-G, offering a trade-off regarding the informativeness of extracted PS encoding and the computation cost to generate them.
-Additional implementation details are provided in Appendix D.
+The functions $\epsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$ and $\Theta_{r}(\cdot)$ are modeled as multilayer perceptrons (MLPs).
+The choice of functions $\mathbf{G}$ defines two instantiations of the framework, HOPSE-M and HOPSE-G, offering a trade-off between representational power  [^3] Since HOPSE-G uses message passing to generate the embeddings of the Hasse graphs it is not completely message passing free..
+and computation cost.
+Additional implementation details are provided in Appendix˜D.
 HOPSE-M.
@@
 Specifically, $\mathbf{G}=\{g_{1},g_{2},g_{3},g_{4}\}$ includes: (i) eigenvector-based encodings, (ii) electrostatic potential-based encodings, (iii) random-walk-based encodings, and (iv) heat kernel signatures.
-These encodings are chosen to provide a fair comparison with HOPSE-G, as the GPSE model used in that variant was trained on the same set of positional and structural descriptors [13].
-While informative, several of these encodings—especially those requiring eigendecomposition such as (i), (ii), and (iv)—are computationally intensive, with complexity ${\mathcal{O}}(n^{3})$ with $n$ the number of graph nodes.
+Section˜F.2 contains an in depth description of the choice of PSEs.
 HOPSE-G.
-This variant leverages the pretrained GPSE network [13] to encode structural and positional information through a lightweight message-passing architecture.
+This variant leverages the pretrained GPSE network (Cantürk et al., 2023) to encode structural and positional information.
 In this realization, $\mathbf{G}=\{g_{1}\}$ with $g_{1}=\textsc{GPSE}(\cdot)$.
 Unlike manual encodings that rely on computationally intensive operations such as eigendecomposition, GPSE produces latent descriptors using standard message passing.
-This makes HOPSE-G particularly suitable for large combinatorial complexes with many cells, offering a scalable alternative without significant loss in expressiveness.
+Domain specific tasks can thus benefit from models pretrained in that domain.
+Choice of ${\mathcal{N}}_{S}$.
+The choice of the set of neighborhood functions ${\mathcal{N}}_{S}$ used by HOPSE is a critical design choice in the HOPSE framework.
+Following the theoretical motivation (Appendix˜B) extensive ablation studies show the empirical performance (Appendix˜H) and runtime behavior (Appendix˜I) of the models.
 Computational complexity.
-The presented implementation of HOPSE is designed for computational efficiency: no training is required up to and including Equation 2, allowing the matrices $\mathbf{X}_{r,k}$ to be fully precomputed in a preprocessing phase.
-By offloading all structure-dependent computations to preprocessing, HOPSE has complexity that scales linearly with the number of cells while retaining high expressive power, as established in Theorem 4.1.
-The efficiency of this design is further validated by experimental runtime results in Table 3.
-A comprehensive complexity analysis and a comparison with GCCN—a highly general higher-order architecture—are provided in Appendix E.
+The HOPSE framework is efficient.
+No training is required up to and including Equation˜2, allowing the matrices $\mathbf{X}_{r,k}$ to be fully precomputed in a preprocessing phase.
+By decoupling the learning procedure of features and structure, HOPSE has complexity that scales linearly with the number of cells.
 5 Experiments
-This section empirically evaluates the proposed HOPSE framework, benchmarked against state-of-the-art TNNs.
-Both instantiations of the framework, HOPSE-M and HOPSE-G (see Section 4.3), are assessed on graph molecular datasets, as well as on expressivity and topological tasks.
-The standardized experimental protocol introduced in TopoBench [42] is adopted to ensure fair and reproducible comparisons through a comprehensive optimization and evaluation suite.
+This section empirically evaluates the HOPSE framework against state-of-the-art TNNs.
+Both instantiations of the framework, HOPSE-M and HOPSE-G (see Section 4.2), are assessed on graph molecular datasets and topological tasks.
+This work adopts the standardized experimental protocol introduced in TopoBench (Telyatnikov et al., 2024) to ensure fair and reproducible comparisons through a comprehensive optimization and evaluation suite.
 The experimental setup is carefully designed to systematically address the following research questions:
   1. Q1.
-    Can the proposed HOPSE framework and its realizations (HOPSE-M and HOPSE-G) effectively approximate the performance of TNNs that leverage HOMP on standard TDL datasets?
+    Is HOMP required for standard TDL evaluation benchmarks ?
   2. Q2.
-    Can HOPSE achieve competitive performance on synthetic datasets specifically constructed to evaluate the ability to capture and utilize topological properties?
+    Is HOMP a hard requirement for tasks related to topological properties?
   3. Q3.
-    How expressive HOPSE can get in practical scenarios compared with other GNNs and TNNs?
-  4. Q4.
-    To what extent does the reduced complexity of HOPSE, relative to HOMP-based methods, yield practical improvements in execution cost and scalability on standard TDL benchmarks?
+    Does the lower computational complexity result in a reduced training and inference time ?
 5.1 Evaluation setup
 Datasets.
-First, the framework is evaluated on five standard benchmarking datasets commonly used in the TDL literature [42], comprising four graph classification tasks and one graph regression task.
-The classification tasks use datasets from the TUDataset collection [^3] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [35].
-For regression, the ZINC dataset is used [27].
-Following standard practice in TDL, structural lifting is applied to lift these graph datasets into higher-order domains—specifically, cycle-based lifting for the cell domain and clique complex lifting for the simplicial domain.
-Next, MANTRA family [2] represents purely topological datasets and tasks.
+First, the framework is evaluated on five standard benchmarking datasets commonly used in the TDL literature (Telyatnikov et al., 2024), comprising four graph classification tasks and one graph regression task.
+The classification tasks use datasets from the TUDataset collection [^4] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluation.MUTAG, PROTEINS, NCI1, and NCI109 (Morris et al., 2020), and the regression is ZINC-12K (Irwin et al., 2012).
+Following standard practice in TDL, structural lifting is applied to transform these graph datasets into higher-order domains—specifically, cycle-based lifting for the cell domain and clique complex lifting for the simplicial domain.
+Next, the MANTRA set of datasets (Ballester et al., 2025) represents purely topological datasets and tasks.
 It comprises over 43,000 triangulations of 2-manifolds.
-Among the multiple classification tasks that MANTRA datasets offer, this analysis focuses on the following three: predicting the homeomorphism name, predicting the orientability of a triangulated surface, and assigning a numerical label to the first three Betti numbers $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$.
-Lastly, the evaluation includes the CLS dataset [36], used to assess the expressivity of GNNs [16], which contains 150 graphs across 10 isomorphism classes.
-For further details, see Appendix F.
+Among the multiple classification tasks that MANTRA datasets offer, this analysis focuses on the following three: predicting the homeomorphism type (NAME), predicting the orientability of a triangulated surface (ORIENT), and assigning a numerical label to the first three Betti numbers $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$.
+For further discussion of the chosen benchmarks refer to Section˜F.2.
 Models.
-The SaNN architecture is re-implemented due to unavailable source code [19] and included in all benchmarks.
-For the MANTRA family [2], baseline results are reported for the first time under the TopoBench protocol, ensuring consistency with the broader evaluation framework.
-Full details of the hyperparameter search and selected configurations are provided in Appendix F.2.
+The benchmark includes several models from the simplicial and cellular domains.
+In the simplicial domain: SCN (Yang et al., 2022), SCCNN (Yang and Isufi, 2023), and GCCN (Papillon et al., 2024).
+In the cellular domain: CCCN (Hajij et al., 2020), CWN (Bodnar et al., 2021), and GCCN.
+This work also provides a re-implementation of the SaNN architecture Gurugubelli and Chepuri (2024), as its source code was unavailable.
+Additionally, the first baseline results for the MANTRA family (Ballester et al., 2025) under the TopoBench protocol.
+Full details on the hyperparameter search and selected configurations are provided in Appendix F.3.
 Training and evaluation.
-To ensure fair comparisons, five stratified splits with a 50%/25%/25% train/validation/test ratio are considered for each dataset and task –except on ZINC, which uses predefined splits [27].
-The optimal hyperparameter configuration is selected based on the average validation performance across five random seeds.
-Evaluation metrics include classification accuracy (MUTAG, PROTEINS, NCI1, NCI109), mean absolute error (MAE; ZINC), and F1 score (MANTRA).
-For each dataset, the mean and standard deviation are computed across the five test sets and reported in Table 1.
 Publicly available results are adopted from Telyatnikov et al.
-[42] and Papillon et al.
-[39] where applicable for graph datasets.
+(2024) and Papillon et al.
+(2024) where applicable for graph datasets.
+While for the rest, the TopoBench evaluation protocol was applied, which consists of stratified splits with 50%/25%/25% train/validation/test ratios, except on ZINC-12K, which uses predefined splits (Irwin et al., 2012).
+The performance for each dataset is the result of the test performance of the highest validation, all averaged over 5 seeds.
+Across all tables, results are shown as mean $\pm$ standard deviation.
+The best results per domain are highlighted in bold, the best overall in red and those within one standard deviation of the best are in blue.
+The evaluation metrics are classification accuracy for MUTAG, PROTEINS, NCI1, NCI109, mean absolute error (MAE) for ZINC-12K, and F1 score for MANTRA.
+This standardized protocol ensures all methods are evaluated on an even playing field, which may lead to results that differ from some prior publications.
+Please refer to Section˜F.2 for further details.
+Additionally, Appendix˜K contains the full parameter budget of the best models.
 5.2 Experimental Results
-Table 1: Accuracy on benchmark datasets. Results are shown as mean $\pm$ standard deviation. The best results are highlighted in bold, and those within one standard deviation of the best are in blue.
-Graph Tasks Topological Tasks Model MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) Graph GCN 69.79 $\pm$ 6.80 75.70 $\pm$ 2.14 72.86 $\pm$ 0.69 72.20 $\pm$ 1.22 0.62 $\pm$ 0.01 42.14 $\pm$ 2.72 47.94 $\pm$ 0.00 46.86 $\pm$ 4.50 0.00 $\pm$ 0.00 GAT 72.77 $\pm$ 2.77 76.34 $\pm$ 1.66 75.00 $\pm$ 0.99 73.80 $\pm$ 0.73 0.61 $\pm$ 0.01 18.09 $\pm$ 0.65 47.94 $\pm$ 0.00 7.45 $\pm$ 0.05 0.00 $\pm$ 0.00 GIN 79.57 $\pm$ 6.13 75.20 $\pm$ 3.30 74.26 $\pm$ 0.96 74.42 $\pm$ 0.70 0.57 $\pm$ 0.04 76.14 $\pm$ 0.14 56.28 $\pm$ 0.45 88.13 $\pm$ 0.00 0.93 $\pm$ 1.21 Simplicial SCN 73.62 $\pm$ 6.13 75.27 $\pm$ 2.14 74.49 $\pm$ 1.03 75.70 $\pm$ 1.04 0.53 $\pm$ 0.04 79.48 $\pm$ 1.36 69.55 $\pm$ 0.97 76.45 $\pm$ 3.06 5.45 $\pm$ 2.31 SCCNN 76.17 $\pm$ 6.63 74.19 $\pm$ 2.86 76.60 $\pm$ 1.75 77.12 $\pm$ 1.07 0.36 $\pm$ 0.02 95.08 $\pm$ 0.56 86.29 $\pm$ 1.23 90.20 $\pm$ 0.20 65.82 $\pm$ 2.70 SaNN 79.57 $\pm$ 3.56 75.05 $\pm$ 3.52 68.75 $\pm$ 1.64 68.17 $\pm$ 0.88 0.55 $\pm$ 0.01 81.76 $\pm$ 1.37 61.65 $\pm$ 0.55 88.46 $\pm$ 0.09 39.22 $\pm$ 2.80 GCCN 85.96 $\pm$ 4.66 74.91 $\pm$ 2.51 76.85 $\pm$ 1.50 75.76 $\pm$ 1.28 0.35 $\pm$ 0.01 86.76 $\pm$ 1.27 76.60 $\pm$ 1.67 84.20 $\pm$ 4.80 41.82 $\pm$ 20.19 HOPSE-M 86.38 $\pm$ 6.31 75.13 $\pm$ 3.66 77.35 $\pm$ 0.81 76.34 $\pm$ 1.40 0.39 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G 85.53 $\pm$ 2.77 76.27 $\pm$ 2.10 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.43 $\pm$ 0.01 81.75 $\pm$ 1.26 62.17 $\pm$ 0.98 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25 Cell CCCN 77.02 $\pm$ 9.32 73.33 $\pm$ 2.30 76.67 $\pm$ 1.48 75.35 $\pm$ 1.50 0.34 $\pm$ 0.02 - - - - CWN 80.43 $\pm$ 1.78 76.13 $\pm$ 2.70 73.93 $\pm$ 1.87 73.80 $\pm$ 2.06 0.34 $\pm$ 0.01 - - - - GCCN 86.38 $\pm$ 6.49 74.41 $\pm$ 1.77 78.23 $\pm$ 1.47 77.19 $\pm$ 0.21 0.19 $\pm$ 0.00 - - - - HOPSE-M 88.09 $\pm$ 5.55 76.99 $\pm$ 2.63 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.25 $\pm$ 0.01 - - - - HOPSE-G 88.09 $\pm$ 1.90 76.42 $\pm$ 2.19 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01 - - - -
-Accuracy comparison.
-Table 1 presents a performance comparison across topological domains: graph, simplicial, and cell.
-The proposed HOPSE-M and HOPSE-G models demonstrate consistently strong performance, particularly in the simplicial domain, answering both Q1.
+Table 1: Performance (ACC,MAE) and Total training time in seconds (T) on Graph Tasks.
+Model Dataset MUTAG PROTEINS NCI1 NCI109 ZINC-12K ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) MAE ($\downarrow$) T ($\downarrow$) Graph GCN $69.79\pm 0.06$.80) $4\pm 1.$ $75.70\pm 0.02$.14) $8\pm 2.$ $72.86\pm 0.00$.69) $53\pm 20.$ $72.20\pm 0.01$.22) $37\pm 9.$ $0.62\pm 0.00$.01) $146\pm 28.$ GCN+PSEs $84.26\pm 0.03$.56) $3\pm 1.$ $75.77\pm 0.02$.98) $23\pm 5.$ $74.88\pm 0.00$.54) $106\pm 47.$ $74.70\pm 0.00$.79) $122\pm 39.$ $0.43\pm 0.00$.03) $903\pm 40.$ GAT $2.77\pm 0.02.77$) $4\pm 1.$ $76.34\pm 0.01$.66) $8\pm 2.$ $75.00\pm 0.01$.00) $57\pm 17.$ $73.80\pm 0.00$.73) $56\pm 9.$ $0.61\pm 0.00$.01) $171\pm 64.$ GAT+PSEs $80.43\pm 0.05$.08) $5\pm 1.$ $75.48\pm 0.03$.44) $36\pm 15.$ $77.86\pm 0.01$.02) $210\pm 28.$ $75.57\pm 0.00$.60) $212\pm 89.$ $0.39\pm 0.00$.02) $708\pm 83.$ Simplicial SCCNN $76.17\pm 0.06$.63) $14\pm 3.$ $74.19\pm 0.02$.86) $53\pm 11.$ $76.60\pm 0.01$.75) $307\pm 83.$ $77.12\pm 0.01$.07) $354\pm 106.$ $0.36\pm 0.00$.02) $2060\pm 408.$ GCCN $85.96\pm 0.04$.66) $61\pm 7.$ $74.91\pm 0.02$.51) $66\pm 21.$ $76.85\pm 0.01$.50) $904\pm 180.$ $75.76\pm 0.01$.28) $538\pm 39.$ $0.35\pm 0.00$.01) $3603\pm 475.$ HOPSE-M (Our) $86.38\pm 0.06$.31) $10\pm 3.$ $75.13\pm 0.03$.66) $44\pm 17.$ $77.35\pm 0.00$.81) $201\pm 88.$ $76.34\pm 0.01$.40) $179\pm 79.$ $0.39\pm 0.00$.02) $1332\pm 721.$ HOPSE-G (Our) $85.53\pm 0.02$.77) $5\pm 1.$ $76.27\pm 0.02$.10) $33\pm 8.$ $77.70\pm 0.01$.40) $130\pm 22.$ $76.36\pm 0.01$.75) $107\pm 28.$ $0.43\pm 0.00$.01) $478\pm 107.$ Cell GCCN $86.38\pm 0.06$.49) $61\pm 18.$ $4.41\pm 0.01.77$) $59\pm 18.$ $78.23\pm 0.01$.47) $523\pm 119.$ $77.19\pm 0.00$.21) $386\pm 76.$ $0.19\pm 0.00$.01) $3301\pm 440.$ HOPSE-M (Our) $88.09\pm 0.05$.55) $7\pm 3.$ $76.99\pm 0.02$.63) $35\pm 17.$ $78.09\pm 0.01$.44) $156\pm 90.$ $77.27\pm 0.01$.01) $147\pm 81.$ $0.24\pm 0.00$.01) $760\pm 533.$ HOPSE-G (Our) $88.09\pm 0.01$.90) $7\pm 1.$ $76.42\pm 0.02$.19) $32\pm 5.$ $79.07\pm 0.00$.65) $153\pm 8.$ $77.81\pm 0.01$.26) $104\pm 38.$ $0.21\pm 0.00$.01) $1230\pm 527.$
+Table 2: Performance (F1) and Total training time in seconds (T) on Topological Tasks.
+Model Task NAME ORIENT $\beta_{1}$ $\beta_{2}$ F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) Graph GIN $76.14\pm 0.00$.14) $3568\pm 44.$ $56.28\pm 0.00$.45) $3536\pm 46.$ $88.13\pm 0.00$.00) $3674\pm 176.$ $0.93\pm 0.01$.21) $3674\pm 176.$ GIN+PSEs $94.05\pm 0.00$.15) $2174\pm 306.$ $82.53\pm 0.00$.78) $2128\pm 247.$ $90.28\pm 0.00$.05) $3011\pm 562.$ $68.43\pm 0.00$.59) $3011\pm 562.$ Simplicial SCCNN $95.08\pm 0.00$.56) $5521\pm 319.$ $86.29\pm 0.01$.23) $3536\pm 46.$ $90.20\pm 0.00$.20) $6720\pm 10.$ $65.82\pm 0.02$.70) $6720\pm 10.$ GCCN $86.76\pm 0.01$.27) $5898\pm 1961.$ $76.60\pm 0.01$.67) $4564\pm 631.$ $84.20\pm 0.04$.80) $7689\pm 2522.$ $41.82\pm 0.20$.19) $7689\pm 2522.$ HOPSE-M (Our) $96.27\pm 0.00$.28) $1530\pm 1366.$ $89.24\pm 0.00$.22) $1886\pm 1449.$ $91.02\pm 0.01$.43) $3179\pm 1756.$ $82.29\pm 0.00$.46) $3179\pm 1756.$ HOPSE-G (Our) $81.75\pm 0.01$.26) $3324\pm 1113.$ $62.17\pm 0.00$.98) $2795\pm 112.$ $88.28\pm 0.00$.08) $3876\pm 1141.$ $35.37\pm 0.02$.25) $3876\pm 1141.$
+HOMP is not a hard requirement for performance.
+Table˜1 and Table˜2 present the performance in graph and topological tasks respectively.
+The performance metric is the left most column for each dataset.
+In both scenarions, HOPSE-M and HOPSE-G models demonstrate consistently strong performance.
+This results illustrates the both Q1.
 and Q2.
-affirmatively.
+are negative, HOMP is not a requirement to capture higher-order relationships in traditional TDL benchmarks or in tasks related to topological properties.
 On standard graph classification benchmarks used in TDL, both HOPSE variants match or outperform most baselines, with HOPSE-M ranking among the top models and achieving state-of-the-art results in several cases (PROTEINS, NCI1, NCI109).
-While the performance gap on ZINC highlights the advantage of message-passing in fine-grained molecular regression tasks, HOPSE still remains competitive without relying on the HOMP procedure.
+While the performance gap on ZINC highlights the advantage of message-passing in fine-grained molecular regression tasks, HOPSE still remains competitive without relying embbedings generated with HOMP.
 The most significant gains are observed on the pure topological datasets.
@@
 This suggests that HOPSE can effectively encode and leverage topological structure.
-In summary, the HOPSE framework achieves strong performance across both graph and synthetic topological benchmarks, demonstrating that a carefully designed structural encoding pipeline can approximate or surpass topological message-passing methods.
-Detailed analysis of the results is provided in Appendix F.
-Empirical expressivity.
-Table 2: Accuracy on CSL ($\uparrow$).
-Model Mean Max Min MP-GNNs[9] 10.00 $\pm$ 0.00 10.00 10.00 SaNN 26.00 $\pm$ 6.83 33.33 20.00 SCCNN 25.33 $\pm$ 7.67 33.33 16.67 GCCN 52.78 $\pm$ 34.28 100.00 3.33 3WLGNN [32] 97.80 $\pm$ 10.92 100.00 30.00 HOPSE-M 100.00 $\pm$ 0.00 100.00 100.00 HOPSE-G 98.00 $\pm$ 4.47 100.00 90.00
-Table 2 reports the mean, minimum, and maximum test classification accuracy over five folds for the best-performing baselines, covering all domains.
-The results show that SaNN and SCCNN slightly outperform baseline MP-GNNs, while GCCN performs better on the cell domain but still falls short of 3WLGNN.
-HOPSE-M perfectly distinguishes all isomorphism classes, and HOPSE-G achieves the second-best position ( Q3.).
-Computation time comparison.
-To highlight the efficiency of HOPSE, Table 3 reports the total runtime in seconds.
-Notably, either HOPSE-G or HOPSE-M achieves the fastest overall runtime across both simplicial and cell domains.
-On graph datasets, the speedup can reach up to 7x—for example, HOPSE-G runs in 130 seconds compared to 904 seconds for GCCN on NCI1, while also delivering better performance.
-On average, HOPSE-M and HOPSE-G outperform the most competitive baselines with speedups of 1.5x and 3x over SCCN, and 3x and 5x over GCCN, respectively, while achieving top results on 6 out of 9 datasets.
-Overall, both models significantly reduce runtime while delivering competitive or superior accuracy (Q4.).
-Table 3: Runtime in seconds. Results are shown as mean $\pm$ standard deviation. The fastest model per domain is in bold and those that are within one standard deviation of the best are in blue.
-Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ($\uparrow$) ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 4 $\pm$ 1 8 $\pm$ 2 53 $\pm$ 20 37 $\pm$ 9 147 $\pm$ 28 3538 $\pm$ 38 3534 $\pm$ 49 7250 $\pm$ 234 GAT 4 $\pm$ 1 8 $\pm$ 2 57 $\pm$ 17 56 $\pm$ 9 171 $\pm$ 65 3589 $\pm$ 84 3544 $\pm$ 44 3612 $\pm$ 52 GIN 5 $\pm$ 1 9 $\pm$ 2 61 $\pm$ 24 50 $\pm$ 8 168 $\pm$ 107 3568 $\pm$ 44 3536 $\pm$ 46 3674 $\pm$ 176 Simplicial SCN 8 $\pm$ 2 43 $\pm$ 13 276 $\pm$ 63 226 $\pm$ 66 1210 $\pm$ 572 5115 $\pm$ 653 4773 $\pm$ 193 6992 $\pm$ 2049 SCCNN 14 $\pm$ 3 54 $\pm$ 11 307 $\pm$ 83 354 $\pm$ 106 2060 $\pm$ 408 5521 $\pm$ 319 6920 $\pm$ 1524 6720 $\pm$ 210 SaNN 9 $\pm$ 2 50 $\pm$ 12 250 $\pm$ 18 211 $\pm$ 66 1263 $\pm$ 252 5572 $\pm$ 1053 4619 $\pm$ 103 5220 $\pm$ 421 GCCN 61 $\pm$ 7 66 $\pm$ 21 904 $\pm$ 180 538 $\pm$ 39 3603 $\pm$ 475 5898 $\pm$ 1961 4564 $\pm$ 631 7689 $\pm$ 2522 HOPSE-M 7 $\pm$ 2 38 $\pm$ 8 164 $\pm$ 79 219 $\pm$ 26 1061 $\pm$ 405 3596 $\pm$ 519 3409 $\pm$ 162 5377 $\pm$ 185 HOPSE-G 5 $\pm$ 0 33 $\pm$ 8 130 $\pm$ 22 107 $\pm$ 28 478 $\pm$ 107 3324 $\pm$ 1113 2795 $\pm$ 112 3876 $\pm$ 1141 Cell CCCN 12 $\pm$ 2 42 $\pm$ 7 372 $\pm$ 109 272 $\pm$ 21 1621 $\pm$ 142 - - - CWN 11 $\pm$ 1 54 $\pm$ 18 302 $\pm$ 63 295 $\pm$ 46 1390 $\pm$ 97 - - - GCCN 61 $\pm$ 18 59 $\pm$ 18 523 $\pm$ 119 386 $\pm$ 76 3301 $\pm$ 440 - - - HOPSE-M 7 $\pm$ 1 37 $\pm$ 7 147 $\pm$ 24 122 $\pm$ 18 1313 $\pm$ 132 - - - HOPSE-G 7 $\pm$ 1 32 $\pm$ 5 153 $\pm$ 8 104 $\pm$ 38 1230 $\pm$ 527 - - -
-Choice of ${\mathcal{N}}_{C}$.
-The neighborhood set ${\mathcal{N}}_{C}$ is a critical design choice in the HOPSE framework, influencing both the expressivity and computational behavior of the model.
-Performed ablation studies (Appendix F) reveal that richer (i.e.
-larger and diverse) neighborhood configurations consistently yield top or near-top accuracy across complex datasets—e.g.
-MANTRA’s topological tasks, ZINC—as shown in Tables 6 and 7.
-Importantly, these gains come without significant runtime penalties: as demonstrated in Tables 10 and 11, per-epoch runtime remains almost stable, and total runtime often decreases due to faster convergence.
-Notably, the cell domain strongly aligns with expressivity and efficiency, especially on molecular tasks.
-These findings suggest a practical guideline: practitioners aiming for strong performance without exhaustive search over all neighborhood configurations can safely opt for larger ${\mathcal{N}}_{C}$ sets.
-This choice is especially advantageous for complex tasks, balancing predictive power, convergence speed, and runtime efficiency (Q1., Q2., Q4.).
+It is also noteworthy that Table˜2 shows that GIN+PSEs always outperforms GCCNs and in $\beta_{1},\beta_{2}$ it outperforms SCCN.
+This results put into question if HOMP is usefull at all for topological tasks.
+For performance over all models refer to Appendix˜G, for ablations over different neighborhoods to Appendix˜H and for ablations over PSEs to Appendix˜J.
+Speed does not need to compromise performance.
+Tables˜2 and 1 show the total runtime of training in the rightmost column of each dataset.
+[^5] This does not included preprocessing times, see Section I.3 for more in depth explanation..
+In the category of models operating over simplicial or cell complexes, a HOPSE realization always achieves the fastest overall runtime.
+On graph datasets, the speedup can reach up to $7\times$—for example, HOPSE-G runs in 130 seconds compared to 904 seconds for GCCN on NCI1, while also delivering better performance.
+HOPSE’s efficiency becomes apparant as tasks scale in size and complexity.
+In NCI1, NCI109 and ZINC it is from $3\times$ to $7\times$ faster than other models with comparable or better performance.
+In topological tasks this is exacerbated, where HOPSE-M is up to $1.5\times$ more efficienct than GIN+PSEs while outperforming all other models.
+These results show that the theoretical complexity does result in empirically reduced training time, even when comparing to the case of GNNs.
+For runtimes over all models refer to Appendix˜G, for ablations over different neighborhoods to Appendix˜I and for model sizes to Appendix˜K.
 6 Concluding remarks
 Conclusions.
-This work introduces HOPSE, a scalable framework designed to extract and leverage higher-order structural and positional information for arbitrary combinatorial representations.
-Theoretically, HOPSE matches the expressivity of state-of-the-art TNNs while avoiding the HOMP paradigm, thereby substantially reducing the computational cost of training and inference.
-Moreover, model scaling is linear with dataset size.
-These claims are supported by experiments on molecular graph benchmarks, as well as expressivity and topological tasks, where both instantiations—HOPSE-M and HOPSE-G—match or surpass state-of-the-art performance while achieving up to 7×7$\times$7 × speed-ups over HOMP-based models.
-By offering a compelling trade-off between scalability and expressivity, this work not only enables broader applicability of TDL but also facilitates the integration of higher-order interactions into other areas of deep learning—mirroring the widespread adoption of graph-based methods [47, 44, 24, 25].
+This work introduces HOPSE, a scalable and compute efficient framework designed to extract and leverage structural information of the Hasse decomposition to characterize arbitrary combinatorial representations.
+HOPSE performs competitively with HOMP without performing any type of message passing.
+Moreover, model scaling is linear with dataset size due parallel learning of structure and attributes.
+These claims are supported by experiments on molecular graph benchmarks, as well as and topological tasks, where both instantiations—HOPSE-M and HOPSE-G—match or surpass state-of-the-art performance while achieving up to $7\times$ speed-ups over HOMP-based models.
+By providing a characterization of a complex object via it’s constituting objects, Hasse graphs, HOPSE provides a new method for solving topological tasks.
+More importantly, in puts into question the necessity of HOMP.
+The cost of operating over augmented neighborhood structures blows up complexity without improving the model’s ability to learn the topology of the data.
+In a novel comparison, the ablations over GNNs with PSEs agains HOMP show that in several cases there is no need for the latter.
+Nevertheless, this work not only enables broader applicability of TDL but also facilitates the integration of higher-order interactions into other areas of deep learning—mirroring the widespread adoption of graph-based methods (Zhou et al., 2023; Wang et al., 2025; Han et al., 2022; 2023).
 Limitations & future work.
 A deeper investigation of HOPSE encodings—along with a comprehensive comparison against alternative methods across diverse domains—lies beyond the scope of this work but remains an important and intriguing direction for future research.
-Other promising avenues include exploring graph-based strategies for generating higher-order positional and structural encodings, or combining encodings derived from multiple topological domains for enhanced expressivity.
-Additionally, this work does not address whether integrating the proposed encodings into GNNs, TNNs, or transformer-based architectures leads to substantial improvements across a range of downstream tasks.
-Finally, and from a more practical standpoint, HOPSE-G realization of the framework leverages a pretrained model to extract latent structural and positional features, but it would be interesting to explore if a joint end-to-end training may further enhance performance –even at the cost of additional computational overhead.
+Other promising avenues include exploring graph-based strategies for generating higher-order positional and structural encodings, or better methods for characterizing Hasse graphs like the ones based on homomorphism counts.
+Additionally, this work does not address whether directly manufacturing encodings on a combinatorial complex provides a better characterization the joint characterization using it’s Hasse decomposition.Finally, and from a more practical standpoint, HOPSE-G realization of the framework leverages a pretrained model to extract latent structural and positional features, but it would be interesting to explore if a joint end-to-end training may further enhance performance –even at the cost of additional computational overhead.
 Appendix A Extended background
@@
 A featured graph is a tuple $\mathcal{G}_{F}=(V,E,F_{V},F_{E})$, where $F_{V}:V\to\mathbb{R}^{d_{v}}$ is a function that maps each node to a feature vector in $\mathbb{R}^{d_{v}}$ and $F_{E}:E\to\mathbb{R}^{d_{e}}$ is a function that maps each edge to a feature vector in $\mathbb{R}^{d_{e}}$.
-A topological domain is a graph generalization that captures pairwise and higher-order relationships between entities [7, 4].
-When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions [23, 38].
+A topological domain is a graph generalization that captures pairwise and higher-order relationships between entities [Bick et al., 2023, Battiston et al., 2021].
+When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions [Hajij et al., 2023, Papillon et al., 2023].
 Definition A.2 (Set-type relation).
@@
 Simplicial complexes extend graphs by incorporating hierarchical part-whole relationships through the multi-scale construction of cells.
-In this structure, nodes correspond to rank 00-cells, which can be combined to form edges (rank $1$-cells).
+In this structure, nodes correspond to rank 0-cells, which can be combined to form edges (rank $1$-cells).
 Edges can then be grouped to form faces (rank 2 cells), and faces can be combined to create volumes (rank $3$-cells), continuing in this manner.
 Consequently, the faces of a simplicial complex are triangles, volumes are tetrahedra, and higher-dimensional cells follow the same pattern.
-A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of $(k-1)$-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
+A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of (k1)(k-1)-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
 Definition A.4 (Simplicial Complex).
@@
 Unlike simplicial complexes, cell complexes are not limited to simplexes; faces can involve more than three nodes, allowing for a more flexible representation.
-This increased flexibility grants cell complexes greater expressivity compared to simplicial complexes [10, 8].
+This increased flexibility grants cell complexes greater expressivity compared to simplicial complexes Bodnar et al.
+[2021], Bodnar [2023].
 Definition A.5 (Cell complex).
@@
 Featured topological domains.
-A featured graph is a graph whose nodes or edges are equipped with feature functions [41].
+A featured graph is a graph whose nodes or edges are equipped with feature functions [Sanchez-Lengeling et al., 2021].
 Similarly, each topological element (e.g., simplex or cell) can carry feature vectors to represent some information.
@@
 Definition A.8 (Neighborhood function).
-Given a combinatorial complex ${\mathcal{T}}=(\mathcal{V},\mathcal{C},{\textnormal{k}})$, a neighborhood $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a combinatorial complex is a function that assigns to each cell $\sigma\in\mathcal{C}$ a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$.
-Each neighborhood can be used to create a Strictly Augmented Hasse graph.
+Given a combinatorial complex ${\mathcal{T}}=(\mathcal{V},\mathcal{C},{\textnormal{k}})$, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a combinatorial complex is a function that assigns to each cell $\sigma\in\mathcal{C}$ a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$.
+Each neighborhood function can be used to create a Strictly Augmented Hasse graph.
 Definition A.9 (Strictly augmented Hasse graph).
-Each particular neighborhood $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ on $\mathcal{T}$ [39, 23], defined as the directed graph whose nodes and edges are given, respectively, by $\mathcal{C}_{\mathcal{N}}=\{\sigma\in\mathcal{C}\,|\,\mathcal{N}(\sigma)\neq\emptyset\}$ and $\mathcal{E}_{\mathcal{N}}=\{(\tau,\sigma)\,|\,\tau\in\mathcal{N}(\sigma)\}$.
-Examples of neighborhood functions are incidences (connecting cells with different ranks) and adjacencies (connecting cells with the same rank), although other neighborhoods can be defined for specific tasks [3].
+Each particular neighborhood function $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ on $\mathcal{T}$ [Papillon et al., 2024, Hajij et al., 2023], defined as the directed graph whose nodes and edges are given, respectively, by $\mathcal{C}_{\mathcal{N}}=\{\sigma\in\mathcal{C}\,|\,\mathcal{N}(\sigma)\neq\emptyset\}$ and $\mathcal{E}_{\mathcal{N}}=\{(\tau,\sigma)\,|\,\tau\in\mathcal{N}(\sigma)\}$.
+Examples of neighborhood functions are incidences (connecting cells with different ranks) and adjacencies (connecting cells with the same rank), although other neighborhood functions can be defined for specific tasks [Battiloro et al., 2025].
 More in detail, given a combinatorial complex $\mathcal{T}$}:
@@
 (Incidences) The incidence neighborhood of a $t$-cell $\sigma\in\mathcal{T}$ with respect to a rank $s$ is defined as
-$\begin{split}\mathcal{I}_{s\to t}(\sigma)=\left\{\begin{array}[]{ll}\{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=s,\sigma\subset\tau\}&\text{ if $t<s$;}\\ \{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=s,\tau\subset\sigma\}&\text{ if $t>s$.}\end{array}\right.\end{split}$ (7)
+[…]
 A $s$-cell $\tau\in\mathcal{T}$ is said to be co-incident of $\sigma$ if it contains $\sigma$ and $s>t$; analogously, a $s$-cell $\tau\in\mathcal{T}$ is incident of $\sigma$ if it is is contained by $\sigma$ and $s<t$.
@@
 In this case, two $t$-cells $\sigma$ and $\tau$ are said to be adjacent if they are both contained in a $s$-cell $\delta\in\mathcal{T}$.
-Appendix B Counting neighborhoods and message passing routes
+A.3 Morphisms between CCs
+Before defining isomorphism between CCs, the definition of homomorphism between CCs is needed
+Definition A.12 (CC-Homomorphism (Definition 10 Hajij et al. [2023]).
+A homomorphism from a $CC(S_{1},\mathcal{X}_{1},\mathrm{rk}_{1})$ to another $CC(S_{2},\mathcal{X}_{2},\mathrm{rk}_{2})$ is a function $f:\mathcal{X}_{1}\to\mathcal{X}_{2}$ that sasfies the following conditions:
+  1. 1.
+    If $x,y\in\mathcal{X}_{1}$ satisfy $x\subseteq y$, then $f(x)\subseteq f(y)$
+  2. 2.
+    If $x\in\mathcal{X}_{1}$, then $\mathrm{rk}_{1}(x)\geq\mathrm{rk}_{2}(f(x))$.
+However, it is not enough to work in the attributed setting. A CC with attributes assigned to cells is
+Definition A.13 (Attributed CC [Papillon et al., 2024]).
+An attributed combinatorial complex $({\mathcal{C}},\ell)$ is a CC denoted $\mathcal{C}$} equipped with a cell coloring function $\ell:{\mathcal{C}}\to\Sigma$ where $\Sigma$ is an arbitrary normed space.
+For a cell $\sigma$ that is part of $\mathcal{C}$}, $\ell(\sigma)$ denotes the label or color of $\sigma$.
+Now note that there also exist homomorphisms induced by the neighborhood functions that can be applied to the CCs. The definition of such is
+Definition A.14 (CC-Homomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ [Papillon et al., 2024]).
+A homomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ with a neighborhood function $\mathcal{N}_{1}$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ with a neighborhood function $\mathcal{N}_{2}$ is a function $f:\mathcal{C}_{1}\to\mathcal{C}_{2}$ that satisfies the following condition: If $\sigma,\tau\in\mathcal{C}_{1}$ and $\tau\in\mathcal{N}_{1}(\sigma)$ then $f(\tau)\in\mathcal{N}_{2}(f(\sigma))$.
+This definition extends to attributed CCs with the following condition: if $\sigma,\tau\in\mathcal{C}_{1}$ have the same label then their image in $\mathcal{C}_{2}$ also has the same label, i.e for $\ell:\mathcal{C}_{1}\to\Sigma$ and $\ell^{\prime}:\mathcal{C}_{2}\to\Sigma$, if $\ell(\sigma)=\ell(\tau)$ then $f(\ell^{\prime}(\sigma))=f(\ell^{\prime}(\tau))$.
+The neighborhood functions also induce Hasse graphs.
+In fact, homomorphisms between CCs induced by neighborhood functions are one-to-one to homomorphisms of the Hasse graphs induced by the same neighborhood functions.
+Proposition A.15 (Proposition B.4 in [Papillon et al., 2024]).
+For every CC-Homomorphism $f$ from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ induced by the neighborhood functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, with domains $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, there exists a unique homomorphism between the strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$.
+The generality of combinatorial complexes complicates their direct isomorphic comparison.
+Interestingly, they can be decomposed and evaluated on their constituting pieces, Hasse graphs.
+Thus, isomorphisms can be induced by neighborhood functions, which have a one-to-one correspondance with Hasse graphs, which give an equivalence relationship between CCs.
+This type of equivalence relationship is definited as
+Definition A.16 (CC-Isomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ [Papillon et al., 2024]).
+A isomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ with a neighborhood function $\mathcal{N}_{1}$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ with a neighborhood function $\mathcal{N}_{2}$ is an invertible CC-Homomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ whose inverse is a CC-Isomorphism induced by $(\mathcal{N}_{2},\mathcal{N}_{1})$.
+In the labeled CC case, then the map also preserves labels.
+Definition A.17 (CC-Isomorphism).
+A isomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ is a bijection $f$ that fulfils:
+  • $\forall\sigma,\tau\in\mathcal{C}_{1}$, if $\sigma\subseteq\tau$ then $f(\sigma)\subseteq f(\tau)$,
+  • $\forall\sigma\in\mathcal{C}_{1}$ it holds that $\mathrm{rk}_{1}(\sigma)=\mathrm{rk}_{2}(f(\sigma))$.
+If there exists such an $f$, the CCs are deemed isomorphic, denoted $CC_{1}\simeq CC_{2}$.
+A.4 Characterization and invariants
+One of the ways to determine if graphs are different is thorugh graph invariants.
+These are functions that evaluate to the same output if the graphs are the same under an equivalence relationship, i.e.
+isomorphism.
+Definition A.18 (Graph invariant).
+A graph invariant is a function $\xi$ on a graph that assigns the same value to isomorphic graphs.
+For a fixed $G$ and any other $G^{\prime}$ such that $G\simeq G^{\prime}$ we have $\xi(G)=\xi(G^{\prime})$.
+A graph invariant distinguishes non-isomorphic graphs $G$ and $G^{\prime}$, denoted $G\not\equiv_{\xi}G^{\prime}$ if $\xi(G)\neq\xi(G^{\prime})$ and conversely if $G$ and $G^{\prime}$ are isomorphic then $G\equiv_{\xi}G^{\prime}$ if $\xi(G)=\xi(G^{\prime})$.
+An equivalent concept can be formulated in terms of combinatorial complexes, as we have a notion of equivalence from Definition˜A.17.
+Definition A.19 (CC-Invariant).
+A CC-Invariant is a function $\xi$ on a CC that assigns the same value to isomorphic CCs.
+For a fixed combinatorial complex $CC_{1}$ and any other $CC_{2}$ such that $CC_{1}\simeq CC_{2}$ we have $\xi(CC_{1})=\xi(CC_{2}^{\prime})$.
+As with graph invariants, this function helps distinguish non-isomorphic CCs.
+Now notice that a CC-Invariant induced by a neighborhood tuple $(\mathcal{N}_{1},\mathcal{N}_{2})$ is simply a graph invariant on the resuling Hasse induced by the neighborhoods.
+With this in mind, Hasse graphs can be used as piece-wise invariants to compose an invariant over the whole complex.
+Appendix B Theoretical motivation
+Proposition B.1.
+Let $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ and $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$.
+Let $\mathcal{N}_{1}$ be a neighborhood function on $\mathcal{CC}_{1}$ and $\mathcal{N}_{2}$ be a neighborhood function on $\mathcal{CC}_{2}$.
+For every CC-Isomorphism $f:\mathcal{C}_{1}\to\mathcal{C}_{2}$ induced by $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, there exists a unique graph isomorphism between their respective strictly augmented Hasse graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
+Proof.
+Note that this is trivial since every isomorphism is also a homomorphism and by Proposition˜A.15 it is unique. ∎
+First, we denote two special kinds of neighborhood functions.
+An incidence neighborhood function $\mathcal{N}$ is such that $\forall\sigma,\tau\in\mathcal{C}$, where $\sigma\in\mathcal{N}(\tau)$ or $\tau\in\mathcal{N}(\sigma)$ only one of these conditions is true:
+  1. 1.
+    $\mathrm{rk}_{1}(\sigma)>\mathrm{rk}_{1}(\tau)$,
+  2. 2.
+    $\mathrm{rk}_{1}(\sigma)<\mathrm{rk}_{1}(\tau)$.
+Conversly, a adjacency neighborhood function $\mathcal{N}$ is such that $\forall\sigma,\tau\in\mathcal{C}$, where $\sigma\in\mathcal{N}(\tau)$ or $\tau\in\mathcal{N}(\sigma)$ then $\mathrm{rk}_{1}(\sigma)=\mathrm{rk}_{1}(\tau)$.
+Proposition B.2.
+Let $CC_{1}$ and $CC_{2}$ be two CCs where $CC_{1}\not\simeq CC_{2}$.
+There exists a pair of neighborhood functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, on $CC_{1}$ and $CC_{2}$ respectively, such that the induced Hasse graphs $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are non-isomorphic.
+Proof.
+Assume that for any choice of functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ then $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are isomorphic.
+This means that there is an isomorphism $\hat{f}:\mathcal{G}_{\mathcal{N}_{1}}\to\mathcal{G}_{\mathcal{N}_{2}}$.
+This means that
+$\forall u,v\in V(\mathcal{G}_{\mathcal{N}_{1}})\text{ s.t. }(u,v)\in E(\mathcal{G}_{\mathcal{N}_{1}})\implies(\hat{f}(u),\hat{f}(v))\in E(\mathcal{G}_{\mathcal{N}_{2}})$ (9)
+and
+$\forall u^{\prime},v^{\prime}\in V(\mathcal{G}_{\mathcal{N}_{2}})\text{ s.t.
+}(u^{\prime},v^{\prime})\in E(\mathcal{G}_{\mathcal{N}_{2}})\implies(\hat{f}^{-1}(u),\hat{f}^{-1}(v^{\prime}))\in E(\mathcal{G}_{\mathcal{N}_{1}}).$ (10)
+Note that each isomorphism between Hasse graph also uniquely defines a CC-Isomorphism $f$ induced by the neighborhood functions $(\mathcal{N}_{1},\mathcal{N}_{2})$.
+Now, notice that only for a specific family of neighborhood functions, this maps encode rank information.
+This family are the incidence neighborhood functions of the form $\mathcal{I}_{s\to t}$ where $s\neq t$.
+Since $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are directed graphs that means that rank is presserved in $f$.
+In other words, this implies that there is a CC-Isomorphism $g$ since for any restriction over the domain $\mathcal{C}_{1}$ induced by any arbitrary neighborhood function, there always exists a map to $\mathcal{C}_{2}$ where
+$\forall\sigma,\tau\in\mathcal{C}_{1}\text{ if }\sigma\subseteq\tau\text{ then }f(\sigma)\subseteq f(\tau)$
+holds trivially by selecting either an adjacency or an incidence since these neighborhood functions preserve set-composition and
+$\forall\sigma\in\mathcal{C}_{1}$ (11)
+holds since $f$ preserves the rank. However, by our initial construction $CC_{1}\not\simeq CC_{2}$ and thus we reach a contradiction. ∎
+Appendix C Counting neighborhoods and message passing routes
 (b) ${\mathcal{I}}_{1\to 0}$
@@
 This discussion, while emphasizing the expressive power of combinatorial representations, highlights the challenges they pose when scaling to real-world problems.
-B.1 Neighborhood Types in Combinatorial Complexes
-Combinatorial representations allow for a variety of neighborhood types, depending on how cells of different ranks interact.
-For a 2-dimensional combinatorial complex (with 00-, $1$-, and $2$-cells), Figure 3 illustrates all possible neighborhood types.
+C.1 Neighborhood types in combinatorial complexes
+Combinatorial representations allow for a variety of neighborhood functions, depending on how cells of different ranks interact.
+For a 2-dimensional combinatorial complex (with 0-, 11-, and $2$-cells), Figure˜3 illustrates all possible neighborhood functions.
 These relations can be grouped into two main groups:
@@
     • (m) ${\mathcal{A}}_{2,1}$: faces $\rightarrow$ faces (via edges)
-B.2 Counting neighborhoods
-The combinatorial complex depicted in Figure 3 has rank $R=\dim({\mathcal{T}})=2$, which results in 12 distinct neighborhood types.
-These neighborhoods can be divided into the two categories from above: (i) incidence relations and (ii) adjacency relations.
+C.2 Counting neighborhoods
+The combinatorial complex depicted in Figure˜3 has rank $R=\dim({\mathcal{T}})=2$, which results in 12 distinct neighborhood functions.
+These neighborhood functions can be divided into the two categories from above: (i) incidence relations and (ii) adjacency relations.
 Under these assumptions, the number of incidence relations is $|I|=(R+1)R$, as each of the $R+1$ ranks can be connected to any of the remaining $R$ ranks.
@@
 $|{\mathcal{N}}_{{\mathcal{T}}}|=2(R+1)R$
-distinct neighborhood types. For the example in Figure 3, where $R=2$, this formula recovers the 12 neighborhoods observed.
-B.3 Message passing routes
-The number of neighborhoods increases quadratically with the maximum rank, which may initially seem manageable, especially since most practical datasets involve low-rank structures (e.g., $R=2$ or $R=3$).
-However, the true combinatorial complexity arises when considering how these neighborhoods can be composed.
+distinct neighborhood functions. For the example in Figure˜3, where $R=2$, this formula recovers the 12 neighborhoods observed.
+C.3 Message passing routes
+The number of neighborhood functions increases quadratically with the maximum rank, which may initially seem manageable, especially since most practical datasets involve low-rank structures (e.g., $R=2$ or $R=3$).
+However, the true combinatorial complexity arises when considering how the results neighborhoods can be composed.
 To formalize this notion, the following definition is introduced.
-Definition B.1 (Message passing route).
-A message passing route is an ordered set of neighborhoods that a model uses to propagate information across a data structure.
+Definition C.1 (Message passing route).
+A message passing route is an ordered set of neighborhood functions that a model uses to propagate information across a data structure.
 Information may flow intra-rank, through adjacency relations, or inter-rank, through incidence relations.
-For example, a network might use the message passing route $[{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ to transmit information from faces to edges, then from edges to nodes, from nodes back to faces, and finally between nodes connected by an edge.
+For example, a network might use the message passing route ${\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ to transmit information from faces to edges, then from edges to nodes, from nodes back to faces, and finally between nodes connected by an edge.
 Given a fixed combinatorial complex, the number of possible message passing routes is infinite, since any sequence can reuse the same neighborhood relations and grow to arbitrary length.
 In what follows, the analysis is restricted to a subset of message passing routes that are particularly relevant for the design of higher-order neural architectures.
-Definition B.2 (Minimal message passing route).
+Definition C.2 (Minimal message passing route).
 A minimal message passing route is a message passing route of minimal length that updates all ranks in a higher-order representation at least once.
@@
 Moreover, minimal length implies simplicity in design, yielding the most compact routes that preserve this full dependency.
-For instance, the previously introduced route $[{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ can be transformed into a minimal message passing route by removing the adjacency ${\mathcal{A}}_{0,1}$.
+For instance, the previously introduced route ${\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ can be transformed into a minimal message passing route by removing the adjacency ${\mathcal{A}}_{0,1}$.
 In this reduced route, information from all cells becomes accessible to all others within two message passing steps.
-In fact, consider the case of 00-cells: after the first step, they depend on $1$-cells; after the second step, since $1$-cells now incorporate information from $2$-cells, the 00-cells become indirectly connected to them as well.
+In fact, consider the case of 0-cells: after the first step, they depend on $1$-cells; after the second step, since $1$-cells now incorporate information from $2$-cells, the 0-cells become indirectly connected to them as well.
 By definition, minimal message passing routes rely exclusively on incidence relations, as adjacencies do not propagate information inter-rank.
@@
 This construction yields exactly one valid minimal route per permutation, resulting in a total of $N=(R+1)!$ such routes.
-In the case of graphs, where $R=1$, this results in exactly two minimal message passing routes—an intuitively expected outcome.
-B.4 Beyond minimal routes: including adjacencies
+In the case of graphs, where $R=1$, this results in exactly two minimal message passing routes.
+C.4 Beyond minimal routes: including adjacencies
 Minimal message passing routes represent only a small subset of the possible routes that can be constructed using the full set of neighborhoods in a combinatorial complex.
-A natural extension involves augmenting these minimal routes with adjacency operations—a common practice in many TDL architectures.
+A natural extension involves augmenting these minimal routes with adjacency operations, a common practice in many TDL architectures.
 For simplicity, the analysis here is restricted to the case where adjacencies are applied strictly after all incidence relations.
@@
 When combined with the $(R+1)!$ minimal incidence-based routes, the total number of extended message passing routes becomes $\hat{N}=(R+1)!\cdot(R+1)^{R}$.
-Appendix C Proofs
-Before proving Theorem 4.1, the following paragraphs show that HOPSE is at least as powerful as the WL test.
-Although this is a weaker result than the one from Theorem 4.1, the proof is more intuitive, making it a valuable tool to understand the HOPSE model.
-First, a beneficial result from Gurugubelli and Chepuri [20] regarding the WL test is reported.
-Lemma C.1.
-For a graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, the update of WL coloring of a node $\sigma\in{\mathcal{V}}$ at iteration K can be represented by the following equation
-$c_{\sigma}^{(K)}=\phi\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{U}}^{(1)}({\sigma}),...,{\mathcal{U}}^{(K)}({\sigma})\}\}\mskip-5.0mu\}\right),$ (9)
-where $\phi(\cdot)$ is some injective function and ${\mathcal{U}}^{(k)}(\sigma)$ denotes the set of upper adjacent $t$-hop neighbors of node $\sigma$.
-Using this lemma, the following theorem can now be tackled.
-Theorem C.2.
-HOPSE is at least as powerful as the WL test in distinguishing non-isomorphic graphs if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
-Proof.
-Let ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ be two graphs.
-By considering their nodes as zero rank cells and their edges as rank-1 cells, they can be considered as CC without loss of generalization.
-Given two nodes (zero rank cells) $\sigma_{1}\in{\mathcal{T}}_{1}$ and $\sigma_{2}\in{\mathcal{T}}_{2}$ let $c_{\sigma_{1}}^{(K)}$, $c_{\sigma_{2}}^{(K)}$ be the WL coloring of the two nodes after $K$ steps of the WL algorithm, and $h_{\sigma_{1}}^{(K)}$ and $h_{\sigma_{2}}^{(K)}$ be the embeddings obtained with HOPSE using a set of neighborhoods $\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$.
-Showing that for all graphs ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ and for all nodes $\sigma_{1}$ and $\sigma_{2}$, there exists a version of HOPSE such that $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies $c_{\sigma_{1}}^{(K)}=c_{\sigma_{2}}^{(K)}$ proves that HOPSE is at least as powerful as the WL isomorphism test.
-This is because it would mean that the WL test fails to assign different embeddings whenever HOPSE fails.
-As the focus is solely on the embeddings of the nodes in the considered graphs, the target rank is fixed to $r_{k}=0$, which is omitted in the remainder of the proof for clarity.
-Furthermore, since only neighborhoods ${\mathcal{N}}_{k}$ with target rank 0 are relevant, the proof is restricted to the case ${\mathcal{N}}_{k}={\mathcal{U}}^{(k)}$ for $k\in{1,\dots,K}$, where ${\mathcal{U}}^{(k)}$ denotes the upper adjacent $k$-hop neighborhood.
-Without loss of generality, the analysis can be restricted to strictly augmented multi-attributed Hasse graphs defined inSection 4 of the form ${\mathcal{J}}_{{\mathcal{N}}}=({\mathcal{C}}_{{\mathcal{N}}},{\mathcal{E}}_{{\mathcal{N}}},{\mathcal{F}}_{{\mathcal{N}}})$, where ${\mathcal{F}}_{{\mathcal{N}}}={\mathcal{F}}$ for all ${\mathcal{N}}\in N^{\rightarrow 0}$, and ${\mathcal{F}}$ denotes the initial node features of the graph.
-As only a single feature type is considered, the index $k=0$ is omitted in the remainder of the proof for clarity.
-For any node $\sigma$, by combining equation 4 to equation 5 the HOPSE embeddings can be written as
-$h_{\sigma}^{(K)}=\Theta\Big{(}\varepsilon(\mathbf{Z}),\{f\big{(}\psi(\{g(\sigma,{\mathcal{J}}_{{\mathcal{N}}_{j}})|{\mathcal{N}}_{j}\in N^{\rightarrow 0}\})\big{)}\}\Big{)}.$ (10)
-Since $g({\mathcal{J}}_{{\mathcal{N}}})$ is the function that creates the encodings for all target cells in the graph by creating a row for each cell, $g(\sigma,{\mathcal{J}}_{{\mathcal{N}}})$ is defined as the row-wise version of $g({\mathcal{J}}_{{\mathcal{N}}})$.
-The functions $g(\sigma,{\mathcal{J}}_{{\mathcal{N}}})$ can be any injective functions, but by considering only functions that depend on the 1-hop neighbors of $\sigma$ in the corresponding Hasse graph, it is immediate to see that $g(\sigma,{\mathcal{J}}_{{\mathcal{U}}_{j}})=\xi(\{\mskip-5.0mu\{{\mathcal{F}}[\tau]|\forall\tau\in{\mathcal{U}}^{(j)}(\sigma)\}\mskip-5.0mu\})$, given that ${\mathcal{N}}_{j}={\mathcal{U}}^{(j)}$.
-Since by hypothesis $g(\cdot)$ is injective, also $\xi(\cdot)$ is.
-Since $\varepsilon(\cdot)$ is also injective by hypothesis equation 10 can be rewritten as
-$\begin{split}h_{\sigma}^{(K)}&=\hat{\Theta}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in{\mathcal{U}}^{(1)}(\sigma)\}\mskip-5.0mu\},...,\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in{\mathcal{U}}^{(K)}(\sigma)\}\mskip-5.0mu\}\right)=\\ &=\hat{\Theta}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{U}}^{(1)}(\sigma),...,{\mathcal{U}}^{(K)}(\sigma)\}\}\mskip-5.0mu\}\right),\end{split}$ (11)
-where the fact that $c_{\sigma}^{(0)}=\mathbf{Z}$ was used, which means that for the WL coloring algorithm the nodes’ features are the initial features.
-Also, since the functions $\Theta(\cdot)$, $f(\cdot)$, $\psi(\cdot)$, and $\xi(\cdot)$ are injective, $\hat{\Theta}(\cdot)$ is also injective since it is a composition of injective functions.
-Since $\hat{\Theta}$ is injective $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies that the arguments of $\hat{\Theta}$ are equal.
-But those arguments are the exact arguments for the WL colorings of nodes $\sigma_{1}$ and $\sigma_{2}$ reported in equation 9, which means that $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies $c_{\sigma_{1}}^{(K)}=c_{\sigma_{2}}^{(K)}$.
-As discussed previously, this implies that HOPSE is at least as powerful as the WL test.
-∎
-The definition of the CCWL test, which generalizes the WL test to combinatorial complexes, is taken from [39].
-Definition C.3 (The CC Weisfeiler-Leman (CCWL) test on labeled combinatorial complexes).
-Let $({\mathcal{T}},F)$ be a labelled combinatorial complex. Let ${\mathcal{N}}$ be a neighborhood on ${\mathcal{C}}$. The scheme proceeds as follows:
-  • Initialization: Cells $\sigma$ are initialized with the labels given by F, i.e.: for all $\sigma\in{\mathcal{C}}$: $c_{\sigma,F}^{(0)}=F(\sigma)$.
-  • Refinement: Given colors of cells at iteration $k$, the refinement step computes the color of cell $\sigma$ at the next iteration $c_{\sigma,F}^{(k+1)}$ using a perfect HASH function as follows:
-    $\begin{split}c_{\mathcal{N}}^{(k)}(\sigma)&=\{\mskip-5.0mu\{c_{\tau,F}^{(k)}|\forall\tau\in{\mathcal{N}}(\sigma)\}\mskip-5.0mu\},\\ c_{\sigma,F}^{(k+1)}&=\text{HASH}\left(c_{\sigma,F}^{(k)},c_{\mathcal{N}}^{(k)}(\sigma)\right).\end{split}$ (12)
-  • Termination: The algorithm stops when an iteration leaves the coloring unchanged.
-To show that HOPSE is more powerful than the CCWL test, the following lemma is proven first.
-Lemma C.4.
-For a combinatorial complex $({\mathcal{T}},F)$ the update of CCWL coloring of a cell $\sigma\in{\mathcal{C}}$ at iteration $T$ can be represented as the following equation
-$c_{\sigma}^{(K)}=\phi\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}({\sigma}),...,{\mathcal{N}}^{(K)}({\sigma})\}\}\mskip-5.0mu\}\right),$ (13)
-where $\phi(\cdot)$ is some injective function, and ${\mathcal{N}}^{(k)}$ denotes the $t$-hops version of the neighborhood ${\mathcal{N}}$.
-Proof.
-The proof is by induction. For $k=0$, the equation holds trivially, since the initial coloring for cell $\sigma$ is $c_{\sigma}^{(0)}$. Assuming that the $k-1$ update is true
-$c_{\sigma}^{(t-1)}=\phi_{0}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\right).$
-By substituting this in equation 12 the $k$ update can be written as
-$\begin{split}c_{\sigma}^{(k)}=\text{HASH}\Big{(}&\phi_{0}\big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\\ \{\mskip-5.0mu\{&\phi_{0}\big{(}c_{\upsilon}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\forall\upsilon\in{\mathcal{N}}^{(1)}(\sigma)\}\mskip-5.0mu\}\Big{)}.\end{split}$
-Now, since HASH and $\phi_{0}$ are injective functions, it follows that
-$\begin{split}c_{\sigma}^{(k)}&=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\},\\ &\hskip 14.22636pt\{\mskip-5.0mu\{\big{(}c_{\upsilon}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau 1\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\forall\upsilon\in{\mathcal{N}}^{(1)}(\sigma)\}\mskip-5.0mu\}\Big{)}\\ &=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(k)}(\sigma)\}\}\mskip-5.0mu\}\Big{)}\\ &=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(k)}(\sigma)\}\}\mskip-5.0mu\}\Big{)}\end{split}$
-This is the result needed, so the theorem is proven by induction. ∎
-This result is the combinatorial complex equivalent of Lemma C.1, and it can now be used to prove the following theorem.
-Theorem C.5.
-HOPSE is at least as powerful as the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $i\in{0,...,R}$.
-Proof.
-Given that Lemma C.4 is equivalent to Lemma C.1 with ${\mathcal{N}}_{i}={\mathcal{N}}^{(i)}$ for $i\in 0,...,K$, this proof follows the same steps as the proof for Theorem C.2. ∎
-Figure 4: a.
-Two combinatorial complexes with nodes, edges, and faces.
-${\mathcal{T}}_{1}$ is a icosahedron while ${\mathcal{T}}_{2}$ is composed of five tetrahedrons.
-b.
-Their respective Hasse graphs given the choice of the neighborhood ${\mathcal{A}}_{2,1}$, which connects faces that share an edge.
-Image adapted from [39].
-Now the two main theorems of the paper can be proven.
-Theorem (Theorem 4.1).
-HOPSE is more powerful than the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
-Proof.
-Since HOPSE matches the expressive power of CCWL, as established in Theorem C.5, to prove this theorem is sufficient to present a case in which two combinatorial complexes are indistinguishable by the CCWL test but are successfully distinguished by the HOPSE algorithm.
-The example in Figure 4 illustrates two such CCs.
-Using the neighborhood ${\mathcal{A}}_{2,1}$ (the faces are adjacent if they share an edge), the corresponding Hasse graphs ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$ are constructed, both of which are 3-regular.
-It is known that the WL test cannot distinguish between regular graphs of the same degree [29].
-According to Proposition B.11 in [39], the CCWL test on these CCs is equivalent to the WL test on ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$, and thus also fails to distinguish them.
-However, the HOPSE algorithm can separate the two cases.
-In fact, by choosing the function $g_{1}(\cdot)$ to return the size of the connected component to which a node belongs, the difference between the two HOPSE encodings is apparent.
-∎
-Theorem (Theorem 4.2).
-HOPSE is cell permutation equivariant if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are cell permutation equivariant for all $k\in{1,\dots,K}$ and all $r\in{0,\dots,R}$.
-Proof.
-To demonstrate that HOPSE is cell permutation equivariant, it suffices to show that the composition of permutation equivariant functions remains permutation equivariant.
-Since, by assumption, all functions that compose HOPSE are permutation equivariant, the result follows directly.
-Let $\pi(\cdot)$ be the permutation operator.
-Then $p(\cdot)$ is permutation equivariant if $\pi(p(\cdot))=p(\pi(\cdot))$.
-Given two permutaiton equivariant functions $p(\cdot)$ and $q(\cdot)$ their composition $r(\cdot)=(p\circ q)(\cdot)$ is also permutation equivariant:
-$r(\pi(\cdot))=(p\circ q)(\pi(\cdot))=p(q(\pi(\cdot)))=p(\pi(q(\cdot)))=\pi(p(q(\cdot)))=\pi(r(\cdot)).$
 Appendix D Framework realization
-This section contains an in-depth description of the algorithmic procedure for extracting structural and positional information according two the proposed realizations HOPSE-M and HOPSE-G.
-Complementary to Section 4.3 the following is an in-depth explanation of the procedure for computing the positional and structural encodings of both realizations.
-The first step in the pipeline is the definition of the family of functions $(g_{0},\dots,g_{K})$ and the same target aggregations $\psi_{k}(\cdot)$.
-Both the variations have a different way of calculating the output embeddings so both of the cases are shown.
-D.1 Precomputation of HOPSE-M
-In this case four hand-picked structural and positional encodings were used.
+This section contains an in-depth explanation of the empirical choices of PSEs that realize the theoretical results of this work, serving as a complement to Section˜4.2.
+It is important to recall that the previous theoretical results hold for any choice of functions $g_{k}(\cdot)$ and $\psi_{k}(\cdot)$ (characterization and aggregation, respectively), as long as they fulfill the injectivity and permutation equivariance constraints.
+The fact that a certain choice of functions does not completely fulfill the injectivity or permutation equivariance constraints does not imply that they have to be completely disregarded.
+These functions are still relevant to study the empirical reach of the theoretical results in a more constrained setting.
+More importantly, fully injective and permutation equivariant functions need to be learned (or fine-tuned) while training, greatly impacting the computational advantage of this method.
+As such, these endeavours are defered to future works.
+D.1 Handcrafted PSEs
+In the literature, functions over graphs that characterize their structure are commonly referred to as positional and structural encodings (PSEs).
+While it is true that not all PSEs are meant to act as fully characterizing functions, their purpose is to represent certain characteristics of the graph that are implicitly present in its structure.
+Constructing PSEs that fully characterize graphs up to isomorphism and preserve injectivity results in the graph canonization problem Dong et al.
+[2023].
+Consequently, there will always be a trade-off between the goodness of the characterization and the computational complexity.
+With this in mind, the following choices of characterization functions stem from the literature as commonly used methods to augment the performance of GNNs.
+Additionally, while not fully injective, they are permutation equivariant, completely fulfilling that theoretical criterion.
 Let ${\mathcal{G}}=(V,E)$ be an undirected graph with $|V|=n$ nodes.
-The individual channels $g_{k}({\mathcal{G}})$ are defined as follows:
+The characterization functions $g_{k}:{\mathcal{G}}\to\mathbb{R}^{d\times d}$ with arbitrary codomains used in this work are defined as follows:
   1. 1.
-    $g_{1}(\cdot)=LapPE_{i}=[\phi_{1},\dots,\phi_{i}]\in\mathbb{R}^{n\times i},$ the first $i$ nontrivial eigenvectors of the graph Laplacian (up to frequency $i$).
+    $g_{1}(\cdot)=\mathrm{LapPE}_{i}=[\phi_{1},\dots,\phi_{i}]\in\mathbb{R}^{n\times i},$ the first $i$ nontrivial eigenvectors of the graph Laplacian (up to frequency $i$).
   2. 2.
-    $g_{2}(\cdot)=RWSE_{K}\in\mathbb{R}^{n\times K},$ the Random Walk Structural Encodings, obtained by stacking the diagonal entries of the $t$-step random-walk transition matrices for $t=1,\dots,K$ (here $K=60$).
+    $g_{2}(\cdot)=\mathrm{RWSE}_{K}\in\mathbb{R}^{n\times K},$ the Random Walk Structural Encodings, obtained by stacking the diagonal entries of the tt‑step random‑walk transition matrices for $t=1,\dots,K$ (here $K=60$).
   3. 3.
-    $g_{3}(\cdot)=HKdiagSE_{K}\in\mathbb{R}^{n\times K},$ the Heat Kernel Diagonal Structural Encodings, given by $[\exp(-\lambda_{j}\,t)]_{j=1,\dots,n;\,t=1,\dots,K}$ for the Laplacian eigenvalues $\{\lambda_{j}\}$.
+    $g_{3}(\cdot)=\mathrm{HKdiagSE}_{K}\in\mathbb{R}^{n\times K},$ the Heat Kernel Diagonal Structural Encodings, given by $\exp(-\lambda_{j}\,t)]_{j=1,\dots,n;\,t=1,\dots,K}$ for the Laplacian eigenvalues $\{\lambda_{j}\}$.
   4. 4.
-    $g_{4}(\cdot)=ElstaticPE\in\mathbb{R}^{n\times d_{4}},$ the Electrostatic Potential Positional Encodings, computed by solving the discrete Poisson equation on ${\mathcal{G}}$ with suitable boundary conditions.
-Importantly, since each $g_{k}$ depends only on the adjacency structure of ${\mathcal{G}}$, none of these channels requires access to any node feature vectors.
-For more indepth explanation of each PS encoding, see [13].
-D.2 Precomputation of HOPSE-G
+    $g_{4}(\cdot)=\mathrm{ElstaticPE}\in\mathbb{R}^{n\times d},$ the Electrostatic Potential Positional Encodings, computed by solving the discrete Poisson equation on $\mathcal{G}$} with suitable boundary conditions.
+Importantly, since each $g_{k}(\cdot)$ depends only on the adjacency structure of $\mathcal{G}$}, none of these channels requires access to any node feature vectors.
+For more indepth explanation of each PSE, see Cantürk et al.
+[2023].
+D.2 Pretrained encodings
 As mentioned previously, this method relies on the use of GPSE.
@@
 The final output $\mathbf{h}_{r,k}^{(L)}$ yields the updated channel feature.
-All MLPs share depth $L$ but may vary in hidden-layer widths $\{d_{r,\ell}\}$.
+All MLPs share depth $L$ but may vary in hidden‐layer widths $\{d_{r,\ell}\}$.
 This unified design—with layer normalization before activation—ensures stable gradient flow and flexible expressivity across channels and ranks.
 Appendix E Complexity analysis
-To analyze the time complexity (in terms of FLOPs) of the HOPSE framework, distinguishing between two components is needed: (i) operations that can be precomputed once before training, and (ii) operations that must be performed during each training iteration.
+The computational complexity of HOPSE scales linearly with the number of cells.
+This is possible by offloading the structure-dependent computations to a one-time preprocessing step.
+The resulting efficiency, confirmed by the experimental results in Table˜4, makes HOPSE well-suited for large-scale graph data.
+The following sections break down the process into two distinct parts (i.) preprocessing and (ii.) training time complexity.
 Preprocessing overhead.
-For positional and structural encodings, although the functions $g_{k}(\cdot)$ can, in principle, be any function—including trainable neural networks—the implementation is restricted to either graph-theoretic positional encodings or those generated by a pretrained GPSE model.
+For positional and structural encodings, although the functions $g_{k}(\cdot)$ can, in principle, be any function (including trainable neural networks) the implementation is restricted to either graph-theoretic positional encodings or those generated by a pretrained GPSE model.
 This allows the tensors $\mathbf{X}_{r,k}$ (as defined in equation 2) to be pre-computed before training begins.
@@
 The trainable components of the HOPSE framework are the functions $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$, for all $r=1,\dots,R$ and $k=1,\dots,K$.
-For these functions, HOPSE uses MLPs with $L$ layers
-An MLP with $L$ layers applied to $N$ elements with input and output dimensions $D_{\text{in}}$ and $D_{\text{out}}$ has a computational complexity of ${\mathcal{O}}(ND_{\text{in}}D_{\text{out}})$.
+HOPSE M and G use MLPs with $L$ layers to model these functions.
+These MLPs applied to $N$ elements with input and output dimensions $D_{\text{in}}$ and $D_{\text{out}}$ have a computational complexity of ${\mathcal{O}}(ND_{\text{in}}D_{\text{out}})$.
 For simplicity, let us consider fixed input and output dimensions for each rank.
-The complexity of generating the initial features $\hat{\mathbf{Z}}_{r}$ in equation 4 is
+The complexity of generating the initial features $\hat{\mathbf{Z}}_{r}$ in Equation˜4 is
 ${\mathcal{O}}(N_{r}D_{r}^{2}L),$
 where $N_{r}$ is the number of cells of rank $r$, and $D_{r}$ is the feature dimension.
-Next, the cost of computing $\mathbf{X}_{r,k}$ for all $K$ positional encodings (as in equation 3) is:
+Next, the cost of computing $\mathbf{X}_{r,k}$ for all $K$ positional encodings (as in Equation˜3) is:
 ${\mathcal{O}}(N_{r}D_{r}^{2}KL),$
 assuming that each function $f_{r,k}$ preserves the feature dimensionality.
-Assuming concatenation is used as the aggregation operation in equation 5, and that the output features have dimension $D_{r}$, the complexity of computing the final embeddings $\mathbf{H}_{r}$ is:
+Assuming concatenation is used as the aggregation operation in Equation˜5, and that the output features have dimension $D_{r}$, the complexity of computing the final embeddings $\mathbf{H}_{r}$ is:
 ${\mathcal{O}}(N_{r}D_{r}^{2}(K+1)L).$
@@
 This is in sharp contrast with most higher-order neural architectures.
-For example, the Generalized Combinatorial Complex Neural Network (GCCN) analyzed in [39] has the following complexity:
+For example, the Generalized Combinatorial Complex Neural Network (GCCN) analyzed in Papillon et al.
+[2024] has the following complexity:
 $C_{\text{GCCN}}=\sum_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}{\mathcal{O}}(\|\mathcal{N}\|_{0}D^{2}+\text{nrows}(\mathcal{N})d_{r^{\prime}}D+\text{nrows}(\mathcal{N}))+\sum_{r^{\prime}\in[0,R^{\prime}]}n_{r^{\prime}}n_{\mathcal{N}_{r^{\prime}}}D,$
@@
 Unlike HOPSE, the GCCN complexity depends on the number of cells and the structure of their relationships.
-Summary.
-This analysis shows that HOPSE achieves high expressive power (as demonstrated in Theorem 4.1) while maintaining a computational complexity that scales linearly with the number of cells.
-This is possible by offloading the structure-dependent computations to a one-time preprocessing step.
-The resulting efficiency, confirmed by the experimental results in Table 3, makes HOPSE well-suited for large-scale graph data.
-Appendix F Extended results
+Appendix F Experimental configuration
 All experiments were performed on a Linux system with 256 CPU cores, 1TB of RAM, and 8 NVIDIA A30 GPUs, each with 24GB of GPU memory.
 The implementation code and reproduction materials will be released upon manuscript acceptance.
-F.1 Training and evaluation.
-Five splits are generated for each dataset, with 50%/25%/25% of the data going to the training, validation, and test sets, respectively; the exception is ZINC, for which the predefined splits are used [27].
-If available, this work considers the results reported in TopoBench for the baseline models.
-In the case of the HOPSE architecture and SaNN baseline, the benchmarking of the corresponding architectures follows the predefined TopoBench suite.
+F.1 Datasets
+The datasets considered are five standard benchmarking datasets commonly used in the TDL literature, comprising four graph classification tasks and one graph regression task.
+The classification tasks use datasets from the TUDataset collection [^6] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [Morris et al., 2020].
+For regression, the ZINC dataset is used [Irwin et al., 2012].
+Following standard practice in the TDL literature [Telyatnikov et al., 2024], liftings are applied to transform these graph datasets into higher-order domains.
+Specifically, the cycle-based lifting was used for the cell domain, and the clique complex lifting for the simplicial domain.
+Additionally, the MANTRA Ballester et al.
+[2025] dataset comprises different triangulations of manifolds.
+A triangulation of a manifold $\mathcal{M}$} is a pair $(K,\iota)$ where $K$ is a simplicial complex and $\iota$ is a homeomorphism between $\mathcal{M}$} and the geometric realization of $K$.
+Concretely, triangulations encode higher-order topological information that the underlying graph, or 0-skeleton, does not encode.
+Since some triangulations are non-homeomorphic but share an underlying graph, the ability to distinguish triangulations requires that the model can access the topological information to make the correct predictions.
+This dataset emphasizes benchmarking a model’s predictive ability on purely topological properties.
+Thus, it aims to assess how much the model encodes topological information beyond the base graph.
+In contrast to the MANTRA paper, this work does not enhance node features with any additional manually derived features but instead initializes features for all cells as ones.
+No other datasets are included, such as the ones proposed in [Eitan et al., 2024], as they are subsumed by MANTRA.
+Since MANTRA contains a wider variety of homeomorphic features, categorizing them by homeomorphism class (NAME) or their Betti numbers is more general and complex.
+Regarding $\beta_{0},\beta_{1},\beta_{2}$, following the Ballester et al.
+[2025] work, in the case of predicting Betti numbers, this work focuses on training a single model as a multivariate regression task to report corresponding results for each Betti number.
+The choice of using F1 over AUROC is based on class imbalances of the Betti numbers $\beta_{0},\beta_{1},\beta_{2}$ and of the predicted properties name (NAME) and orientability ORIENT.
+Additionally the task of $\beta_{0}$ is excluded as all graphs are connected thus rendering the label as $1$ for all cases.
+F.2 Training and evaluation.
+Each dataset splits into three parts, with 50%/25%/25% of the data going to the training, validation, and test sets, respectively.
+The exception is ZINC, for which the predefined splits are used [Irwin et al., 2012].
+The results are the average over 5 runs using different seeds.
 Following the literature, different performance metrics are reported per dataset.
@@
 For each dataset, the mean and standard deviation of the chosen metric are computed across the five test sets.
-F.2 Hyperparameter search
+Discrepancy in ZINC Results and other datasets
+The discrepancies in our baseline results from common literature arise from our deliberate choice to use the TopoBench Telyatnikov et al.
+[2024] suite, a framework designed for standardized and fair comparison.
+Our goal was to evaluate all methods on an even playing field to rigorously assess the performance-efficiency trade-off.
+This standardization introduces several key differences compared to prior publications [Corso et al., 2020, Dwivedi et al., 2023, Jin et al., 2024, Bouritsas et al., 2022]:
+  • Experimental protocol: A single 50/25/25 train/validation/test split and a constrained hyperparameter search are used for all models (please refer to in Section˜F.3).
+This differs from the k-fold cross-validation and extensive tuning found in other works.
+  • Standardized components: TopoBench uses common encoders and readouts to fairly evaluate the core “backbone” of each model.
+This contrasts with studies using domain-specific knowledge, such as directly encoding molecular rings.
+  • Feature handling: To ensure a fair comparison, TopoBench sometimes disregards certain features (e.g., edge attributes in ZINC or MUTAG) in favour of standardization, which impacts the final reported performance.
+F.3 Hyperparameter search
 The HOPSE architecture is optimized over eight variations in the neighborhood set. Each variation specifies a combination of neighborhoods. The considered neighborhood combinations are as follows:
@@
   • Encoder hidden dimension: {32, 64, 128, 256}
-  • Encoder dropout: {0.25, 0.5}
+  • Model hidden dimensions: {64, 128, 256 }
+  • Encoder dropout: {0, 0.25, 0.5}
   • Number of backbone layers: {1, 2, 3, 4}
@@
 For SCN and SCCNN models, the considered readout types are direct readout (DR) and signal down-propagation (SDP) (refer to Telyatnikov et al. [2024] for in-depth discussion on these options).
-F.3 Datasets
-The datasets considered are five standard benchmarking datasets commonly used in the TDL literature, comprising four graph classification tasks and one graph regression task.
-The classification tasks use datasets from the TUDataset collection [^4] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [35].
-For regression, the ZINC dataset is used [27].
-Following standard practice in the TDL literature [42], liftings are applied to transform these graph datasets into higher-order domains.
-Specifically, the cycle-based lifting was used for the cell domain, and the clique complex lifting for the simplicial domain.
-Additionally, the MANTRA dataset comprises different triangulations of manifolds.
-A triangulation of a manifold ${\mathcal{M}}$ is a pair $(K,\iota)$ where $K$ is a simplicial complex and $\iota$ is a homeomorphism between ${\mathcal{M}}$ and the geometric realization of $K$.
-Concretely, triangulations encode higher-order topological information that the underlying graph, or 00-skeleton, does not encode.
-Since some triangulations are non-homeomorphic but share an underlying graph, the ability to distinguish triangulations requires that the model can access the topological information to make the correct predictions.
-This dataset emphasizes benchmarking a model’s predictive ability on purely topological properties.
-Thus, it aims to assess how much the model encodes topological information beyond the base graph.
-In contrast to the MANTRA paper, this work does not enhance node features with any additional manually derived features but instead initializes features for all cells as ones.
-Regarding $\beta_{0},\beta_{1},\beta_{2}$, following the Ballester et al.
-[2] work, in the case of predicting Betti numbers, this work focuses on training a single model as a multivariate regression task to report corresponding results for each Betti number.
-The choice of using F1 over AUROC is based on class imbalances of the Betti numbers $\beta_{0},\beta_{1},\beta_{2}$ and of the predicted properties name (NAME) and orientability ORIENT.
-F.4 Runtime per epoch
-In these experiments, the runtime per epoch is shown for each dataset.
-Table 4 shows the average runtime per epoch on the benchmarked datasets.
-Our method has the lowest runtime per epoch across all datasets among HOMP methods.
-These results validate the theoretical findings presented.
-In addition, HOPSE-G is faster than HOPSE-M in all benchmarks except NCI109.
-This aligns with complexity and scales linearly with the number $K$ of positional encodings considered.
-Compared to graph message passing methods, such as GCN, GAT, and GIN, the HOPSE framework performs competitively on smaller datasets such as MUTAG, PROTEINS, NCI1, and NCI109.
-On larger datasets, it speeds up up to 4x in ZINC and closer to 2x in MANTRA.
-Table 4: Runtime per epoch in seconds.
-Results are shown as mean and standard deviation.
-The best result per domain and per model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
-The HOPSE framework is more efficient than all other HOMP-based methods across all datasets.
-In the larger datasets such as ZINC and MANTRA, it is more efficient than traditional message passing over graphs.
-The variation between the simplicial and cell domain for both models is relatively low
-Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 0.03 $\pm$ 0.00 0.05 $\pm$ 0.00 0.26 $\pm$ 0.01 0.27 $\pm$ 0.02 1.24 $\pm$ 0.01 6.24 $\pm$ 0.11 6.24 $\pm$ 0.10 6.43 $\pm$ 0.08 GAT 0.04 $\pm$ 0.00 0.07 $\pm$ 0.00 0.34 $\pm$ 0.02 0.33 $\pm$ 0.02 1.24 $\pm$ 0.01 6.38 $\pm$ 0.14 6.16 $\pm$ 0.12 6.44 $\pm$ 0.15 GIN 0.03 $\pm$ 0.00 0.06 $\pm$ 0.00 0.27 $\pm$ 0.02 0.27 $\pm$ 0.02 1.19 $\pm$ 0.00 6.37 $\pm$ 0.12 6.18 $\pm$ 0.06 6.46 $\pm$ 0.23 Simplicial SCN 0.09 $\pm$ 0.00 0.42 $\pm$ 0.01 1.60 $\pm$ 0.03 1.59 $\pm$ 0.02 6.93 $\pm$ 0.09 6.83 $\pm$ 0.08 7.35 $\pm$ 0.38 6.92 $\pm$ 0.08 SCCNN 0.09 $\pm$ 0.00 0.43 $\pm$ 0.01 1.65 $\pm$ 0.02 1.66 $\pm$ 0.03 10.40 $\pm$ 0.06 8.47 $\pm$ 0.63 8.25 $\pm$ 0.12 9.24 $\pm$ 0.15 SaNN 0.03 $\pm$ 0.00 0.18 $\pm$ 0.00 0.64 $\pm$ 0.02 0.65 $\pm$ 0.02 0.63 $\pm$ 0.03 7.21 $\pm$ 0.08 7.77 $\pm$ 0.15 7.34 $\pm$ 0.10 GCCN 0.04 $\pm$ 0.00 3.06 $\pm$ 0.02 0.43 $\pm$ 0.02 0.39 $\pm$ 0.02 - 6.79 $\pm$ 0.24 6.13 $\pm$ 0.09 7.70 $\pm$ 0.05 HOPSE-M 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00 0.32 $\pm$ 0.00 0.47 $\pm$ 0.03 0.35$\pm$ 0.01 5.13 $\pm$ 0.14 5.60 $\pm$ 0.55 5.54 $\pm$ 0.05 HOPSE-G 0.02 $\pm$ 0.00 0.10 $\pm$ 0.02 0.31 $\pm$ 0.00 0.37 $\pm$ 0.02 0.28 $\pm$ 0.01 4.48 $\pm$ 0.14 4.23 $\pm$ 0.12 4.48 $\pm$ 0.11 Cell CCCN 0.11 $\pm$ 0.00 0.32 $\pm$ 0.01 1.33 $\pm$ 0.02 1.33 $\pm$ 0.02 6.14 $\pm$ 0.04 - - - CWN 0.10 $\pm$ 0.00 0.37 $\pm$ 0.02 1.38 $\pm$ 0.03 1.37 $\pm$ 0.01 5.86 $\pm$ 0.05 - - - HOPSE-M 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.34 $\pm$ 0.00 0.33 $\pm$ 0.01 0.47 $\pm$ 0.02 - - - HOPSE-G 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.40 $\pm$ 0.02 0.37 $\pm$ 0.00 0.32 $\pm$ 0.01 - - -
-F.5 Ablation on neighborhood set ${\mathcal{N}}_{C}$
-To gain more insight into how the choice of neighborhoods ${\mathcal{N}}_{C}$ impacts the performance of the HOPSE framework, this section shows an ablation over the considered sets of neighborhoods.
-The neighborhood sets listed in the tables are defined in Table 5.
-Each configuration differs in the number and type of neighborhoods considered, directly influencing computational complexity and model expressivity.
-Specifically, the Adj family explores adjacency-based neighborhoods of increasing order: Adj-1 uses a single neighborhood, while Adj-2 and Adj-3 consider two and three neighborhoods, respectively.
-The Inc configurations mix the base adjacency (${\mathcal{A}}_{0,1}$) and incidence, with Inc-1 and Inc-2 using three neighborhoods, and Inc-3 extending to five.
-The Mix configurations, Mix-1 and Mix-2, are the biggest, combining a total of eight and six neighborhoods, respectively.
-Mix-1 uses all previous adjacencies and incidences combined while Mix-2 includes more complex neighborhoods.
-One of such is ${\mathcal{A}}_{0,2}$, which means the adjacencies of 00-cells with $2$-cells through $1$-cells.
-Effectively, this means that the structure of neighborhoods between non-adjacent ranks (i.e., a difference of $1$ in the rank number) is considered.
-As will be seen later in the section, considering these non-standard neighborhoods does account for a difference in performance for the more complex benchmarks.
+Appendix G Complete model performance and runtimes
+Table 3: Performance on benchmark datasets.
+Graph Tasks Topological Tasks Model MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC-12K ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) Graph GCN 69.79 $\pm$ 6.80 75.70 $\pm$ 2.14 72.86 $\pm$ 0.69 72.20 $\pm$ 1.22 0.62 $\pm$ 0.01 42.14 $\pm$ 2.72 47.94 $\pm$ 0.00 46.86 $\pm$ 4.50 0.00 $\pm$ 0.00 GAT 72.77 $\pm$ 2.77 76.34 $\pm$ 1.66 75.00 $\pm$ 0.99 73.80 $\pm$ 0.73 0.61 $\pm$ 0.01 18.09 $\pm$ 0.65 47.94 $\pm$ 0.00 7.45 $\pm$ 0.05 0.00 $\pm$ 0.00 GIN 79.57 $\pm$ 6.13 75.20 $\pm$ 3.30 74.26 $\pm$ 0.96 74.42 $\pm$ 0.70 0.57 $\pm$ 0.04 76.14 $\pm$ 0.14 56.28 $\pm$ 0.45 88.13 $\pm$ 0.00 0.93 $\pm$ 1.21 Simplicial SCN 73.62 $\pm$ 6.13 75.27 $\pm$ 2.14 74.49 $\pm$ 1.03 75.70 $\pm$ 1.04 0.53 $\pm$ 0.04 79.48 $\pm$ 1.36 69.55 $\pm$ 0.97 76.45 $\pm$ 3.06 5.45 $\pm$ 2.31 SCCNN 76.17 $\pm$ 6.63 74.19 $\pm$ 2.86 76.60 $\pm$ 1.75 77.12 $\pm$ 1.07 0.36 $\pm$ 0.02 95.08 $\pm$ 0.56 86.29 $\pm$ 1.23 90.20 $\pm$ 0.20 65.82 $\pm$ 2.70 SaNN 79.57 $\pm$ 3.56 75.05 $\pm$ 3.52 68.75 $\pm$ 1.64 68.17 $\pm$ 0.88 0.55 $\pm$ 0.01 81.76 $\pm$ 1.37 61.65 $\pm$ 0.55 88.46 $\pm$ 0.09 39.22 $\pm$ 2.80 GCCN 85.96 $\pm$ 4.66 74.91 $\pm$ 2.51 76.85 $\pm$ 1.50 75.76 $\pm$ 1.28 0.35 $\pm$ 0.01 86.76 $\pm$ 1.27 76.60 $\pm$ 1.67 84.20 $\pm$ 4.80 41.82 $\pm$ 20.19 HOPSE-M (Our) 86.38 $\pm$ 6.31 75.13 $\pm$ 3.66 77.35 $\pm$ 0.81 76.34 $\pm$ 1.40 0.39 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G (Our) 85.53 $\pm$ 2.77 76.27 $\pm$ 2.10 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.43 $\pm$ 0.01 81.75 $\pm$ 1.26 62.17 $\pm$ 0.98 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25 Cell CCCN 77.02 $\pm$ 9.32 73.33 $\pm$ 2.30 76.67 $\pm$ 1.48 75.35 $\pm$ 1.50 0.34 $\pm$ 0.02 - - - - CWN 80.43 $\pm$ 1.78 76.13 $\pm$ 2.70 73.93 $\pm$ 1.87 73.80 $\pm$ 2.06 0.34 $\pm$ 0.01 - - - - GCCN 86.38 $\pm$ 6.49 74.41 $\pm$ 1.77 78.23 $\pm$ 1.47 77.19 $\pm$ 0.21 0.19 $\pm$ 0.00 - - - - HOPSE-M (Our) 88.09 $\pm$ 5.55 76.99 $\pm$ 2.63 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.24 $\pm$ 0.01 - - - - HOPSE-G (Our) 88.09 $\pm$ 1.90 76.42 $\pm$ 2.19 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01 - - - -
+Table 4: Total end-to-end training runtime in seconds.
+Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC-12K NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 4 $\pm$ 1 8 $\pm$ 2 53 $\pm$ 20 37 $\pm$ 9 147 $\pm$ 28 3538 $\pm$ 38 3534 $\pm$ 49 7250 $\pm$ 234 GAT 4 $\pm$ 1 8 $\pm$ 2 57 $\pm$ 17 56 $\pm$ 9 171 $\pm$ 65 3589 $\pm$ 84 3544 $\pm$ 44 3612 $\pm$ 52 GIN 5 $\pm$ 1 9 $\pm$ 2 61 $\pm$ 24 50 $\pm$ 8 168 $\pm$ 107 3568 $\pm$ 44 3536 $\pm$ 46 3674 $\pm$ 176 Simplicial SCN 8 $\pm$ 2 43 $\pm$ 13 276 $\pm$ 63 226 $\pm$ 66 1210 $\pm$ 572 5115 $\pm$ 653 4773 $\pm$ 193 6992 $\pm$ 2049 SCCNN 14 $\pm$ 3 54 $\pm$ 11 307 $\pm$ 83 354 $\pm$ 106 2060 $\pm$ 408 5521 $\pm$ 319 6920 $\pm$ 1524 6720 $\pm$ 210 SaNN 9 $\pm$ 2 50 $\pm$ 12 250 $\pm$ 18 211 $\pm$ 66 1263 $\pm$ 252 5572 $\pm$ 1053 4619 $\pm$ 103 5220 $\pm$ 421 GCCN 61 $\pm$ 7 66 $\pm$ 21 904 $\pm$ 180 538 $\pm$ 39 3603 $\pm$ 475 5898 $\pm$ 1961 4564 $\pm$ 631 7689 $\pm$ 2522 HOPSE-M (Ours) 7 $\pm$ 2 38 $\pm$ 8 164 $\pm$ 79 219 $\pm$ 26 1061 $\pm$ 405 3596 $\pm$ 519 3409 $\pm$ 162 5377 $\pm$ 185 HOPSE-G (Ours) 5 $\pm$ 0 33 $\pm$ 8 130 $\pm$ 22 107 $\pm$ 28 478 $\pm$ 107 3324 $\pm$ 1113 2795 $\pm$ 112 3876 $\pm$ 1141 Cell CCCN 12 $\pm$ 2 42 $\pm$ 7 372 $\pm$ 109 272 $\pm$ 21 1621 $\pm$ 142 - - - CWN 11 $\pm$ 1 54 $\pm$ 18 302 $\pm$ 63 295 $\pm$ 46 1390 $\pm$ 97 - - - GCCN 61 $\pm$ 18 59 $\pm$ 18 523 $\pm$ 119 386 $\pm$ 76 3301 $\pm$ 440 - - - HOPSE-M (Ours) 7 $\pm$ 1 37 $\pm$ 7 147 $\pm$ 24 122 $\pm$ 18 1313 $\pm$ 132 - - - HOPSE-G (Ours) 7 $\pm$ 1 32 $\pm$ 5 153 $\pm$ 8 104 $\pm$ 38 1230 $\pm$ 527 - - -
+Appendix H Performance ablation across different neighborhoods
+The choice of neighborhood functions set ${\mathcal{N}}_{C}$ is a crucial and non-trivial choice for the performance of HOPSE.
+This section contains the performance of HOPSE for each choice of neighborhood encoded by a distinct name, as reported in Table˜5.
 Table 5: Taxonomy of neighborhood combinations
@@
 Inc-3 ${\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
-Mix-1 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1}$
+Mix-1 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
 Mix-2 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{0,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{A}}_{2,0}$
-F.5.1 Analysis of performance across neighborhoods
-Analysis of performance for different domains.
-The results in Table 6 and Table 7 demonstrate insightful patterns in how neighborhood selection impacts performance depending on domain type and dataset complexity.
-Comparing performance between HOPSE-M and HOPSE-G variants across different neighborhood configurations further suggests that the pre-training strategy employed in GPSE provides rich neighborhood information, particularly for the molecular datasets, as the GPSE encoder has been pre-trained on molecular data.
-This is most evident in the cell domain results for ZINC, where HOPSE-G with Mix-2 achieves a 16% improvement compared to the best HOPSE-M configuration.
-Table 6: Performance per neighborhood on simplicial.
-Results are shown as mean $\pm$ standard deviation.
-The best result per neighborhood is in bold and results within one standard deviation of the best per neighborhood are highlighted in blue.
-The performance differences between neighborhood configurations are minimal, with most results falling within one standard deviation of each other.
-As datasets grow bigger and more complex, such as ZINC and the MANTRA benchmarks, the neighborhood choice starts to matter.
-The results show that for simplicial models, the most informative neighborhoods for ZINC are the extended groups of adjacencies Adj-2 and Adj-3.
-However, for the MANTRA tasks, the most important neighborhood is Mix-2
-Graph Tasks Topological Tasks Nbhd MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) HOPSE-M Adj-1 86.38 $\pm$ 6.31 74.70 $\pm$ 4.04 77.35 $\pm$ 0.81 75.41 $\pm$ 1.17 0.40 $\pm$ 0.01 79.58 $\pm$ 2.88 60.93 $\pm$ 3.03 88.35 $\pm$ 0.09 28.25 $\pm$ 2.71 Adj-2 85.53 $\pm$ 7.12 73.19 $\pm$ 3.32 76.26 $\pm$ 1.59 75.59 $\pm$ 0.41 0.39 $\pm$ 0.02 80.35 $\pm$ 3.96 74.63 $\pm$ 1.77 89.97 $\pm$ 0.18 64.06 $\pm$ 2.27 Adj-3 85.96 $\pm$ 5.75 75.13 $\pm$ 3.66 76.44 $\pm$ 1.44 75.22 $\pm$ 2.06 0.39 $\pm$ 0.02 84.72 $\pm$ 1.23 78.19 $\pm$ 2.70 90.13 $\pm$ 0.49 70.10 $\pm$ 1.86 Inc-1 85.11 $\pm$ 6.20 73.98 $\pm$ 2.97 76.40 $\pm$ 0.99 76.09 $\pm$ 1.50 0.40 $\pm$ 0.01 84.01 $\pm$ 6.84 65.10 $\pm$ 9.66 90.06 $\pm$ 0.09 65.02 $\pm$ 1.58 Inc-2 84.26 $\pm$ 6.13 73.69 $\pm$ 2.76 76.40 $\pm$ 1.45 75.39 $\pm$ 1.13 0.42 $\pm$ 0.05 80.67 $\pm$ 2.88 69.51 $\pm$ 6.86 89.51 $\pm$ 0.94 64.28 $\pm$ 3.08 Inc-3 85.11 $\pm$ 5.83 73.55 $\pm$ 3.38 76.56 $\pm$ 1.40 76.34 $\pm$ 1.40 0.42 $\pm$ 0.02 82.56 $\pm$ 4.15 77.35 $\pm$ 2.12 89.89 $\pm$ 0.13 63.12 $\pm$ 1.68 Mix-1 84.26 $\pm$ 5.12 73.69 $\pm$ 3.21 76.44 $\pm$ 1.62 75.57 $\pm$ 1.09 0.40 $\pm$ 0.01 90.37 $\pm$ 1.29 72.64 $\pm$ 7.53 90.21 $\pm$ 0.18 67.14 $\pm$ 2.18 Mix-2 83.83 $\pm$ 6.49 73.41 $\pm$ 2.95 75.14 $\pm$ 1.59 74.87 $\pm$ 0.94 0.41 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G Adj-1 82.98 $\pm$ 7.22 74.05 $\pm$ 2.17 77.18 $\pm$ 1.45 76.21 $\pm$ 0.65 0.44 $\pm$ 0.01 80.43 $\pm$ 1.03 60.92 $\pm$ 1.18 88.18 $\pm$ 0.03 24.59 $\pm$ 1.87 Adj-2 85.53 $\pm$ 2.77 74.05 $\pm$ 3.08 77.51 $\pm$ 1.15 74.91 $\pm$ 1.17 0.43 $\pm$ 0.00 80.19 $\pm$ 1.13 61.88 $\pm$ 0.83 88.13 $\pm$ 0.02 20.31 $\pm$ 2.57 Adj-3 85.11 $\pm$ 3.36 73.26 $\pm$ 4.20 77.32 $\pm$ 1.19 74.79 $\pm$ 0.91 0.43 $\pm$ 0.01 79.11 $\pm$ 1.53 61.86 $\pm$ 1.10 88.11 $\pm$ 0.02 19.08 $\pm$ 3.57 Inc-1 84.68 $\pm$ 3.50 73.19 $\pm$ 3.52 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.44 $\pm$ 0.00 80.70 $\pm$ 0.80 62.17 $\pm$ 0.98 87.31 $\pm$ 1.17 20.47 $\pm$ 3.22 Inc-2 82.13 $\pm$ 5.75 76.27 $\pm$ 2.10 75.66 $\pm$ 1.52 74.17 $\pm$ 1.99 0.45 $\pm$ 0.02 78.92 $\pm$ 1.25 59.57 $\pm$ 1.56 88.08 $\pm$ 0.08 20.28 $\pm$ 3.48 Inc-3 77.87 $\pm$ 4.90 75.48 $\pm$ 3.15 76.15 $\pm$ 1.77 75.14 $\pm$ 0.75 0.44 $\pm$ 0.01 78.78 $\pm$ 0.74 60.35 $\pm$ 1.05 88.12 $\pm$ 0.04 18.82 $\pm$ 3.54 Mix-1 79.15 $\pm$ 7.88 75.70 $\pm$ 3.24 76.50 $\pm$ 1.09 75.35 $\pm$ 0.97 0.45 $\pm$ 0.01 79.76 $\pm$ 0.51 60.33 $\pm$ 1.60 88.07 $\pm$ 0.09 23.48 $\pm$ 3.44 Mix-2 83.40 $\pm$ 8.56 74.05 $\pm$ 3.39 77.02 $\pm$ 0.92 74.75 $\pm$ 1.47 0.46 $\pm$ 0.01 81.75 $\pm$ 1.26 61.32 $\pm$ 2.67 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25
-Efficacy of neighborhoods by domain.
-One important observation is how neighborhood efficacy varies between the simplicial and cell domains.
-While in the simplicial domain, adding mixed neighborhoods occasionally leads to performance degradation for some datasets (e.g., PROTEINS where Inc-1 outperforms Mix configurations for HOPSE-M), in the cell domain, more comprehensive neighborhood sets either improve or do not significantly impact performance.
-This domain-dependent behavior suggests that the optimal neighborhood configuration should be ideally selected based on task complexity and the topological representation chosen for the data; however, choosing comprehensive neighborhood sets on average allows for achieving good performance while not harming the computational cost (see Section F.5.2 for details on computational analysis).
-These findings demonstrate that the choice of neighborhood configuration represents a critical design decision in topological learning frameworks.
+For small and simple datasets (MUTAG, PROTEINS, NCI1, NCI109), the performance differences between neighborhood configurations are minimal, with most results falling within one standard deviation of each other.
+As datasets grow bigger and more complex, such as ZINC-12K and the MANTRA, the neighborhood choice starts to matter.
+The results show that for simplicial models, the best performing neighborhoods for ZINC-12K are the extended groups of adjacencies Adj-2 and Adj-3.
+However, for the MANTRA tasks, the most important neighborhood is Mix-2.Note that the per neighborhood ablation results differ from the main table due to some neighborhoods having a higher validation which does not lead to ahigher test.These findings demonstrate that the choice of neighborhood configuration represents a critical design decision in the topological learning frameworks.
 While simpler configurations may suffice for some graph benchmarks, an effective strategy is to consider a greater neighbourhood set ${\mathcal{N}}_{C}$ to capture the full expressive power of topological structures.
-It is especially evident on complex tasks, topological benchmarking datasets (MANTRA), and relatively large datasets like ZINC.
-Table 7: Performance per neighborhood on cell.
-Results are shown as mean $\pm$ standard deviation.
-The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
-In smaller and simpler datasets such as MUTAG, PROTEINS, NCI1, and NCI109, the best neighborhoods are not always the most encompassing, but they perform closely to others.
-However, on a more complex dataset such as ZINC there is a significant difference between simpler neighborhoods like the Adj[1-3] and Mix-2, particularly in the case of HOPSE-G.
-Graph Tasks Nbhd MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) HOPSE-M Adj-1 88.09 $\pm$ 5.55 75.77 $\pm$ 2.67 76.85 $\pm$ 0.83 76.34 $\pm$ 1.02 0.29 $\pm$ 0.01 Adj-2 82.13 $\pm$ 4.15 74.55 $\pm$ 2.65 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.27 $\pm$ 0.01 Adj-3 85.53 $\pm$ 3.50 75.27 $\pm$ 2.19 77.57 $\pm$ 1.48 75.37 $\pm$ 0.79 0.28 $\pm$ 0.01 Inc-1 86.81 $\pm$ 4.61 76.99 $\pm$ 2.63 77.32 $\pm$ 1.62 75.70 $\pm$ 1.36 0.27 $\pm$ 0.02 Inc-2 83.40 $\pm$ 3.81 75.56 $\pm$ 2.26 75.89 $\pm$ 2.14 75.43 $\pm$ 1.98 0.29 $\pm$ 0.02 Inc-3 82.98 $\pm$ 3.36 74.70 $\pm$ 3.05 77.14 $\pm$ 0.46 74.46 $\pm$ 1.52 0.25 $\pm$ 0.01 Mix-1 81.28 $\pm$ 5.90 75.13 $\pm$ 3.09 77.26 $\pm$ 1.70 76.13 $\pm$ 1.38 0.25 $\pm$ 0.01 Mix-2 84.26 $\pm$ 6.13 72.62 $\pm$ 2.55 76.69 $\pm$ 1.25 76.57 $\pm$ 1.55 0.26 $\pm$ 0.01 HOPSE-G Adj-1 82.13 $\pm$ 7.76 76.27 $\pm$ 2.70 78.31 $\pm$ 0.84 77.48 $\pm$ 0.99 0.26 $\pm$ 0.00 Adj-2 87.66 $\pm$ 6.96 76.42 $\pm$ 2.19 78.83 $\pm$ 1.40 77.48 $\pm$ 0.64 0.24 $\pm$ 0.00 Adj-3 88.09 $\pm$ 1.90 75.91 $\pm$ 1.89 78.35 $\pm$ 1.11 76.94 $\pm$ 1.46 0.23 $\pm$ 0.00 Inc-1 85.11 $\pm$ 6.56 74.62 $\pm$ 3.99 77.57 $\pm$ 1.67 76.81 $\pm$ 1.23 0.25 $\pm$ 0.01 Inc-2 82.55 $\pm$ 8.82 74.91 $\pm$ 2.11 78.15 $\pm$ 0.75 76.40 $\pm$ 1.86 0.27 $\pm$ 0.01 Inc-3 84.68 $\pm$ 7.88 74.12 $\pm$ 3.39 76.81 $\pm$ 0.59 75.26 $\pm$ 1.16 0.26 $\pm$ 0.00 Mix-1 88.09 $\pm$ 3.23 76.20 $\pm$ 2.62 79.03 $\pm$ 0.73 76.88 $\pm$ 0.95 0.23 $\pm$ 0.00 Mix-2 85.11 $\pm$ 6.02 75.63 $\pm$ 2.35 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01
-F.5.2 Analysis of time efficiency across neighborhoods
-This section presents the effects of the choice of ${\mathcal{N}}_{C}$ on the total runtime and per-epoch performance of the HOPSE framework.
-The results in Table 8 and Table 9 report the per-epoch runtime for each neighborhood configuration across both the simplicial and cell domains, under the HOPSE-M and HOPSE-G models, respectively.
-While Table 10 and Table 11 present the total runtime to convergence for eight different neighbourhoods for simplicial and cell domains, and HOPSE-M and HOPSE-G correspondingly.
-Table 8: Runtime per epoch per neighborhood on simplicial.
-Results are shown as mean $\pm$ standard deviation.
-The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
-Simpler configurations such as Adj-1 consistently yield the lowest runtime across domains, particularly for HOPSE-M, where Adj-1 achieves the fastest epoch times in 5 out of 8 datasets,
-Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.32 $\pm$ 0.00 0.32 $\pm$ 0.00 0.30 $\pm$ 0.01 4.12 $\pm$ 0.08 4.41 $\pm$ 0.11 4.48 $\pm$ 0.16 Adj-2 0.03 $\pm$ 0.01 0.12 $\pm$ 0.00 0.36 $\pm$ 0.01 0.36 $\pm$ 0.00 0.35 $\pm$ 0.01 4.37 $\pm$ 0.13 4.15 $\pm$ 0.09 4.54 $\pm$ 0.21 Adj-3 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.37 $\pm$ 0.00 0.39 $\pm$ 0.00 0.37 $\pm$ 0.01 4.42 $\pm$ 0.11 4.41 $\pm$ 0.22 4.91 $\pm$ 0.07 Inc-1 0.03 $\pm$ 0.00 0.14 $\pm$ 0.01 0.44 $\pm$ 0.01 0.40 $\pm$ 0.01 0.40 $\pm$ 0.01 6.28 $\pm$ 1.34 5.14 $\pm$ 0.24 5.41 $\pm$ 0.17 Inc-2 0.03 $\pm$ 0.01 0.12 $\pm$ 0.01 0.36 $\pm$ 0.00 0.39 $\pm$ 0.00 0.34 $\pm$ 0.01 4.41 $\pm$ 0.05 4.29 $\pm$ 0.06 4.87 $\pm$ 0.12 Inc-3 0.02 $\pm$ 0.00 0.16 $\pm$ 0.00 0.46 $\pm$ 0.01 0.47 $\pm$ 0.03 0.42 $\pm$ 0.01 5.89 $\pm$ 0.31 5.71 $\pm$ 0.23 5.70 $\pm$ 0.23 Mix-1 0.03 $\pm$ 0.00 0.17 $\pm$ 0.01 0.51 $\pm$ 0.00 0.51 $\pm$ 0.02 0.48 $\pm$ 0.01 6.00 $\pm$ 0.29 6.25 $\pm$ 0.31 6.51 $\pm$ 0.23 Mix-2 0.02 $\pm$ 0.00 0.15 $\pm$ 0.00 0.45 $\pm$ 0.01 0.44 $\pm$ 0.00 0.40 $\pm$ 0.01 5.13 $\pm$ 0.14 5.60 $\pm$ 0.55 5.54 $\pm$ 0.05 HOPSE-G Adj-1 0.02 $\pm$ 0.00 0.10 $\pm$ 0.01 0.28 $\pm$ 0.01 0.28 $\pm$ 0.00 0.26 $\pm$ 0.02 3.19 $\pm$ 0.09 3.18 $\pm$ 0.12 3.55 $\pm$ 0.12 Adj-2 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.30 $\pm$ 0.00 0.31 $\pm$ 0.01 0.28 $\pm$ 0.01 3.53 $\pm$ 0.12 3.41 $\pm$ 0.10 3.67 $\pm$ 0.12 Adj-3 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.33 $\pm$ 0.02 0.34 $\pm$ 0.01 0.29 $\pm$ 0.02 3.79 $\pm$ 0.14 3.84 $\pm$ 0.14 3.98 $\pm$ 0.07 Inc-1 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.38 $\pm$ 0.01 0.37 $\pm$ 0.02 0.34 $\pm$ 0.01 4.44 $\pm$ 0.14 4.23 $\pm$ 0.12 4.64 $\pm$ 0.08 Inc-2 0.02 $\pm$ 0.00 0.10 $\pm$ 0.00 0.31 $\pm$ 0.00 0.31 $\pm$ 0.00 0.28 $\pm$ 0.01 3.74 $\pm$ 0.17 3.66 $\pm$ 0.10 3.78 $\pm$ 0.06 Inc-3 0.02 $\pm$ 0.00 0.13 $\pm$ 0.00 0.41 $\pm$ 0.00 0.40 $\pm$ 0.01 0.38 $\pm$ 0.03 5.17 $\pm$ 0.40 4.75 $\pm$ 0.23 5.13 $\pm$ 0.06 Mix-1 0.02 $\pm$ 0.00 0.16 $\pm$ 0.01 0.46 $\pm$ 0.02 0.46 $\pm$ 0.01 0.44 $\pm$ 0.01 5.50 $\pm$ 0.24 5.74 $\pm$ 0.22 5.84 $\pm$ 0.38 Mix-2 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.38 $\pm$ 0.00 0.38 $\pm$ 0.01 0.36 $\pm$ 0.02 4.48 $\pm$ 0.14 4.55 $\pm$ 0.19 4.84 $\pm$ 0.11
-Table 9: Runtime per epoch per neighborhood on cell.
-Results are shown as mean $\pm$ standard deviation.
-The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
-Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.31 $\pm$ 0.01 0.32 $\pm$ 0.01 0.30 $\pm$ 0.01 Adj-2 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.34 $\pm$ 0.00 0.33 $\pm$ 0.00 0.30 $\pm$ 0.01 Adj-3 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.35 $\pm$ 0.00 0.34 $\pm$ 0.00 0.32 $\pm$ 0.01 Inc-1 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.40 $\pm$ 0.01 0.40 $\pm$ 0.02 0.37 $\pm$ 0.02 Inc-2 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00 0.35 $\pm$ 0.01 0.34 $\pm$ 0.01 0.32 $\pm$ 0.01 Inc-3 0.03 $\pm$ 0.01 0.15 $\pm$ 0.00 0.42 $\pm$ 0.01 0.43 $\pm$ 0.01 0.41 $\pm$ 0.01 Mix-1 0.03 $\pm$ 0.00 0.16 $\pm$ 0.01 0.50 $\pm$ 0.01 0.54 $\pm$ 0.01 0.47 $\pm$ 0.02 Mix-2 0.02 $\pm$ 0.00 0.13 $\pm$ 0.00 0.47 $\pm$ 0.02 0.46 $\pm$ 0.00 0.39 $\pm$ 0.01 HOPSE-G Adj-1 0.02 $\pm$ 0.00 0.09 $\pm$ 0.01 0.25 $\pm$ 0.01 0.26 $\pm$ 0.01 0.24 $\pm$ 0.00 Adj-2 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.29 $\pm$ 0.01 0.27 $\pm$ 0.01 0.26 $\pm$ 0.01 Adj-3 0.02 $\pm$ 0.00 0.10 $\pm$ 0.00 0.32 $\pm$ 0.01 0.32 $\pm$ 0.02 0.27 $\pm$ 0.01 Inc-1 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.37 $\pm$ 0.01 0.34 $\pm$ 0.00 0.33 $\pm$ 0.00 Inc-2 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.32 $\pm$ 0.01 0.30 $\pm$ 0.05 0.28 $\pm$ 0.02 Inc-3 0.02 $\pm$ 0.00 0.14 $\pm$ 0.01 0.40 $\pm$ 0.00 0.37 $\pm$ 0.00 0.36 $\pm$ 0.02 Mix-1 0.03 $\pm$ 0.00 0.14 $\pm$ 0.00 0.44 $\pm$ 0.01 0.45 $\pm$ 0.00 0.42 $\pm$ 0.00 Mix-2 0.02 $\pm$ 0.00 0.14 $\pm$ 0.00 0.40 $\pm$ 0.02 0.37 $\pm$ 0.01 0.32 $\pm$ 0.01
-Runtime per epoch per neighborhood.
-As expected Table 8 and Table 9 show that the runtime per epoch increases as the number of neighborhoods in ${\mathcal{N}}_{\mathcal{C}}$ increases.
-This aligns with the complexity analysis presented in Appendix E.
-Still, the relative increase in computational time is at most 80%, which is acceptable when considering how many more higher-order relations are being considered.
-The efficiency-expressivity trade-off revealed in these results has important implications for the practical deployment.
-While intuition might suggest that more complex neighborhood configurations would always incur substantial computational penalties, our findings demonstrate that this is not necessarily true.
-Despite incorporating significantly more structural information, complex neighborhood configurations like Mix-1 and Mix-2 introduce only a modest increase in per-epoch runtime, typically less than 2×2$\times$2 × compared to the simplest configurations.
-This makes them viable options even in computationally constrained environments.
-Total runtime per neighborhood.
-The total runtime required to reach convergence, presented in Table 10 and Table 11, reveals more nuanced patterns.
-Unlike per-epoch runtime, total runtime encapsulates both computational cost per epoch and the number of epochs to convergence, which are indirectly influenced by the expressiveness of the neighborhood structure.
-Interestingly, richer neighborhood configurations, despite their potentially higher per-epoch cost, often lead to a similar relative increase in time needed as the per-epoch case.
-This means that the number of epochs needed for convergence stays relatively constant.
-Table 10: Total runtime per neighborhood on simplicial.
-Results are shown as mean $\pm$ standard deviation.
-The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
-HOPSE configurations, especially with mixed neighborhood sets, provide an excellent balance between efficiency and expressiveness.
-Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 6.76 $\pm$ 1.38 44.60 $\pm$ 17.17 164.10 $\pm$ 79.48 125.46 $\pm$ 19.54 1238.09 $\pm$ 378.91 2652.62 $\pm$ 67.51 3028.14 $\pm$ 41.05 5327.90 $\pm$ 1334.91 Adj-2 8.66 $\pm$ 0.51 31.26 $\pm$ 15.10 165.10 $\pm$ 31.84 131.47 $\pm$ 24.84 1512.34 $\pm$ 310.43 3036.30 $\pm$ 1105.71 2535.98 $\pm$ 45.80 4053.39 $\pm$ 718.36 Adj-3 6.87 $\pm$ 1.54 38.41 $\pm$ 8.07 175.02 $\pm$ 49.19 146.19 $\pm$ 27.88 $1$$\pm$ 404.64 2734.33 $\pm$ 156.56 2689.60 $\pm$ 38.86 4266.24 $\pm$ 651.78 Inc-1 8.52 $\pm$ 1.47 41.92 $\pm$ 5.67 194.90 $\pm$ 64.73 206.44 $\pm$ 16.44 1183.26 $\pm$ 135.82 4351.87 $\pm$ 1203.66 3342.21 $\pm$ 88.20 5172.86 $\pm$ 1084.20 Inc-2 10.37 $\pm$ 2.21 40.03 $\pm$ 20.40 161.40 $\pm$ 39.23 180.17 $\pm$ 22.35 895.12 $\pm$ 295.76 2704.44 $\pm$ 23.84 2998.32 $\pm$ 790.12 4859.76 $\pm$ 819.26 Inc-3 8.81 $\pm$ 3.00 39.99 $\pm$ 6.31 209.70 $\pm$ 38.55 218.69 $\pm$ 25.83 1332.58 $\pm$ 316.15 3493.64 $\pm$ 62.70 4076.33 $\pm$ 671.42 5465.95 $\pm$ 1816.61 Mix-1 9.81 $\pm$ 2.93 54.94 $\pm$ 13.50 214.50 $\pm$ 7.54 198.45 $\pm$ 54.21 1288.78 $\pm$ 590.16 4201.85 $\pm$ 851.36 4315.76 $\pm$ 1097.48 6542.09 $\pm$ 1272.01 Mix-2 7.63 $\pm$ 1.71 43.67 $\pm$ 18.29 229.05 $\pm$ 65.67 189.19 $\pm$ 45.39 1434.43 $\pm$ 547.60 3596.38 $\pm$ 519.37 3408.67 $\pm$ 162.45 5377.27 $\pm$ 185.49 HOPSE-G Adj-1 4.56 $\pm$ 0.62 25.49 $\pm$ 6.16 92.97 $\pm$ 27.48 104.94 $\pm$ 28.21 542.03 $\pm$ 94.68 1891.07 $\pm$ 17.81 1885.35 $\pm$ 28.06 2194.20 $\pm$ 42.70 Adj-2 4.92 $\pm$ 0.37 32.03 $\pm$ 9.91 118.32 $\pm$ 17.84 111.18 $\pm$ 18.27 480.02 $\pm$ 100.70 2172.54 $\pm$ 216.62 2096.15 $\pm$ 33.94 2323.75 $\pm$ 222.16 Adj-3 5.58 $\pm$ 1.28 39.85 $\pm$ 17.90 120.88 $\pm$ 31.47 134.65 $\pm$ 49.81 478.06 $\pm$ 107.13 2302.00 $\pm$ 47.08 2385.63 $\pm$ 176.04 2997.61 $\pm$ 824.53 Inc-1 5.80 $\pm$ 0.27 32.71 $\pm$ 8.43 130.06 $\pm$ 22.37 106.98 $\pm$ 27.98 649.05 $\pm$ 114.28 2580.45 $\pm$ 47.63 2471.89 $\pm$ 47.46 3716.82 $\pm$ 1192.83 Inc-2 6.44 $\pm$ 0.54 32.98 $\pm$ 8.25 128.01 $\pm$ 22.39 99.39 $\pm$ 33.07 661.57 $\pm$ 133.52 2489.63 $\pm$ 444.47 2240.50 $\pm$ 43.84 2568.75 $\pm$ 76.80 Inc-3 6.85 $\pm$ 0.98 51.18 $\pm$ 11.23 174.20 $\pm$ 57.02 148.67 $\pm$ 37.09 778.78 $\pm$ 117.23 3038.25 $\pm$ 65.40 2930.43 $\pm$ 430.59 3602.56 $\pm$ 555.32 Mix-1 6.85 $\pm$ 0.72 51.94 $\pm$ 9.59 209.07 $\pm$ 28.40 150.07 $\pm$ 36.64 911.69 $\pm$ 165.22 3408.18 $\pm$ 515.16 3379.79 $\pm$ 111.99 4321.64 $\pm$ 1094.92 Mix-2 5.16 $\pm$ 0.36 35.97 $\pm$ 4.38 168.30 $\pm$ 38.75 142.36 $\pm$ 34.54 716.46 $\pm$ 53.19 3324.16 $\pm$ 1113.48 2794.73 $\pm$ 345.73 3876.34 $\pm$ 1141.15
-Table 11: Total runtime per neighborhood on cell.
-Results are shown as mean $\pm$ standard deviation.
-The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
+Table 6: Performance ablation of $\mathcal{N}_{S}$ on simplicial domain.
+Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) HOPSE-M Adj-1 $83.40\pm 0.07$.12) $76.13\pm 0.03$.09) $75.97\pm 0.01$.18) $75.70\pm 0.01$.17) $0.43\pm 0.00$.01) $85.36\pm 0.08$.46) $83.76\pm 0.00$.94) $90.32\pm 0.00$.09) $68.79\pm 0.00$.78) Adj-2 $85.96\pm 0.05$.55) $76.85\pm 0.01$.75) $76.48\pm 0.01$.44) $75.37\pm 0.00$.98) $0.42\pm 0.00$.03) $94.45\pm 0.00$.82) $89.24\pm 0.00$.22) $91.72\pm 0.01$.17) $78.45\pm 0.01$.92) Adj-3 $84.68\pm 0.05$.51) $76.13\pm 0.03$.01) $75.64\pm 0.00$.76) $75.06\pm 0.00$.84) $0.43\pm 0.00$.01) $96.01\pm 0.00$.54) $88.60\pm 0.00$.77) $91.13\pm 0.01$.20) $79.12\pm 0.01$.11) Inc-1 $84.26\pm 0.05$.55) $76.06\pm 0.03$.05) $75.51\pm 0.01$.03) $75.20\pm 0.00$.93) $0.42\pm 0.00$.01) $96.27\pm 0.00$.28) $87.34\pm 0.00$.74) $90.76\pm 0.01$.57) $78.32\pm 0.01$.13) Inc-2 $85.11\pm 0.07$.22) $75.27\pm 0.01$.88) $75.51\pm 0.00$.53) $74.42\pm 0.00$.80) $0.43\pm 0.00$.02) $95.96\pm 0.00$.29) $86.35\pm 0.02$.15) $91.22\pm 0.00$.06) $79.08\pm 0.00$.75) Inc-3 $84.68\pm 0.06$.96) $76.13\pm 0.01$.84) $76.01\pm 0.00$.64) $74.77\pm 0.01$.21) $0.42\pm 0.00$.01) $96.07\pm 0.00$.19) $85.74\pm 0.01$.03) $91.07\pm 0.00$.10) $77.73\pm 0.01$.69) Mix-1 $82.98\pm 0.03$.98) $76.56\pm 0.02$.84) $75.66\pm 0.00$.32) $75.84\pm 0.00$.50) $0.43\pm 0.00$.01) $93.84\pm 0.01$.76) $87.42\pm 0.00$.64) $91.63\pm 0.00$.93) $80.45\pm 0.01$.06) Mix-2 $83.83\pm 0.04$.90) $76.06\pm 0.02$.31) $75.21\pm 0.01$.30) $75.90\pm 0.00$.48) $0.44\pm 0.00$.01) $95.71\pm 0.01$.41) $89.23\pm 0.00$.15) $90.87\pm 0.01$.07) $82.29\pm 0.00$.46) HOPSE-G Adj-1 $82.98\pm 0.07$.22) $74.05\pm 0.02$.17) $77.18\pm 0.01$.45) $76.21\pm 0.00$.65) $0.44\pm 0.00$.01) $80.43\pm 0.01$.03) $60.92\pm 0.01$.18) $88.18\pm 0.00$.03) $24.59\pm 0.01$.87) Adj-2 $85.53\pm 0.02$.77) $74.05\pm 0.03$.08) $77.51\pm 0.01$.15) $74.91\pm 0.01$.17) $43\pm 0.00.00$) $80.19\pm 0.01$.13) $61.88\pm 0.00$.83) $88.13\pm 0.00$.02) $20.31\pm 0.02$.57) Adj-3 $85.11\pm 0.03$.36) $73.26\pm 0.04$.20) $77.32\pm 0.01$.19) $74.79\pm 0.00$.91) $0.43\pm 0.00$.01) $79.11\pm 0.01$.53) $61.86\pm 0.01$.10) $88.11\pm 0.00$.02) $19.08\pm 0.03$.57) Inc-1 $84.68\pm 0.03$.50) $73.19\pm 0.03$.52) $77.70\pm 0.01$.40) $76.36\pm 0.01$.75) $44\pm 0.00.00$) $80.70\pm 0.00$.80) $62.17\pm 0.00$.98) $87.31\pm 0.01$.17) $0.47\pm 0.03.22$) Inc-2 $82.13\pm 0.05$.75) $76.27\pm 0.02$.10) $75.66\pm 0.01$.52) $74.17\pm 0.01$.99) $0.45\pm 0.00$.02) $78.92\pm 0.01$.25) $59.57\pm 0.01$.56) $88.08\pm 0.00$.08) $20.28\pm 0.03$.48) Inc-3 $77.87\pm 0.04$.90) $75.48\pm 0.03$.15) $6.15\pm 0.01.77$) $75.14\pm 0.00$.75) $0.44\pm 0.00$.01) $78.78\pm 0.00$.74) $60.35\pm 0.01$.05) $88.12\pm 0.00$.04) $18.82\pm 0.03$.54) Mix-1 $79.15\pm 0.07$.88) $75.70\pm 0.03$.24) $76.50\pm 0.01$.09) $75.35\pm 0.00$.97) $0.45\pm 0.00$.01) $79.76\pm 0.00$.51) $60.33\pm 0.01$.60) $88.07\pm 0.00$.09) $23.48\pm 0.03$.44) Mix-2 $83.40\pm 0.08$.56) $74.05\pm 0.03$.39) $77.02\pm 0.00$.92) $74.75\pm 0.01$.47) $0.46\pm 0.00$.01) $81.75\pm 0.01$.26) $61.32\pm 0.02$.67) $88.28\pm 0.00$.08) $35.37\pm 0.02$.25)
+Table 7: Performance ablation of $\mathcal{N}_{S}$ on cell domain.
+Graph Tasks $\mathcal{N}_{S}$ MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) HOPSE-M Adj-1 $83.40\pm 0.04$.85) $75.05\pm 0.02$.33) $76.09\pm 0.01$.99) $76.38\pm 0.00$.73) $0.27\pm 0.00$.01) Adj-2 $83.83\pm 0.03$.87) $77.28\pm 0.02$.44) $76.71\pm 0.00$.42) $75.80\pm 0.01$.71) $0.26\pm 0.00$.01) Adj-3 $85.53\pm 0.04$.85) $76.06\pm 0.03$.24) $76.05\pm 0.01$.75) $75.68\pm 0.00$.93) $25\pm 0.00.00$) Inc-1 $83.83\pm 0.05$.12) $76.99\pm 0.01$.94) $77.88\pm 0.01$.71) $76.32\pm 0.01$.40) $0.26\pm 0.00$.01) Inc-2 $86.38\pm 0.05$.34) $75.84\pm 0.02$.94) $76.63\pm 0.01$.78) $76.03\pm 0.01$.45) $0.26\pm 0.00$.01) Inc-3 $80.43\pm 0.05$.51) $75.70\pm 0.03$.55) $76.77\pm 0.00$.21) $75.78\pm 0.01$.66) $0.25\pm 0.00$.02) Mix-1 $84.26\pm 0.06$.49) $76.63\pm 0.01$.99) $77.80\pm 0.00$.82) $77.08\pm 0.01$.35) $0.24\pm 0.00$.01) Mix-2 $83.83\pm 0.04$.15) $77.06\pm 0.02$.43) $76.52\pm 0.02$.26) $76.22\pm 0.01$.64) $0.24\pm 0.00$.01) HOPSE-G Adj-1 $82.13\pm 0.07$.76) $76.27\pm 0.02$.70) $78.31\pm 0.00$.84) $77.48\pm 0.00$.99) $26\pm 0.00.00$) Adj-2 $87.66\pm 0.06$.96) $76.42\pm 0.02$.19) $78.83\pm 0.01$.40) $77.48\pm 0.00$.64) $24\pm 0.00.00$) Adj-3 $88.09\pm 0.01$.90) $75.91\pm 0.01$.89) $78.35\pm 0.01$.11) $76.94\pm 0.01$.46) $23\pm 0.00.00$) Inc-1 $85.11\pm 0.06$.56) $74.62\pm 0.03$.99) $77.57\pm 0.01$.67) $76.81\pm 0.01$.23) $0.25\pm 0.00$.01) Inc-2 $82.55\pm 0.08$.82) $74.91\pm 0.02$.11) $78.15\pm 0.00$.75) $76.40\pm 0.01$.86) $0.27\pm 0.00$.01) Inc-3 $4.68\pm 0.07.88$) $74.12\pm 0.03$.39) $76.81\pm 0.00$.59) $75.26\pm 0.01$.16) $26\pm 0.00.00$) Mix-1 $88.09\pm 0.03$.23) $76.20\pm 0.02$.62) $79.03\pm 0.00$.73) $76.88\pm 0.00$.95) $23\pm 0.00.00$) Mix-2 $85.11\pm 0.06$.02) $75.63\pm 0.02$.35) $79.07\pm 0.00$.65) $77.81\pm 0.01$.26) $0.21\pm 0.00$.01)
+Appendix I Runtime ablation across different neighborhoods
+I.1 Total runtime
 On HOPSE-G, some more complex neighborhoods will result in faster time, even though they contain more information and thus parameters to be learned.
 This is one of the effects of convergence speed.
-This can be seen in neighborhoods Mix-1 and Mix-2.
-On HOPSE-M, there is no such drastic difference
-Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 7.03 $\pm$ 0.86 37.52 $\pm$ 2.06 144.11 $\pm$ 32.07 142.87 $\pm$ 45.01 1128.14 $\pm$ 166.57 Adj-2 7.66 $\pm$ 0.82 35.68 $\pm$ 7.88 147.04 $\pm$ 24.07 121.89 $\pm$ 18.30 901.44 $\pm$ 203.17 Adj-3 9.78 $\pm$ 2.94 37.85 $\pm$ 10.19 153.94 $\pm$ 29.39 131.39 $\pm$ 36.31 1414.14 $\pm$ 325.91 Inc-1 7.86 $\pm$ 0.54 37.06 $\pm$ 7.05 174.19 $\pm$ 33.02 190.50 $\pm$ 35.45 1173.43 $\pm$ 198.41 Inc-2 6.41 $\pm$ 1.43 36.82 $\pm$ 11.82 220.47 $\pm$ 55.22 128.32 $\pm$ 14.86 1020.04 $\pm$ 131.15 Inc-3 8.30 $\pm$ 0.78 48.24 $\pm$ 18.26 169.79 $\pm$ 41.60 169.78 $\pm$ 37.03 1183.44 $\pm$ 217.72 Mix-1 8.83 $\pm$ 1.18 51.25 $\pm$ 13.72 205.54 $\pm$ 25.51 210.50 $\pm$ 11.62 1313.43 $\pm$ 132.16 Mix-2 7.42 $\pm$ 1.37 44.34 $\pm$ 13.40 203.46 $\pm$ 48.75 193.17 $\pm$ 44.83 1676.92 $\pm$ 398.90 HOPSE-G Adj-1 5.09 $\pm$ 0.93 26.71 $\pm$ 4.79 88.31 $\pm$ 11.27 95.95 $\pm$ 28.16 683.86 $\pm$ 125.70 Adj-2 4.98 $\pm$ 0.47 31.85 $\pm$ 5.43 96.33 $\pm$ 19.91 93.98 $\pm$ 15.00 887.83 $\pm$ 160.74 Adj-3 6.64 $\pm$ 0.89 31.13 $\pm$ 5.56 100.65 $\pm$ 14.14 113.39 $\pm$ 18.75 595.39 $\pm$ 107.46 Inc-1 5.75 $\pm$ 1.45 34.11 $\pm$ 7.89 110.19 $\pm$ 19.33 105.67 $\pm$ 27.00 767.42 $\pm$ 85.39 Inc-2 7.65 $\pm$ 2.32 24.65 $\pm$ 6.74 105.02 $\pm$ 28.61 126.14 $\pm$ 37.05 721.57 $\pm$ 60.10 Inc-3 7.87 $\pm$ 1.47 44.42 $\pm$ 9.81 181.18 $\pm$ 96.45 100.22 $\pm$ 11.50 896.86 $\pm$ 162.01 Mix-1 7.61 $\pm$ 0.68 46.14 $\pm$ 10.02 239.59 $\pm$ 38.24 168.35 $\pm$ 50.01 671.40 $\pm$ 948.29 Mix-2 5.27 $\pm$ 0.57 36.29 $\pm$ 13.05 152.73 $\pm$ 8.26 104.12 $\pm$ 37.76 1229.56 $\pm$ 527.27
-Optimal configurations.
-For simpler graph classification tasks, simpler neighborhood configurations may offer the best efficiency.
-In contrast, more expressive neighborhood sets often provide better efficiency-accuracy trade-offs for complex topological tasks and larger datasets.
-These findings demonstrate that the choice of neighborhood configuration represents a critical design decision not just for model accuracy, but also for computational efficiency.
-The provided analysis presents a principled approach to neighborhood selection based on task requirements and domain characteristics, establishing an essential foundation for applying TDL to new fields.
-F.5.3 Preprocessing time per neighborhood
-Preprocessing time analysis
-The computational costs of extracting positional and structural encodings are summarized in Tables 12 and 13.
-While preprocessing times scale with the complexity of the selected neighborhoods and not only with the number of cells, two points need to be emphasized:
-Table 12: Preprocessing time per neighborhood on simplicial.
-Results are shown as mean time in seconds.
-The best result per domain and model is in bold.
-All Topological Tasks have the same time as they share the same preprocessing time.
-Times increase with neighborhood since with Adj-1 being the fastest as it’s the neighborhood of the underlying graph.
-Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 2.42 35.15 53.37 53.82 142.01 333.65 333.65 333.65 Adj-2 3.24 49.49 72.94 72.80 195.71 823.68 823.68 823.68 Adj-3 3.26 58.49 74.16 73.40 199.37 973.49 973.49 973.49 Inc-1 3.62 79.37 92.40 92.20 229.92 1121.78 1121.78 1121.78 Inc-2 4.39 80.91 111.65 111.49 280.18 1544.42 1544.42 1544.42 Inc-3 5.58 123.97 147.31 148.10 367.95 2233.27 2233.27 2233.27 Mix-1 7.27 165.71 191.94 193.45 489.55 3724.43 3724.43 3724.43 Mix-2 4.99 95.34 118.46 117.35 318.12 2495.14 2495.14 2495.14 HOPSE-G Adj-1 4.54 45.94 103.91 103.47 297.61 913.38 913.38 913.38 Adj-2 7.81 65.54 172.20 172.97 496.73 1676.79 1676.79 1676.79 Adj-3 7.76 92.15 174.23 174.49 502.85 2742.26 2742.26 2742.26 Inc-1 8.05 93.32 181.82 183.44 532.93 2654.36 2654.36 2654.36 Inc-2 11.21 90.87 245.87 248.52 714.05 2813.64 2813.64 2813.64 Inc-3 14.53 142.22 324.62 326.91 952.95 4432.76 4432.76 4432.76 Mix-1 21.02 210.41 474.58 472.70 1381.15 7329.51 7329.51 7329.51 Mix-2 14.28 160.55 314.86 319.78 932.91 5241.30 5241.30 5241.30
-Table 13: Preprocessing time per neighborhood on cell.
-Results are shown as mean time in seconds.
-The best result per domain and model is in bold.
-Similarly to the case of simplicial, the fastest neighborhoods to precompute are the smallest ones.
-Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 2.81 27.16 59.68 60.07 163.66 Adj-2 3.75 40.70 81.96 82.60 221.18 Adj-3 4.44 50.97 96.94 97.38 267.01 Inc-1 4.97 74.99 120.83 121.33 310.83 Inc-2 4.98 75.22 121.62 122.38 312.12 Inc-3 7.07 116.00 182.49 179.43 461.41 Mix-1 9.66 162.30 242.08 244.40 627.80 Mix-2 7.04 89.45 159.44 160.62 441.84 HOPSE-G Adj-1 5.08 38.04 113.21 108.81 315.07 Adj-2 8.34 58.50 189.71 184.54 525.01 Adj-3 11.10 75.56 253.63 248.51 727.66 Inc-1 11.89 85.26 264.02 259.75 734.69 Inc-2 11.81 120.32 292.64 257.43 740.39 Inc-3 18.51 132.25 420.45 403.67 1163.78 Mix-1 28.02 188.99 641.07 621.33 1776.32 Mix-2 20.47 140.60 472.04 460.79 1326.40
-  1. 1.
-    One time preprocessing: Encodings are precomputed once per dataset, enabling efficient experimentation.
-For instance, preprocessing all neighborhoods for ZINC in the cell domain takes 25–30 minutes (e.g., 1,776 seconds for Mix-1 in HOPSE-G), but this cost is amortized over all subsequent training runs.
-The practitioners can freely test configurations like Adj-3 or Inc-2 without recomputing features, streamlining hyperparameter optimization.
-Finally, the precomputed encodings can be shared, reducing redundant computation.
-  2. 2.
-    Parallel computation: The reported precomputation time results can be easily optimized by parallel precomputation.
-F.6 Runtime vs accuracy
-Figure 5 provides a comparative overview of model performance in terms of F1 score against total training runtime for four benchmark tasks: $\beta_{1}$ (a), $\beta_{2}$ (b), NAME (c), and ORIENT (d).
-Figure 6 shows the same arrangement but from the perspective of average runtime per epoch against performance.
-Additionally, Figure 6 shows the convergence speed of any particular model.
-These results clearly show that HOPSE is the best model for an optimal trade-off between runtime and accuracy.
-[…]
-[…]
-Appendix G Licenses
+This can be seen in neighborhoods Mix-1 and Inc-[1,3].
+On HOPSE-M, there is no such drastic difference in HOPSE configurations, especially with mixed neighborhood sets, which provide an excellent balance between efficiency and expressiveness.
+Table 8: Total training runtime (seconds) ablation of $\mathcal{N}_{S}$ on simplicial domain.
+Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 $9.62\pm 0.00$.40) $49.15\pm 0.39$.24) $160.18\pm 0.29$.76) $145.63\pm 0.13$.40) $1488.50\pm 2.83$.52) $1754.92\pm 11.74$.29) $2281.97\pm 4.78$.17) $2498.81\pm 1.86$.09) Adj-2 $10.97\pm 0.01$.91) $49.32\pm 0.16$.14) $141.54\pm 0.34$.29) $177.84\pm 0.49$.45) $2855.15\pm 4.79$.89) $2455.75\pm 4.27$.45) $2478.94\pm 5.20$.52) $2331.82\pm 3.46$.01) Adj-3 $12.66\pm 0.02$.91) $47.95\pm 0.23$.11) $225.36\pm 0.45$.43) $195.86\pm 0.36$.31) $599.04\pm 0.42.11$) $3576.27\pm 3.02$.11) $3043.11\pm 8.31$.36) $2561.17\pm 1.63$.73) Inc-1 $11.50\pm 0.02$.44) $58.85\pm 0.13$.74) $169.46\pm 0.42$.22) $220.22\pm 0.32$.95) $2655.44\pm 4.36$.38) $2963.63\pm 5.10$.20) $2287.67\pm 3.48$.83) $2613.46\pm 1.72$.92) Inc-2 $10.65\pm 0.01$.13) $44.15\pm 0.20$.68) $162.13\pm 0.44$.12) $139.93\pm 0.22$.47) $1923.95\pm 3.35$.87) $3287.92\pm 4.24$.70) $2439.12\pm 4.94$.99) $2178.14\pm 3.33$.52) Inc-3 $11.20\pm 0.01$.06) $61.33\pm 0.21$.24) $236.34\pm 0.21$.93) $194.27\pm 0.37$.37) $2620.02\pm 7.80$.01) $3292.73\pm 4.44$.17) $2825.51\pm 4.50$.02) $2386.66\pm 3.17$.80) Mix-1 $12.16\pm 0.00$.93) $74.15\pm 0.14$.39) $372.57\pm 0.43$.27) $235.41\pm 0.44$.74) $3168.39\pm 5.57$.44) $4205.60\pm 15.93$.84) $6681.04\pm 12.61$.16) $3591.07\pm 4.58$.88) Mix-2 $11.17\pm 0.00$.84) $57.38\pm 0.19$.99) $210.32\pm 0.53$.70) $217.65\pm 0.27$.40) $1630.30\pm 0.75$.76) $3497.96\pm 11.59$.55) $3737.92\pm 2.34$.34) $3590.55\pm 4.68$.47) HOPSE-G Adj-1 $4.56\pm 0.00$.62) $25.49\pm 0.06$.16) $92.97\pm 0.27$.48) $104.94\pm 0.28$.21) $542.03\pm 0.94$.68) $1891.07\pm 0.17$.81) $1885.35\pm 0.28$.06) $2194.20\pm 0.42$.70) Adj-2 $4.92\pm 0.00$.37) $32.03\pm 0.09$.91) $118.32\pm 0.17$.84) $111.18\pm 0.18$.27) $480.02\pm 1.00$.70) $2172.54\pm 2.16$.62) $2096.15\pm 0.33$.94) $2323.75\pm 2.22$.16) Adj-3 $5.58\pm 0.01$.28) $39.85\pm 0.17$.90) $120.88\pm 0.31$.47) $134.65\pm 0.49$.81) $478.06\pm 1.07$.13) $2302.00\pm 0.47$.08) $2385.63\pm 1.76$.04) $2997.61\pm 8.24$.53) Inc-1 $5.80\pm 0.00$.27) $32.71\pm 0.08$.43) $130.06\pm 0.22$.37) $106.98\pm 0.27$.98) $649.05\pm 1.14$.28) $2580.45\pm 0.47$.63) $2471.89\pm 0.47$.46) $3716.82\pm 11.92$.83) Inc-2 $6.44\pm 0.00$.54) $32.98\pm 0.08$.25) $128.01\pm 0.22$.39) $99.39\pm 0.33$.07) $661.57\pm 1.33$.52) $2489.63\pm 4.44$.47) $2240.50\pm 0.43$.84) $2568.75\pm 0.76$.80) Inc-3 $6.85\pm 0.00$.98) $51.18\pm 0.11$.23) $174.20\pm 0.57$.02) $148.67\pm 0.37$.09) $778.78\pm 1.17$.23) $3038.25\pm 0.65$.40) $2930.43\pm 4.30$.59) $3602.56\pm 5.55$.32) Mix-1 $6.85\pm 0.00$.72) $51.94\pm 0.09$.59) $209.07\pm 0.28$.40) $150.07\pm 0.36$.64) $911.69\pm 1.65$.22) $3408.18\pm 5.15$.16) $3379.79\pm 1.11$.99) $4321.64\pm 10.94$.92) Mix-2 $5.16\pm 0.00$.36) $35.97\pm 0.04$.38) $168.30\pm 0.38$.75) $142.36\pm 0.34$.54) $716.46\pm 0.53$.19) $3324.16\pm 11.13$.48) $2794.73\pm 3.45$.73) $3876.34\pm 11.41$.15)
+Table 9: Total training runtime (seconds) of $\mathcal{N}_{S}$ on cell domain.
+Graph Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC-12K HOPSE-M Adj-1 $8.61\pm 0.02$.06) $38.36\pm 0.12$.45) $153.99\pm 0.20$.76) $155.94\pm 0.40$.02) $856.03\pm 2.24$.92) Adj-2 $5.51\pm 0.00$.81) $32.18\pm 0.10$.10) $169.57\pm 0.35$.97) $153.25\pm 0.21$.41) $1013.70\pm 1.20$.74) Adj-3 $5.50\pm 0.00$.41) $33.48\pm 0.06$.54) $175.46\pm 0.22$.23) $153.43\pm 0.17$.47) $922.75\pm 0.89$.20) Inc-1 $6.98\pm 0.01$.24) $41.18\pm 0.15$.30) $144.62\pm 0.29$.66) $192.74\pm 0.45$.31) $952.75\pm 1.66$.32) Inc-2 $6.81\pm 0.00$.88) $36.17\pm 0.09$.11) $154.09\pm 0.55$.24) $135.15\pm 0.29$.51) $827.17\pm 0.77$.46) Inc-3 $7.15\pm 0.01$.07) $44.15\pm 0.16$.85) $234.99\pm 0.25$.73) $157.33\pm 0.30$.17) $934.61\pm 0.49$.49) Mix-1 $8.88\pm 0.01$.71) $30.30\pm 0.02$.29) $163.48\pm 0.27$.28) $222.96\pm 0.59$.41) $1421.85\pm 1.27$.42) Mix-2 $5.54\pm 0.00$.54) $36.30\pm 0.08$.16) $129.82\pm 0.21$.18) $161.53\pm 0.50$.71) $1046.86\pm 0.47$.16) HOPSE-G Adj-1 $5.09\pm 0.00$.93) $26.71\pm 0.04$.79) $88.31\pm 0.11$.27) $95.95\pm 0.28$.16) $683.86\pm 1.25$.70) Adj-2 $4.98\pm 0.00$.47) $31.85\pm 0.05$.43) $96.33\pm 0.19$.91) $93.98\pm 0.15$.00) $887.83\pm 1.60$.74) Adj-3 $6.64\pm 0.00$.89) $31.13\pm 0.05$.56) $100.65\pm 0.14$.14) $113.39\pm 0.18$.75) $595.39\pm 1.07$.46) Inc-1 $5.75\pm 0.01$.45) $34.11\pm 0.07$.89) $110.19\pm 0.19$.33) $105.67\pm 0.27$.00) $767.42\pm 0.85$.39) Inc-2 $7.65\pm 0.02$.32) $24.65\pm 0.06$.74) $105.02\pm 0.28$.61) $126.14\pm 0.37$.05) $721.57\pm 0.60$.10) Inc-3 $7.87\pm 0.01$.47) $44.42\pm 0.09$.81) $181.18\pm 0.96$.45) $100.22\pm 0.11$.50) $896.86\pm 1.62$.01) Mix-1 $7.61\pm 0.00$.68) $46.14\pm 0.10$.02) $239.59\pm 0.38$.24) $168.35\pm 0.50$.01) $671.40\pm 9.48$.29) Mix-2 $5.27\pm 0.00$.57) $36.29\pm 0.13$.05) $152.73\pm 0.08$.26) $104.12\pm 0.37$.76) $1229.56\pm 5.27$.27)
+I.2 Runtime per epoch
+Tables˜10 and 11 validate that smaller neighborhoods, and thus less parameters, result in better per epoch runtimes.
+HOPSE-G has lower per epoch runtime due to having to learn less instantiations of $f_{r,k}$ since $k=1$.
+Simpler configurations such as Adj-1 consistently yield the lowest runtimes, particularly for HOPSE-M, where Adj-1 achieves the fastest epoch times in 5 out of 8 datasets.
+Note that as dataset size increases, the epoch time scales linearly, thus while HOPSE-G remains the fastest, it is only so by a increasingly smaller fraction.
+Table 10: Per epoch runtime (seconds) of $\mathcal{N}_{S}$ on simplicial domain.
+Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC-12K NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 $03\pm 0.00.00$) $0.12\pm 0.00$.04) $0.40\pm 0.00$.01) $0.34\pm 0.00$.01) $0.43\pm 0.00$.03) $3.12\pm 0.00$.11) $3.67\pm 0.00$.07) $3.91\pm 0.00$.14) Adj-2 $03\pm 0.00.00$) $0.12\pm 0.00$.01) $0.44\pm 0.00$.01) $0.37\pm 0.00$.02) $0.89\pm 0.00$.07) $3.52\pm 0.00$.10) $4.15\pm 0.00$.09) $4.20\pm 0.00$.24) Adj-3 $03\pm 0.00.00$) $0.15\pm 0.00$.07) $0.54\pm 0.00$.01) $0.39\pm 0.00$.02) $0.62\pm 0.00$.11) $4.67\pm 0.00$.06) $4.40\pm 0.00$.12) $4.36\pm 0.00$.12) Inc-1 $03\pm 0.00.00$) $0.15\pm 0.00$.01) $0.51\pm 0.00$.02) $0.47\pm 0.00$.06) $0.88\pm 0.00$.06) $5.20\pm 0.00$.24) $4.33\pm 0.00$.12) $4.83\pm 0.00$.12) Inc-2 $03\pm 0.00.00$) $0.14\pm 0.00$.02) $0.45\pm 0.00$.01) $0.44\pm 0.00$.02) $0.64\pm 0.00$.05) $4.58\pm 0.00$.07) $4.18\pm 0.00$.10) $4.39\pm 0.00$.19) Inc-3 $03\pm 0.00.00$) $0.17\pm 0.00$.01) $0.54\pm 0.00$.02) $0.48\pm 0.00$.04) $0.87\pm 0.00$.06) $5.62\pm 0.00$.12) $5.43\pm 0.00$.09) $5.80\pm 0.00$.18) Mix-1 $04\pm 0.00.00$) $0.22\pm 0.00$.02) $0.55\pm 0.00$.01) $0.55\pm 0.00$.01) $1.23\pm 0.00$.13) $5.94\pm 0.00$.32) $8.08\pm 0.00$.15) $6.64\pm 0.00$.32) Mix-2 $03\pm 0.00.00$) $0.15\pm 0.00$.01) $0.50\pm 0.00$.02) $0.46\pm 0.00$.01) $0.61\pm 0.00$.04) $4.82\pm 0.00$.24) $4.72\pm 0.00$.26) $5.65\pm 0.00$.42) HOPSE-G Adj-1 0.02(0.00) $0.10\pm 0.00$.01) 0.28(0.01) $28\pm 0.00.00$) $0.26\pm 0.00$.02) $3.19\pm 0.00$.09) $3.18\pm 0.00$.12) $3.55\pm 0.00$.12) Adj-2 $02\pm 0.00.00$) $0.12\pm 0.00$.01) $30\pm 0.00.00$) $0.31\pm 0.00$.01) $0.28\pm 0.00$.01) $3.53\pm 0.00$.12) $3.41\pm 0.00$.10) $3.67\pm 0.00$.12) Adj-3 $02\pm 0.00.00$) $11\pm 0.00.00$) $0.33\pm 0.00$.02) $0.34\pm 0.00$.01) $0.29\pm 0.00$.02) $3.79\pm 0.00$.14) $3.84\pm 0.00$.14) $3.98\pm 0.00$.07) Inc-1 $02\pm 0.00.00$) $0.13\pm 0.00$.01) $0.38\pm 0.00$.01) $0.37\pm 0.00$.02) $0.34\pm 0.00$.01) $4.44\pm 0.00$.14) $4.23\pm 0.00$.12) $4.64\pm 0.00$.08) Inc-2 $02\pm 0.00.00$) $10\pm 0.00.00$) $31\pm 0.00.00$) $31\pm 0.00.00$) $0.28\pm 0.00$.01) $3.74\pm 0.00$.17) $3.66\pm 0.00$.10) $3.78\pm 0.00$.06) Inc-3 $02\pm 0.00.00$) $13\pm 0.00.00$) $41\pm 0.00.00$) $0.40\pm 0.00$.01) $0.38\pm 0.00$.03) $5.17\pm 0.00$.40) $4.75\pm 0.00$.23) $5.13\pm 0.00$.06) Mix-1 $02\pm 0.00.00$) $0.16\pm 0.00$.01) $0.46\pm 0.00$.02) $0.46\pm 0.00$.01) $0.44\pm 0.00$.01) $5.50\pm 0.00$.24) $5.74\pm 0.00$.22) $5.84\pm 0.00$.38) Mix-2 $02\pm 0.00.00$) $0.13\pm 0.00$.01) $38\pm 0.00.00$) $0.38\pm 0.00$.01) $0.36\pm 0.00$.02) $4.48\pm 0.00$.14) $4.55\pm 0.00$.19) $4.84\pm 0.00$.11)
+Table 11: Per epoch runtime (seconds) of $\mathcal{N}_{S}$ on cell domain.
+Graph Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC-12K HOPSE-M Adj-1 $02\pm 0.00.00$) $12\pm 0.00.00$) $0.39\pm 0.00$.01) $0.44\pm 0.00$.05) $0.27\pm 0.00$.01) Adj-2 $02\pm 0.00.00$) $0.10\pm 0.00$.01) $0.29\pm 0.00$.01) $0.38\pm 0.00$.02) $0.27\pm 0.00$.01) Adj-3 $02\pm 0.00.00$) $0.11\pm 0.00$.01) $0.36\pm 0.00$.01) $0.42\pm 0.00$.02) $0.30\pm 0.00$.01) Inc-1 $02\pm 0.00.00$) $0.12\pm 0.00$.01) $0.40\pm 0.00$.01) $0.47\pm 0.00$.02) $0.44\pm 0.00$.01) Inc-2 $03\pm 0.00.00$) $0.12\pm 0.00$.01) $0.31\pm 0.00$.02) $0.42\pm 0.00$.01) $0.31\pm 0.00$.01) Inc-3 $03\pm 0.00.00$) $0.13\pm 0.00$.01) $0.43\pm 0.00$.02) $0.52\pm 0.00$.02) $41\pm 0.00.00$) Mix-1 $03\pm 0.00.00$) $14\pm 0.00.00$) $0.49\pm 0.00$.01) $0.58\pm 0.00$.02) $0.51\pm 0.00$.02) Mix-2 $03\pm 0.00.00$) $0.13\pm 0.00$.01) $0.42\pm 0.00$.01) $0.50\pm 0.00$.02) $45\pm 0.00.00$) HOPSE-G Adj-1 $02\pm 0.00.00$) $0.09\pm 0.00$.01) $0.25\pm 0.00$.01) $0.26\pm 0.00$.01) $24\pm 0.00.00$) Adj-2 $02\pm 0.00.00$) $09\pm 0.00.00$) $0.29\pm 0.00$.01) $0.27\pm 0.00$.01) $0.26\pm 0.00$.01) Adj-3 $02\pm 0.00.00$) $10\pm 0.00.00$) $0.32\pm 0.00$.01) $0.32\pm 0.00$.02) $0.27\pm 0.00$.01) Inc-1 $02\pm 0.00.00$) $11\pm 0.00.00$) $0.37\pm 0.00$.01) $34\pm 0.00.00$) $33\pm 0.00.00$) Inc-2 $02\pm 0.00.00$) $09\pm 0.00.00$) $0.32\pm 0.00$.01) $0.30\pm 0.00$.05) $0.28\pm 0.00$.02) Inc-3 $02\pm 0.00.00$) $0.14\pm 0.00$.01) $40\pm 0.00.00$) $37\pm 0.00.00$) $0.36\pm 0.00$.02) Mix-1 $03\pm 0.00.00$) $14\pm 0.00.00$) $0.44\pm 0.00$.01) $45\pm 0.00.00$) $42\pm 0.00.00$) Mix-2 $02\pm 0.00.00$) $14\pm 0.00.00$) $0.40\pm 0.00$.02) $0.37\pm 0.00$.01) $0.32\pm 0.00$.01)
+I.3 Preprocessing time
+Tables˜13 and 12 show the preprocessing times for both HOPSE-M and HOPSE-G across various neighborhoods.
+While the tables indicate that HOPSE-M can be faster on these specific benchmarks, this empirical result requires careful contextualization and does not reflect the models’ true asymptotic scalability.
+The primary limitation of HOPSE-M is the high asymptotic complexity of its handcrafted encodings.
+Certain methods require operations like eigendecomposition, which scales at $O(n^{3})$ and becomes computationally prohibitive on larger graphs.
+In contrast, HOPSE-G is designed for greater scalability.
+Its preprocessing consists of a single GNN forward pass with more favorable scaling properties.
+Furthermore, HOPSE-G enhances downstream efficiency by consolidating all structural information into a single channel ($K=1$), which reduces the number of learnable projection functions ($f_{r,k}$) and leads to faster training and inference.
+Therefore, when considering both theoretical complexity and end-to-end efficiency, HOPSE-G is the more scalable framework.
+As expected, the tables show that preprocessing times increase with the complexity of the neighborhood definition, with the simple graph-based Adj-1 neighborhood being the fastest.
+Table 12: Preprocessing time (seconds) of $\mathcal{N}_{S}$ on simplicial domain.
+Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 2.42 35.15 53.37 53.82 142.01 333.65 333.65 333.65 Adj-2 3.24 49.49 72.94 72.80 195.71 823.68 823.68 823.68 Adj-3 3.26 58.49 74.16 73.40 199.37 973.49 973.49 973.49 Inc-1 3.62 79.37 92.40 92.20 229.92 1121.78 1121.78 1121.78 Inc-2 4.39 80.91 111.65 111.49 280.18 1544.42 1544.42 1544.42 Inc-3 5.58 123.97 147.31 148.10 367.95 2233.27 2233.27 2233.27 Mix-1 7.27 165.71 191.94 193.45 489.55 3724.43 3724.43 3724.43 Mix-2 4.99 95.34 118.46 117.35 318.12 2495.14 2495.14 2495.14 HOPSE-G Adj-1 4.54 45.94 103.91 103.47 297.61 913.38 913.38 913.38 Adj-2 7.81 65.54 172.20 172.97 496.73 1676.79 1676.79 1676.79 Adj-3 7.76 92.15 174.23 174.49 502.85 2742.26 2742.26 2742.26 Inc-1 8.05 93.32 181.82 183.44 532.93 2654.36 2654.36 2654.36 Inc-2 11.21 90.87 245.87 248.52 714.05 2813.64 2813.64 2813.64 Inc-3 14.53 142.22 324.62 326.91 952.95 4432.76 4432.76 4432.76 Mix-1 21.02 210.41 474.58 472.70 1381.15 7329.51 7329.51 7329.51 Mix-2 14.28 160.55 314.86 319.78 932.91 5241.30 5241.30 5241.30
+Table 13: Preprocessing time (seconds) of $\mathcal{N}_{S}$ on cell domain.
+Graph Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 2.81 27.16 59.68 60.07 163.66 Adj-2 3.75 40.70 81.96 82.60 221.18 Adj-3 4.44 50.97 96.94 97.38 267.01 Inc-1 4.97 74.99 120.83 121.33 310.83 Inc-2 4.98 75.22 121.62 122.38 312.12 Inc-3 7.07 116.00 182.49 179.43 461.41 Mix-1 9.66 162.30 242.08 244.40 627.80 Mix-2 7.04 89.45 159.44 160.62 441.84 HOPSE-G Adj-1 5.08 38.04 113.21 108.81 315.07 Adj-2 8.34 58.50 189.71 184.54 525.01 Adj-3 11.10 75.56 253.63 248.51 727.66 Inc-1 11.89 85.26 264.02 259.75 734.69 Inc-2 11.81 120.32 292.64 257.43 740.39 Inc-3 18.51 132.25 420.45 403.67 1163.78 Mix-1 28.02 188.99 641.07 621.33 1776.32 Mix-2 20.47 140.60 472.04 460.79 1326.40
+Appendix J Ablations over choices of PSEs
+This section introduces an ablation study over the choices of PSEs as ways to characterize neighborhoods.
+The conditions include experiments for each individual PSE, a mix of all PSEs and a mix of all PSE without LapPE.
+The following results show these combinations over HOPSE-M and graph-based models GCN, GAT and GIN.
+J.1 Over HOPSE-M
+Table 14: Ablation in HOPSE-M using different PSEs on Graph Tasks.
+PSE MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) Simplicial EStat $85.11\pm 0.06$.20) $74.70\pm 0.03$.03) $75.21\pm 0.01$.30) $74.04\pm 0.01$.59) $0.53\pm 0.00$.03) HKDiag $80.43\pm 0.04$.61) $76.13\pm 0.03$.09) $75.66\pm 0.00$.32) $75.90\pm 0.00$.48) $0.42\pm 0.00$.06) LapPE $78.30\pm 0.04$.09) $75.41\pm 0.01$.75) $73.99\pm 0.00$.90) $72.99\pm 0.00$.67) $64\pm 0.00.00$) RWSE $83.40\pm 0.07$.12) $75.20\pm 0.03$.48) $76.44\pm 0.00$.93) $74.39\pm 0.00$.86) $0.42\pm 0.00$.01) All w/o LapPE $86.38\pm 0.06$.31) $75.13\pm 0.03$.66) $77.35\pm 0.00$.81) $76.34\pm 0.01$.40) $0.39\pm 0.00$.02) Cell EStat $80.43\pm 0.05$.51) $75.20\pm 0.03$.12) $76.52\pm 0.02$.26) $76.19\pm 0.01$.75) $0.28\pm 0.00$.01) HKDiag $83.40\pm 0.06$.96) $76.99\pm 0.01$.94) $77.80\pm 0.00$.82) $76.53\pm 0.01$.58) $0.25\pm 0.00$.02) LapPE $58.94\pm 0.26$.35) $66.45\pm 0.09$.00) $62.95\pm 0.13$.52) $62.19\pm 0.12$.76) $35\pm 0.00.00$) RWSE $86.38\pm 0.05$.34) $73.33\pm 0.02$.34) $76.05\pm 0.01$.75) $75.45\pm 0.01$.45) $25\pm 0.00.00$) All w/o LapPE $88.09\pm 0.05$.55) $76.99\pm 0.02$.63) $78.09\pm 0.01$.44) $77.27\pm 0.01$.01) $0.25\pm 0.00$.01) All $85.53\pm 0.05$.51) $73.62\pm 0.02$.59) $76.32\pm 0.01$.74) $77.08\pm 0.01$.35) $0.24\pm 0.00$.01)
+Table 15: Ablation in HOPSE-M using different PSEs on Topological Tasks.
+PSE NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}(\uparrow)$ $\beta_{2}(\uparrow)$ EStat $81.69\pm 0.06$.38) $66.34\pm 0.12$.46) $90.15\pm 0.00$.30) $67.94\pm 0.01$.24) HKDiag $79.18\pm 0.06$.13) $77.27\pm 0.00$.91) $89.65\pm 0.00$.51) $59.59\pm 0.07$.16) LapPE $95.51\pm 0.00$.23) $89.23\pm 0.00$.15) $91.02\pm 0.01$.43) $81.47\pm 0.01$.16) RWSE $86.78\pm 0.07$.97) $78.59\pm 0.02$.46) $90.12\pm 0.00$.17) $65.81\pm 0.01$.40) All $96.27\pm 0.00$.28) $89.24\pm 0.00$.22) $90.87\pm 0.01$.07) $82.29\pm 0.00$.46)
+J.2 Over GNNs
+Table 16: Ablation on GNNs using different PSEs on Graph Tasks.
+PSE MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) GCN EStat $78.72\pm 0.04$.76) $74.84\pm 0.02$.96) $73.95\pm 0.01$.30) $73.51\pm 0.01$.32) $0.60\pm 0.00$.01) HKDiag $70.21\pm 0.05$.42) $74.27\pm 0.00$.64) $74.32\pm 0.00$.63) $72.26\pm 0.01$.27) $0.42\pm 0.00$.04) LapPE $60.99\pm 0.25$.49) $66.19\pm 0.07$.71) $62.73\pm 0.13$.28) $61.52\pm 0.12$.02) $0.62\pm 0.00$.01) RWSE $79.15\pm 0.09$.44) $73.62\pm 0.03$.56) $74.88\pm 0.00$.54) $74.70\pm 0.00$.79) $0.43\pm 0.00$.03) All w/o LapPE $77.02\pm 0.04$.09) $75.77\pm 0.02$.98) $73.42\pm 0.01$.37) $73.61\pm 0.01$.09) $0.48\pm 0.00$.03) All $84.26\pm 0.03$.56) $67.50\pm 0.08$.87) $62.03\pm 0.09$.91) $62.91\pm 0.10$.78) $0.58\pm 0.00$.05) GAT EStat $74.89\pm 0.04$.85) $74.84\pm 0.02$.02) $76.19\pm 0.01$.55) $74.48\pm 0.01$.05) $0.59\pm 0.00$.01) HKDiag $73.19\pm 0.04$.15) $74.70\pm 0.01$.55) $77.86\pm 0.01$.02) $75.57\pm 0.00$.60) $0.39\pm 0.00$.02) LapPE $61.23\pm 0.26$.12) $65.84\pm 0.08$.32) $62.55\pm 0.13$.11) $61.55\pm 0.12$.09) $0.61\pm 0.00$.01) RWSE $76.17\pm 0.08$.82) $73.84\pm 0.03$.16) $75.80\pm 0.01$.26) $73.57\pm 0.01$.07) $0.44\pm 0.00$.04) All w/o LapPE $78.30\pm 0.04$.61) $75.48\pm 0.03$.44) $75.76\pm 0.01$.68) $75.57\pm 0.01$.48) $0.47\pm 0.00$.04) All $80.43\pm 0.05$.08) $66.47\pm 0.08$.19) $62.91\pm 0.13$.48) $62.74\pm 0.13$.31) $0.56\pm 0.00$.02) GIN EStat $77.45\pm 0.08$.85) $73.91\pm 0.03$.31) $75.64\pm 0.01$.51) $74.81\pm 0.01$.00) $0.55\pm 0.00$.05) HKDiag $77.45\pm 0.10$.27) $74.77\pm 0.02$.67) $74.86\pm 0.00$.82) $74.39\pm 0.01$.11) $0.36\pm 0.00$.04) LapPE $60.99\pm 0.25$.69) $65.91\pm 0.08$.38) $62.63\pm 0.13$.17) $62.12\pm 0.12$.65) $0.53\pm 0.00$.02) RWSE $74.89\pm 0.10$.68) $72.83\pm 0.03$.23) $73.52\pm 0.02$.87) $74.00\pm 0.01$.69) $0.41\pm 0.00$.14) All w/o LapPE $76.17\pm 0.07$.12) $75.13\pm 0.02$.22) $74.30\pm 0.00$.84) $75.55\pm 0.01$.07) $0.44\pm 0.00$.02) All $83.83\pm 0.03$.56) $73.76\pm 0.03$.38) $75.84\pm 0.01$.51) $73.90\pm 0.01$.59) $0.54\pm 0.00$.10)
+Table 17: Ablation on GNNs using different PSEs on Topological Tasks.
+PSE NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}(\uparrow)$ $\beta_{2}(\uparrow)$ GIN EStat $81.39\pm 0.00$.92) $56.73\pm 0.01$.34) $88.27\pm 0.00$.10) $24.27\pm 0.04$.41) HKDiag $79.73\pm 0.01$.72) $56.27\pm 0.00$.45) $88.39\pm 0.00$.16) $27.50\pm 0.06$.52) LapPE $91.44\pm 0.01$.66) $81.68\pm 0.00$.96) $89.77\pm 0.00$.05) $63.65\pm 0.00$.49) RWSE $77.37\pm 0.02$.15) $68.89\pm 0.01$.48) $88.50\pm 0.00$.24) $32.82\pm 0.09$.71) All w/o LapPE $82.03\pm 0.00$.68) $60.09\pm 0.04$.87) $88.26\pm 0.00$.09) $24.20\pm 0.04$.31) All $94.05\pm 0.00$.15) $82.53\pm 0.00$.78) $90.28\pm 0.00$.05) $68.43\pm 0.00$.59)
+Appendix K Model parameters.
+Table 18: Model sizes (in thousands of parameters) corresponding to the best-performing hyperparameter configurations.
+Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $\beta_{0},\beta_{1},\beta_{2}$ Graph GCN 67.97K 13.19K 6.72K 23.75K 22.59K 3.56K 3.46K 3.49K GAT 22.02K 10.11K 11.20K 11.23K 22.85K 3.62K 3.52K 51.08K GIN 38.40K 13.19K 154.37K 154.50K 10.40K 3.56K 13.06K 13.12K GIN + PSEs 43.97K 29.82K 13.89K 16.22K 905.21K 128.26K 616.70K 485.37K GAT + PSEs 35.77K 13.15K 11.68K 176.38K 130.56K - - - GCN + PSEs 15.33K 12.03K 16.29K 33.47K 140.54K - - - Simplicial SCN 20.03K 10.24K 94.98K 26.08K 24.42K 348.55K 22.53K 348.29K SCCNN 27.11K 26.72K 188.99K 49.54K 1453.82K 362.88K 362.69K 1446.40K SaNN 286.08K 1523.97K 626.95K 632.19K 995.59K 891.14K 890.75K 890.88K GCCN 15.11K 46.27K 68.99K 49.63K 67.42K 1758.60K 953.35K 6957.31K HOPSE-M 1169.67K 1175.81K 2983.94K 809.86K 764.80K 579.46K 1198.08K 4593.92K HOPSE-G 359.04K 365.83K 1397.25K 1795.84K 666.50K 1370.37K 658.95K 2610.95K Cell CCCN 284.29K 34.56K 63.87K 17.67K 287.74K - - - CWN 334.72K 101.12K 124.10K 412.29K 88.06K - - - GCCN 15.11K 46.27K 68.99K 49.63K 39.78K - - - HOPSE-M 558.34K 2957.83K 809.47K 2189.06K 864.64K - - - HOPSE-G 1405.95K 357.89K 1842.69K 1445.63K 690.30K - - -
+To ensure a fair comparison, our hyperparameter search, guided by the TopoBench framework, constrained all models to a comparable maximum parameter budget.
+The final parameter s for the best-performing configuration of each model are detailed in Table˜18.
+Interestingly, the search process often selected more compact models for the graph-based baselines, a finding consistent with the original TopoBench study.
+In contrast, the best-performing models for HOPSE were comparatively larger.
+We assume this is because HOPSE’s architecture is not constrained by the over-smoothing and over-squashing issues that typically limit the depth of message-passing-based methods.
+Despite its empirically larger size, HOPSE maintains a has a very competitive execution runtime, highlighting the scalability of its design.
+We believe this efficiency is a crucial feature for enabling the application of Topological Deep Learning on large and complex datasets.
+Appendix L Licenses
 This work does not distribute existing datasets.
@@
 The MANTRA benchmarking dataset is licensed under the 3-Clause BSD license.
-The CSL dataset does not specify a particular license aside from the CC BY 4.0 from [17].
-Appendix H Author contributions.
+Appendix M Author contributions.
 M.C.